### v1

- DoFs
- degrees of freedom
- 1D
- one-dimensional
- 3D
- three-dimensional
- PCC
- piecewise constant curvature
- PD
- proportional-derivative
- FEM
- finite element method
- MOR
- model order reduction
- ROM
- reduced order model
- EoM
- equation of motion
- PDE
- partial differential equation
- ODE
- ordinary differential equation
- ODEs
- ordinary differential equations
- BVP
- boundary value problem
- CRT
- Cosserat rod theory
- KRT
- Kirchhoff rod theory
- EBRT
- Euler-Bernoulli rod theory
- PCS
- piecewise constant strain
- DER
- Discrete Elastic Rod
- CSM
- continuum soft manipulator
- PCM
- parallel continuum manipulator
- SMA
- shape memory alloy
- TCP
- twisted and coiled polymer
- CTR
- concentric tube robot
- ETR
- eccentric tube robot
- MB
- Model-Based
- MP
- Multi-Point
- TT
- trajectory tracking
- MPC
- Model Predictive Control
- SMC
- Sliding Mode Control
- ESC
- Energy-Shaping Control
- FL
- Feedback Linearization
- PH
- Pythagorean Hodograph
- ML
- Machine Learning
- RL
- Reinforcement Learning
- SL
- Supervised Learning
- ANN
- artificial neural network
- RNN
- recurrent neural network
- MLP
- multi layer perceptron
- DQL
- deep Q-learning
- DDPG
- deep deterministic policy gradient
- PPO
- proximal policy optimization
- DR
- domain randomization
- MIS
- minimally invasive surgery
Carlo Alessi & Egidio Falotico, Brain-Inspired Robotics Laboratory, The BioRobotics Institute, Scuola Superiore Sant’Anna, Pisa, Italy.
Rod models in continuum and soft robot control: a review
Abstract
Continuum and soft robots can positively impact diverse sectors, from biomedical applications to marine and space exploration, thanks to their potential to adaptively interact with unstructured environments. However, the complex mechanics exhibited by these robots pose diverse challenges in modeling and control. Reduced order continuum mechanical models based on rod theories have emerged as a promising framework, striking a balance between accurately capturing deformations of slender bodies and computational efficiency. This review paper explores rod-based models and control strategies for continuum and soft robots. In particular, it summarizes the mathematical background underlying the four main rod theories applied in soft robotics. Then, it categorizes the literature on rod models applied to continuum and soft robots based on deformation classes, actuation technology, or robot type. Finally, it reviews recent model-based and learning-based control strategies leveraging rod models. The comprehensive review includes a critical discussion of the trends, advantages, limits, and possible future developments of rod models. This paper could guide researchers intending to simulate and control new soft robots and provide feedback to the design and manufacturing community.
keywords:
Soft Robotics; Modeling, Control, and Learning for Soft Robots; Biologically-Inspired Robots1 Introduction
Continuum robots are defined as robots with distributed deformations throughout their structure, resulting in infinite degrees of freedom (DoFs). This characteristic gives rise to a hyper-redundant configuration space, allowing the robot tip to reach any point in the three-dimensional (3D) workspace with a virtually infinite number of configurations (Trivedi et al., 2008c). Soft robots, built with soft materials or deformable structures (Laschi et al., 2016), are continuum robots. Generally speaking, soft robots assume a variety of morphologies. In this paper, we refer to both classes with their partial overlap.
Continuum and soft robots deform elastically while safely interacting with unstructured environments thanks to the inherent compliance of soft materials (Laschi and Cianchetti, 2014; Laschi et al., 2016). Novel design and manufacturing technologies are accelerating the prototyping of highly dexterous robotic systems (Rus and Tolley, 2015; Wang et al., 2024). However, the modeling and control of continuum and soft robots are still challenging problems. Indeed, soft materials cause extreme hyper-redundancy and exhibit nonlinear properties such as hysteresis and stress softening. These challenges inspired the development of diverse modeling techniques (Armanini et al., 2023; Gilbert, 2021), each characterized by underlying assumptions, mathematical framework, and a trade-off between accuracy and computational cost. Soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical (Fig. 1). Concurrently, innovative control strategies emerged to leverage these models (George Thuruthel et al., 2018; Della Santina et al., 2023).
Popular modeling methods are inherently data-driven (Laschi et al., 2023). Despite not requiring geometric and physical expertise, they rely on pseudo-random motion data to train artificial neural networks mapping actuation to task space (Chin et al., 2020). This robot-independent method effectively derived efficient forward models employed in learned policies for tracking (Centurelli et al., 2022; Piqué et al., 2022) and throwing (Bianchi et al., 2023). However, they are black-box, could suffer overfitting, and the burden of collecting interaction datasets limits their applicability.
Other approaches, like discrete material models, discretize continuous bodies a priori. For example, pseudo-rigid models represent them with a chain of rigid links connected by joints. While these provide satisfactory results for hyper-redundant arms (Venkiteswaran et al., 2019), they poorly approximate elastic structures. In addition, lumped-mass models employ an array of masses, springs, and dampers (Habibi et al., 2020). This modularity can model complex phenomena but could require intensive system identification.
Alternatively, geometrical models utilize geometrical assumptions on the deformed shape of the soft body. For instance, functional models describe the deformations using polynomials (Della Santina and Rus, 2019). In addition, piecewise constant curvature (PCC) models discretize continuous bodies in circular arcs with constant curvature (Xie et al., 2022). These models were effectively used within proportional-derivative (PD) task space controllers for various dynamic tasks (Fischer et al., 2022). Since most continuum robots have actuators parallel to the centerline producing constant-curvature deformations, PCC is viable for uniform and lightweight robots (Webster III and Jones, 2010). However, the suitability of these approximations could degrade when robots are subject to significant external forces or interactions with the environment.
Finally, continuum mechanical models characterize soft robots with continuous configuration spaces and define deformations in physical terms. They enable the simulation of physical interactions and the study of robot mechanics. The classical finite element method (FEM) accurately represents complex 3D geometries to the extent of high computational costs and involved mathematical formulations (Coevoet et al., 2017). Since such models would impede real-time control applications, their use is often limited to designing and simulating soft robotic components (Xavier et al., 2021) or as a benchmark. Nonetheless, recent model order reduction (MOR) techniques are making 3D mechanical models more affordable (Allard et al., 2007; Dubied et al., 2022). Since continuum and soft robots are slender, other methods employ reduced order models of elastic rods, one-dimensional (1D) objects with a dimension (the length) much larger than the others. Elastic rods effectively describe the deformations of slender bodies (i.e., stretching, shearing, bending, twisting), balancing computational efficiency and an accurate representation of complex mechanics. Therefore, rod theories are promising candidates for soft robotics modeling and control, which are the focus of this paper.
1.1 Article Contribution
In this study, we investigate the role of rod theories in the modeling, simulation, and control of continuum and soft robots. The main contributions are:
1.
review of the fundamentals of four major rod theories with a common formulation, which facilitates the comparison and highlights connections to other modeling strategies;
2.
review and analysis of rod models applied to continuum and soft robots, originally classified on deformation classes; and
3.
review and analysis of recent model-based and learning-based controllers leveraging rod models.
First, we summarize the mathematical fundamentals underlying the four main rod theories applied in soft robotics: Cosserat-Reissner, Kirchhoff-Love, Timoshenko, and Euler-Bernoulli (Sec. 2). The same notation facilitates finding similarities and differences between the rod theories and supports the analysis of deformation classes. Further, it highlights theoretical insights on spatial discretization (i.e., positions vs configurations), connections to other models (e.g., the polynomial curvature), and implications on the control problem. Also, we consider the general modeling of actuation-induced distributed loads with different actuator routings (Sec. 3). Then, we review and discuss rod models applied to continuum and soft robots classified by deformation, and eventually robot types or actuation systems (Sec. 4). The analyzed models could guide researchers intending to simulate new soft robots in selecting the most suited rod theory. In addition, it could provide feedback to the design and manufacturing community. Finally, we review and discuss recent model-based and learning-based controllers that leverage rod models (Sec. 5). Along the sections, we critically discuss trends, research gaps, advantages, and limitations of rod models. The paper concludes with an overview of the main remarks (Sec. 6).
1.2 Related Surveys
Soft robotics has been the subject of extensive surveys, reviews, and perspectives, which represent a wealth of knowledge and a growing interest in a thriving research field.
The seminal paper by Webster III and Jones (2010) reviews PCC kinematic models of continuum robots, emphasizing the robot-specific and robot-independent mappings and providing a perspective on further developments. Conversely, our work focuses on rod models, which generalize PCC approaches in case of significant external forces.
Recently, Armanini et al. (2023) presented a structured overview of all modeling approaches in soft robotics. They propose a classification in four model classes (continuum mechanics, geometrical, discrete material, and surrogate), providing the fundamental theoretical grounds and analyzing their applicability. However, the control problem is only touched marginally. The perspective by Mengaldo et al. (2022) provide a unified formulation for modeling the embodied intelligence of soft robots. It broadly overviews the theories explaining the practical significance of embodied intelligence. Conversely, this paper is a review/survey article specific to rod-based models and controllers. Another work reviews state-of-the-art methods for soft robot modeling and control, emphasizing the use of real-time FEM (Schegg and Duriez, 2022). Also, Zhang et al. (2022) broadly covers a long soft robotics pipeline: design, actuation methods, modeling, and control. It focuses on major technical concerns, providing a perspective on the development of continuum robots. Finally, Sadati et al. (2023) survey MOR techniques and ROM, clarifying their disentangled usage in soft robotics. Within this framework, they broadly overview modeling approaches, constitutive laws, and solution strategies. While they survey the theoretical aspects of modeling, we also provide insights into the experimental validation of rod models.
Regarding control, George Thuruthel et al. (2018) provide a broad overview of model-based and model-free control strategies for continuum soft manipulators, without concern for the modeling. Also, Wang and Chortos (2022) provide a broad review of actuation mechanisms and control strategies, including open-loop, closed-loop, and autonomous methods. They discuss emergent directions in the control-actuator interface, underactuation, and utilization of artificial intelligence. Della Santina et al. (2023) describe the control problem of CSM specifically for the model-based view. They present a unified formulation of soft robot dynamics independent of the modeling technique. Then, they discuss the shape control and tracking problems and explore open challenges (e.g., under-actuation, environment interactions, actuator dynamics, task space control) while surveying the literature. Conversely, we review and discuss rod-based controllers utilizing model-based strategies and deep Reinforcement Learning (RL).
In summary, this review on rod models complements related works with minor intersections.
2 Background on Rod Theories
An elastic rod is a quasi 1D body in which the length is much larger than the radius of its cross-section. Therefore, the only deformation type considered is along one axis, neglecting other deformations like cross-section deformation. In nature, there are many examples of slender elastic bodies (e.g., hairs, muscle fibers, DNA strands, flagella). In the artificial world, notable examples are continuum and soft robots. In most cases, characterizing the behavior of these robots as that of elastic rods provides a good approximation. The four main Rod Theories used in soft robotics are: Cosserat-Reissner, Kirchoff-Clebsch-Love, Euler-Bernoulli, Timoshenko-Ehrenfest. These models differ in the kinematic assumptions and strain modes considered. An elastic rod might exhibit (i) bending, (ii) torsion, (iii) stretching (i.e., elongation or compression), and (iv) shear. Fig. 2 shows the four rod theories and the modeled strain modes. In this work, we adopted frequently the Lie Algebra notation. The reader can find in Appendix B the definition of the employed operators.
2.1 Problem Statement
An Elastic Rod can be kinematically described through a time-variant 3D curve, usually called backbone. The length of this curve can be parameterized by a material curvilinear abscissa . Every point of the backbone can be written w.r.t. an inertial frame . Usually, in literature, the point of the backbone at is called base and at is called tip of the robot. Furthermore, every cross-section has associated geometrical and material quantities: the cross-section area , the second moment of area about the local axes ; the mass density of the rod , the Young Modulus of the material , and the shear Modulus , where is the Poisson coefficient.
2.2 Cosserat Rod Theory
The Cosserat rod theory (CRT) (Cosserat and Cosserat, 1896) describes all strain modes of the rod. This model is useful for describing CSM that interact with cluttered environments, in which shear and elongation deformations play a fundamental role. Each cross-section is associated with a reference system , whose axes are called directors. The relative roto-translation between and is expressed by the homogeneous matrix defined as
where is the rotation matrix that represents relative rotation. The main idea of the CRT is to allow every cross-section to freely rotate and translate relatively.
2.2.1 Kinematics
Let be the strain twist defined as
where are the angular strain modes (bending and torsion), are the linear strain modes (shear and elongation/compression), and is the spatial partial derivative.
Similarly, it is possible to define the velocity twist as
where are the angular and linear velocities of cross-section , and is the time partial derivative. The strain and velocity twists are both expressed in the local frame and represent the evolution of the rod over space and time. Thanks to the mixed partial derivatives equality, it is possible to derive a kinematic relation between the strain and velocity twist, such as
where denotes an integration variable instead of the material abscissa .
From a mathematical point of view, the main idea of the CRT can be well condensed in the Configuration Space, defined as
Definition (5) indicates the infinite DoFs of an elastic rod. The Configuration Space results in a functional space of curves in (Boyer et al., 2020). A Configuration Space with an infinite dimension is a unique feature of soft robots, making control very challenging.
2.2.2 Dynamics
To describe the equations of motion of an elastic rod, it is necessary to define three distributed wrenches along the length of the rod:
The Internal Forces Wrench expresses the internal load applied by the material, including elastic and damping effects. Assuming small strains, a linear viscoelastic constitutive model relates the strain field and the internal forces wrench .
(6) where is the stress-free strain twist, is the stiffness matrix and is the viscosity matrix.
The External Forces Wrench represents the distributed external load applied to the Rod. For instance, the effect of gravity can be computed as
(7) where is the gravity acceleration twist w.r.t. the inertial frame and is the cross-sectional inertia matrix.
The Actuation Forces Wrench expresses the internal active forces exerted by the actuators. Refer to Sec. 3 for more details.
After defining these quantities, it is possible to compute the Dynamics of an elastic rod using the Poincaré equations (Renda et al., 2018a)
The EoM (8) is a set of partial differential equations of a Cosserat Rod in the local frame. It is common also to find the same set of PDE in an explicit form and expressed in the inertial frame, such as
where denotes a variable expressed in the inertial frame, is the second moment of the area expressed in the inertial frame, and and are the contributions of the internal and actuation forces.
2.2.3 Strain Parameterization
The Configuration Space (5) and PDE (8) show the infinite DoFs of an elastic rod. From a control and numerical implementation perspective, it is convenient to find a method to discretize the continuous nature of the soft segment. Renda et al. (2020) and Boyer et al. (2020) proposed a Strain Parameterization, expressing the Kinematics and the Dynamics of the elastic rod in terms of strain twist . Let us again consider the definition of the strain twist (2), which is easily rewritten as . Assuming that the initial condition is known, the function can be univocally determined by the strain twist. The Configuration Space can be considered the Shape Space of the elastic rod, which is a functional space of the -parameterized curves in , such as
This functional space can be generated by an infinite-dimension basis matrix , such as
where and is a vector of generalized coordinates, with . The main idea of the discretization technique is to truncate the basis matrix to a finite number of columns, reducing the discretized Shape Space to . This approach allows the user to choose the degree of approximation or neglect certain strain modes. To solve the PDE, the proposed method uses the Magnus Expansion (Haier et al., 2006), resulting in a convenient Product of Exponentials, which is widely used in classical robotics (Murray et al., 1994)
where denotes the Magnus expansion of the strain twist. Thanks to that, it is possible to rewrite the Differential Kinematics with a well-known form, such as
where denotes the Soft Geometric Jacobian. Similarly, the Dynamics can be rewritten in a classical Lagrangian form
The EoM (14) can be solved using a standard time solver, such as Runge-Kutta or Explicit Euler, as implemented in the SoRoSim simulator (Mathew et al., 2022).
2.2.4 Discrete Elastic Rod
Another method to discretize the continuum nature of elastic rods was introduced in the pioneering work of Gazzola et al. (2018). First, the Authors derived the EoM (9), including an elongation/compression ratio defined as , where is the curvilinear abscissa in the rest configuration. This ratio is present because the length is parameterized using the curvilinear abscissa differently from (9). In the presence of axial stretching, the stretching ratio scales the geometrical quantities
where the bar sign indicates the geometric quantities in the rest configuration. The EoM can be finally written as
It is worth highlighting that (16) contains an additional contribution with respect to (8), which depends on the time derivative of stretching ratio . Furthermore, scaling the geometrical quantities partially relaxes the assumption of rigid cross-sections. To numerically resolve the EoM (16), they extended for a Cosserat rod the spatial discretization algorithm proposed in a previous study (Bergou et al., 2008), which discretizes the rod in a sequence of rigid segments connecting nodes. For each node, the EoM consider the interactions with the other nodes and external forces. In addition, it is possible to associate kinematic and dynamic quantities to each node and segment to solve (16). In the discrete domain, some quantities, such as curvature, must be expressed in an integrated form over the domain (Gazzola et al., 2018). This domain corresponds to the Voronoi region associated with the interior nodes ,
where is the length of the -th segment. Then, the discrete curvature and bending stiffness matrix can be written as
with . Finally, the discretized EoM of the rod can be rewritten in algorithmic form as follows:
where is the point-wise mass associated with the node, is a domain dilation factor, and are the discrete difference and the averaging operator defined in (Gazzola et al., 2018). In the first equation of (19), , and . In contrast, in the second equation, ranges , while still ranges from . This distinction arises from the definition of discrete curvature and bending stiffness within the interior nodes set.
For the time integration of the discretized EoM (19), the authors proposed the use of a Second-Order Position Verlet time integrator, which exhibits a good balance between numerical accuracy and computational cost (Gazzola et al., 2018). To ensure numerical stability, they proposed an empirical law to choose the integration frequency as . The discrete EoM (19) were implemented in the PyElastica simulator (Naughton et al., 2021).
2.3 Kirchhoff Rod Theory
The Kirchhoff rod theory (KRT) (Love, 1906) is a special case of CRT that considers an elastic rod unstretchable and unshearable. It is particularly suitable for CSM that bend around any axis and twist. Notably, it introduces the notion of a directed curve, assigning a specific reference system to each cross-section. The EoM of Kirchhoff rods can be computed by specializing the EoM of Cosserat, that is (8) or (9). In particular, the constraint of an unstretchable and unshearable rod can be written as
This constraint can be added in the EoM (16), resulting in
The linear internal force serves as a Lagrangian multiplier and it is a virtual internal force that constrains the rod from stretching or shearing. In the case of Strain Parameterization, it is sufficient to apply the constraint (20) in the strain twist
Therefore, the basis function matrix only generates the angular strain mode vector .
2.4 Euler-Bernoulli Rod Theory
The Euler-Bernoulli rod theory (EBRT) (Timoshenko, 1983) is one of the simplest rod theories in which the rod can only bend around one axis. It can be considered the 2D case of the KRT without the twist. Here, the assumption is that the slope angle of the rod is equal to the tangent angle of the backbone curve. Below, we report linear and nonlinear versions.
2.4.1 Linear Euler-Bernoulli
Let us consider the rod in the - plane. Unlike the previous theories, the backbone curve is described by the displacement from the -axis. Recalling the assumption of the EBRT, the slope angle of the beam can be written as
From the minimization of the strain energy, the EoM of the rod can be computed as
where is the distributed external moments around the -axis. The EoM (24) is linear, assuming that the constitutive law of the bending moment is
where is the -component of the internal moment . For control purposes, the linearity of (24) facilitates the use of efficient controllers from the classic control theory (Doroudchi et al., 2018).
2.4.2 Nonlinear Euler-Bernoulli
In the nonlinear EBRT (or Euler’s Elastica), the rod length is parameterized by the arclength . Recalling the Fundamental Theorem of the local theory of Curves in the 2D case, every regular curve can be determined by the curvature (do Carmo, 1976). From this concept, Della Santina and Rus (2019) reformulated Elastica in a classical robotic formulation. Let be the -component of the curvature twist . The Cartesian pose of every cross-section is uniquely determined by the curvature, i.e.
Furthermore, from (26), it is possible to derive the EoM following the Lagrangian approach.
Similarly to the Strain Parameterization, they propose to consider the curvature as an infinite sum of monomials in , i.e.
The geometrical meaning of the polynomial curvature is to constrain the shape of the backbone curve to be a Generalized Cornu Spiral, i.e., a curve with a polynomial curvature (Della Santina and Rus, 2019). It is worth highlighting that the (27) is equivalent to write (11) with a polynomial basis function, related only to the curvature (Caradonna et al., 2024). Defining as joint variables, it is possible to write the EoM as
where is transposed orientation jacobian for actuators and are the pure actuators’ torque.
2.5 Timoshenko Rod Theory
The Timoshenko-Ehrenfest rod theory (Timoshenko and Goodier, 1951) extends the EBRT, relaxing the unshearability constraint. Therefore, the equality between the tangent angle and the cross-sections’ angle is no longer valid (i.e., ). For this reason, the EoM of a Timoshenko rod considers the angle as an independent variable. Its dynamics can be written as
The standard formulation assumes a linear constitutive law with bending moment and shear force defined as
where is the Timoshenko shear coefficient, which depends on the cross-section geometry. We can see that the shear force contribution is proportional to the difference between the tangent angle and the angle of cross-section .
3 Modeling Actuation
The distinction between physical and robotic models lies in the inclusion of actuation. In soft robotics, actuation sources greatly differ from those used in rigid robots. In particular, the most used technologies are cables and fluidic chambers, which guarantee distributed active loads .
3.1 The Actuation Matrix
Consider a Cosserat rod with actuators. Let be the distance from the center of the cross-section and the -th actuator (Fig. 3). Each actuator applies an internal active wrench that depends on the actuator routing. We can express the internal active wrench applied by the actuator as
where is the tangent versor to the actuator path and is the magnitude of the i-th actuator. For example , can be
where is the cable tension, is the pneumatic pressure and is the internal cross-section area of the fluidic chamber. Moreover, the expression of the tangent versor is derived as
where is expressed in homogeneous coordinates and the operator extracts the first three rows of a homogeneous vector. The resultant of the contributions of actuators is
where in the internal active wrench and is the actuation matrix
The actuation matrix is crucial for designing control algorithms because it contains valuable information about the strain modes that can be excited by the actuators. From a control perspective, this information is related to the reachability of the system.
3.2 Actuators-Strain Mapping
From the definition of actuation matrix (35), it is possible to map the actuation magnitude vector to the distributed active load . Renda et al. (2020) propose a statics-based method to relate the actuators with excited strain modes. In particular, let be the statics of a Cosserat Rod as
assuming no external forces and the strain twist is discretized with the Strain Parameterization approach, i.e. (11). By invoking the D’Alembert Principle, it is possible to derive
Finally, substituting the constitutive relation (6) and the definition of actuation forces (34), we derive
From this equality, Renda et al. (2024) propose a trivial form of the static equation, choosing the implicit parametrization
Consequentially, the static equation degenerates in a trivial form, such as
The functional basis derived from the implicit parametrization (39) uses the information of actuation routing contained in , to provide the minimum set of functional basis. In particular, these functional bases describe which are the excited strain modes and which is the best function to describe them statically. Furthermore, in (35) considers also the geometrical and the material information of the robot, contained in . Finally, (39) can be a useful starting point to build the optimal functional basis matrix , avoiding redundant or inefficient shape functions, preserving accuracy and computational efficiency.
4 Rod Models for Continuum and Soft Robots
The rod models were classified based on combinations of the four deformation modes. Out of the fifteen possible combinations, only nine classes emerged. Fig. 4 shows the distribution of rod models over deformations classes. Within each deformation class, we divided the papers mainly by actuation, robot type, or topic. In each division, the articles follow a chronological or logical order. An advanced Scopus query aided the selection of the articles (see Appendix A). Fig. 5 shows how different robots might exhibit these deformations induced by actuators or the interaction with the environment. We survey rod models ordering the deformation classes from most to least popular (see Fig. 4).
We summarize the four major deformation classes (i.e., Bend, All Deformation, Bend & Twist, Bend & Stretch) respectively in Table 1-4 to facilitate the literature comprehension. Each table arbitrarily reports representative articles to distill the deformation class and highlight the versatility of rod models.
4.1 Bend
Bending is a prevalent deformation thanks to the slender morphology of continuum and soft robots (Fig. 5). Herein, we review the evolution of rod models undergoing bending through various mechanisms. Table 1 distills the deformation class reporting representative works.
4.1.1 Bend by Cables
The literature presents several solutions to predict the bending of soft robots. The CRT computed in real-time the kinematics of tendon-actuated robots subject to external loads , modeling the tendon actuation as a single point moment applied to the backbone where each tendon is attached, achieving position errors below on a Nitinol rod (Jones et al., 2009). Direct and inverse kinematic models were presented drawing inspiration from the structure of octopus arms (Giorelli et al., 2012). The direct kinematics exploited a Cosserat rod model. Conversely, the inverse kinematics computes the Jacobian matrix iteratively from the task space. The kinematic models overcome the slack problem as they work directly on the cable tensions instead of controlling the cable extension. Actuations validated the model on a prototype immersed in water. The challenge of predicting the bending angle of a planar manipulator driven by four wires and with triangular notches was addressed using a kinematic Timoshenko rod model (Wenlong et al., 2013). A two-node Timoshenko rod element modeled each triangular unit, accurately predicting the bending angles measured on the tensioned manipulator. Concerning planar grasping, it was proposed a dynamic Cosserat rod model for a tendon-driven continuum robotic finger (Dehghani and Moosavian, 2014b). The finger grasping is simulated as an external point force acting on the finger backbone , assuming that the objects do not slip. Also, a singularity-free version of the model equations is derived using the Taylor expansion. A preliminary experimental validation is conducted by attaching different weights on the tip of a superelastic NiTinol rod. Also, the EBRT was derived using the Lagrangian method to predict the shape of a multi-backbone continuum robot (He et al., 2013). The experimental validation obtained position errors lower than 4%L. Moreover, Gao et al. (2016) predicted manipulator deformations due to external forces, tendon forces, or friction. They provide a planar Cosserat rod model of dexterous continuum manipulators, achieving sub-millimeter tip error predictions in the experimental validation. A dynamic model of a cable-driven soft tentacle was developed to explore the relationship between the strain field and the rope length by formulating a rope-driven actuator model (Ma et al., 2021). A kinematic model based on CRT was presented. Then, a Newton-Euler inverse dynamics algorithm for an equivalent CSM computed the matrices of the Lagrangian inverse dynamics model. A computationally efficient analytical tip force estimator was proposed for tendon-driven catheters (Hooshiar et al., 2021). First, they used a Bezier-based shape interpolation assuming constant curvature. Then, they proposed an optimal problem to invert the system statics and compute the tip force . The experimental validation achieved a N mean absolute error at high frequency. Concerning the interaction with objects, a general CRT-based framework investigated the coupled dynamics of tendon-driven cooperative continuum robots interacting with flexible objects (Jalali and Janabi-Sharifi, 2021). The framework is validated experimentally using two cable-driven spring steel arms performing trajectories with a Nitinol object.
Workspace investigations are crucial for dexterous manipulation tasks. A study on estimating the workspace of CSM was conducted by developing a quasi-static Cosserat rod model and introducing a continuation algorithm for a tendon-driven CSM (Amehri et al., 2021). The algorithm maps the exterior and interior workspace boundaries, addressing challenges like actuator saturation, physical limitations, self-collision, or elastic instabilities. The study highlights that complex configurations of CSM can lead to isolated boundaries and voids. Validation through numerical simulations on planar and spatial configurations demonstrated the effectiveness of the proposal.
Other works focus on the stiffness analysis. A Cosserat rod model for a continuum manipulator with variable stiffness was used to compute the tip stiffness (Zhao et al., 2020a). Experimental results demonstrate the ability to adjust the tip stiffness in various directions with an increase of about ten times the minimum stiffness. Also, a novel cable-driven continuum robot with layer jamming sheaths, which control the stiffness in the transverse and axial directions, is designed, modeled, and validated (Fan et al., 2022). An Euler-Bernoulli rod and the Maxwell-Mohr method derived the effective robot stiffnesses, expressing the rod deflection as
where is the moment due to the unit load. Extensive static force-deflection tests are conducted with different layer jamming sheaths, vacuum pressures, and bending angles. A novel approach proposed additional cables to selectively stiffen planar cable-driven robots with limited coupling with the actuating cables (Molaei et al., 2022). This approach increases the control of shape and stiffness of continuum manipulators without significantly affecting the basic kinematics and addressing uncontrolled deformations like buckling. A planar Cosserat rod model based on (Rucker and Webster III, 2011) identified the design conditions to route the additional cables. The simulations show a contained coupling between the length changes of the actuating and stiffening cables. The proposal is validated experimentally on a prototype with two bending cables and two stiffening cables, demonstrating that the secondary tendons increase the stiffness by ten times. Stiffness is also investigated for a simple rod-shaped silicone soft robot modeled with a geometrically exact CRT (Grube and Seifried, 2022). The stiffness properties are investigated by applying horizontal pulling forces up to 3 N to the rod and robot free-ends. They analyzed experimentally a linear viscoelastic material model and the nonlinear Saint Venant-Kirchhoff material model. The analysis of the damping mechanism has shown that linear damping can be assumed for oscillations with small to medium amplitudes since nonlinear material laws improve the model accuracy only slightly. Therefore, nonlinear effects have to be considered for oscillations of higher amplitudes. Xiao et al. (2023) present an on-demand stiffness modulation approach that retains the intrinsic compliance of elastomeric bodies. They use a Nitinol tube as the backbone, which slides through a cable-driven soft body to achieve robot pose or stiffness modulation. They develop and validate a CRT model in different scenarios by varying and the contact force . Experimental results indicate average shape and tip errors below 1%L, while simulations demonstrate that backbone insertion enhances the workspace and reduces compliance by 57.7%. Their design was further validated by object manipulation and minimally invasive surgery (MIS) case studies.
A generalized kinetostatics CRT model with friction considered for multi-module and multi-segment continuum robots is proposed (Wang et al., 2023). The model is applied to a two-module rod-driven continuum robot with winding ropes to analyze its deformation and load characteristics. The experimental validation, with and without loads, achieves less than shape error.
4.1.2 Bend in Soft Pneumatic Actuators
Several works estimate the bending deflection in pneumatic soft robots. An Euler-Bernoulli rod described a bending pneumatic soft actuator (de Payrebrune and O’Reilly, 2016). Through experimental procedures with the robot clamped at one end, the curvature was expressed as a function of the applied pressure . The results show that the curvature field is a function of the applied pressure.
FEM simulations often provide a benchmark. Using FEM and experiments, a five-parameter constitutive relation for a rod-based pneumatic soft arm model was explored (De Payrebrune and O’Reilly, 2017). The constitutive law’s parameters are determined for pneumatic soft arms through FEM simulations under different conditions. Furthermore, various actuator sections are investigated (e.g., rectangular, semicircular). The parameterization is validated by comparing FEM with the rod model for various actuator sections, subject to tip loads . Another work introduced interoperable dynamic modeling for fluidic elastomeric soft fingers (Mbakop et al., 2021b). The model employs an Euler-Bernoulli framework with Pythagorean Hodograph (PH) curves reconstructing the kinematic shape from virtual control points. The shape of the soft finger is recovered as a function of the applied pressure and control points as
Here, is the height from the ventral part of the finger in the cross-section, and is the normalized curvilinear coordinate along the neutral axis of the soft finger. In this way, a relationship is established between control inputs and PH curve control points, enabling dynamic inversion. FEM simulations validated the PH-EBRT approach for single and two-segment soft fingers, while the experimental validation measures cartesian error and bending curvature on 3D-printed soft fingers of varying stiffness. Interestingly, EBRT successfully modeled a compliant robotic fish tail with fluidic actuators (Wu and Shi, 2022b). The model predicts actuator deflections and nonlinear strains. Nonlinear FEM validated the approach by comparing the results in terms of the first, second, and third obtained natural frequencies, showing good agreement. Finally, the bending of a beam-shaped soft actuator with an embedded fluidic network was analyzed by leveraging a dynamic Euler-Bernoulli rod with large deflections and nonlinear strains (Janizadeh Haji and Bamdad, 2023). Its accuracy is compared against nonlinear FEM and 4th Runge–Kutta. Examining multi-mode and nonlinear strain effects revealed instabilities in some frequencies for the nonlinear model. Geometric parameters influence actuator deformations, with increased channel distribution density enhancing deflection. Moreover, larger fluid channel diameter and actuator length require fewer inlet pressures for the same bending, while higher actuator width and height demand higher pressures. Conversely, Sadati et al. (2019) introduced a ROM and a discretized model based on EBRT with absolute states for the STIFF-FLOP. In addition, they improve the model in (Renda et al., 2018a) with an Euler-Bernoulli discretized model with relative states. An experimental comparison revealed that the ROM is more accurate and numerically robust, while absolute states provide higher computational efficiency. In contrast, relative states show higher sensitivity and computational burden. Finally, a lumped-mass model provided a benchmark with the highest computational cost among the considered models.
Bending structures also exhibit stiffness modulation and adaptation. For instance, CRT modeled a soft bending actuator with an added fabric layer, which changes the actuator stiffness, enhancing the bending and preventing longitudinal stretching (Nalkenani et al., 2021). The authors incorporate fabric effects as a boundary condition in the Cosserat model and validate experimentally across actuation frequencies and loading conditions, producing accurate behavior predictions. Also, it was proposed a nonlinear kinematic Cosserat rod model of a pneumatic surgical CSM with stiffness adaptation (Roshanfar et al., 2022). The model incorporates nonlinear material properties through the two-term Mooney-Rivlin hyperelastic constitutive law, whose strain energy density is
Here, and are the deviatoric strains invariants while and are material constants. The simulation and experimental results for tip displacement and stiffness show similar trends with a maximum error of 8.25%.
More recently, the work in (Namdar Ghalati et al., 2022) presented a method to obtain the static shape of a semicylindrical soft fiber-reinforced pneumatic bending actuator under external force constraints. The pressure-angle relationship in free motion is derived using the Neo-Hookean strain energy theory of hyper-elastic materials, i.e.,
Here, is the first invariant of the right Cauchy-Green deformation tensor and , , are the principal stretches (Odgen et al., 2013). They investigated the effects of external forces using the EBRT, which modeled the arm as a cantilever beam under external load.
4.1.3 Bending with Smart Actuators
Bending is achieved by several innovative actuation technologies. Inspired by the forward crawling motion of the Manduca sexta caterpillar, a dynamic Euler-Bernoulli rod model for a soft robot with shape memory alloy (SMA) actuators was introduced to analyze undulation mechanics (Daily-Diamond et al., 2017). The model employs a motion pattern based on the Witch of Agnesi function
Its shape is analogous to the probability density function of the Cauchy distribution. Lastly, experimental data on an SMA actuator’s flexural rigidity, curvature, and recorded caterpillar leg velocity enable studying ground contact forces and the energetics of successful forward locomotion. Differently, CRT solved direct and inverse kinematics of soft robots actuated by twisted and coiled polymers (Pawlowski and Zhao, 2018). Direct kinematics determines the manipulator shape based on TCP temperatures, and inverse kinematics uses an augmented Jacobian approach (Rucker and Webster, 2011). A subsequent work integrates a thermodynamic TCP model and a dynamic Cosserat rod model, with promising forward dynamics results (Pawlowski et al., 2018). Additionally, an Euler-Bernoulli rod model predicted the deflection of a micro soft robot made from a magnetic elastic composite material (Xiang et al., 2019). The 3D-printed robot, having a head and a tail for rolling and crawling, is controlled by two magnetic fields generated by a 3D Helmholtz coil. Experimental validation for rolling and crawling velocity under different input waveforms, amplitudes, and frequencies shows a linear relation of rolling speed to rotating magnetic field frequency. Moreover, crawling speed correlates positively with magnetic field amplitude and frequency, being faster with sinusoidal compared to triangular wave inputs. A recent work developed a dynamic model of a continuum robotic arm with a 3D-printed elastic backbone and orthogonal plates supporting distributed soft cubic actuators, composed of a novel temperature-responsive hydrogel, each with an embedded Joule heater (Doroudchi et al., 2020). Leveraging the CRT, they simplify the PDE into reduced ordinary differential equations by integrating semi-discretized distributed force and rotational inertia components. Assuming robot isotropy and homogeneity, they estimate the Young modulus and damping coefficients through vibrational tests. Experimental verification accurately predicts tip displacements with a maximum of error. Finally, a study proposes a model of CSM with magnetic actuation (Tariverdi et al., 2020). They use Lie group synchronous variational integrators to derive a geometrically exact model based on an extended dynamic Kirchhoff rod and Hamiltonian formulation. The model considers nonlinearities, magnetic field loads, gravity, and internal and external dissipation forces from friction. Experiments involve dynamic motions with metallic and polymeric soft rods. Polymer rods show higher errors due to fabrication uncertainties and nonlinear elastic properties. Increasing the number of nodes improves the system frequency accuracy.
4.1.4 Bending in Parallel Manipulators
A novel parallel continuum manipulator (PCM) constructed with soft panels and its shape modeling approach was presented (Li et al., 2018). They used the nonlinear EBRT to model the deformations of each soft panel. By solving a constrained optimization for the kinematics, they calculated the deformed shapes and obtained the relationship between the panels’ length and the pose of the moving platform. Experiments validated the model reaching good shape predictions.
Li et al. (2023) present a kinematic model of a planar PCM made of two independent flexible panels. The formulation unifies the geometric constraints, the load balance of the moving platform, and the panel’s statics. For each panel, the relation between internal moment and deflection is described by the EBRT via elliptic integrals, which account for inflection points with large deformations. Simulation results compared against FEM reveal that, under large deflections, the panels exhibit two inflection points close to both ends, which affect the distribution of the internal moment along the panel, (25). Stress checking confirms that the deformations are within the elastic region.
4.1.5 Bending in Tubular Robots
Bending is the prominent deformation of tubular robots aimed for MIS. Pattanshetti and Ryu (2019) employ the CRT to model a laser-patterned monolithic tube with varying structural properties along the length. The combination of CRT and a vision-based identification algorithm enables the characterization of the non-linearities and the bending stiffness along the rod with increasing strain. Another work adjusts the piecewise variable-strain (Renda et al., 2020; Boyer et al., 2020) to simulate multi-section concentric tube robots by including the tubes sliding motion (Renda et al., 2021). Rotation motions of the tubes are included as generalized coordinates instead of boundary kinematic conditions. This procedure led to a minimum set of closed-form algebraic equations solvable for both the shape variables, the actuation forces, and the torques. The resulting equations facilitate control, design optimization, and stability assessment. The work tackles one of the main limitations of the sliding structure of standard CTR, the spaghetti problem.
4.1.6 Bending in Simulated Locomotion
Bending is also exploited to investigate in simulation locomotion strategies (Fig. 5). Zhou et al. (2015b) describe a soft robotic limb attached to a mass with an Euler-Bernoulli rod to examine in quasi-static conditions the peristaltic locomotion controlling the intrinsic curvature on a rough horizontal surface. The paper also provides design recommendations for the optimal performance of soft robots with variable curvature and adhesion. Interestingly, they also studied grasping strategies using a gripper with two fingers attached to a palm modeled using two rods (Zhou et al., 2015a). Concerning snake-like locomotion, especially undulatory swimming, is examined by studying a planar inextensible elastic Cosserat rod with adjustable spontaneous curvature, providing an internal actuation mechanism resembling muscular action in a snake (Cicconofri and DeSimone, 2015). Equations of motion are derived for (a) the rod moving along a prescribed curve and (b) the rod sliding longitudinally without fixed lateral slipping. Non-standard boundary conditions arising from constraints enable solving the equations. The one-dimensional equation captures system kinematics and dynamics without restrictive assumptions on trajectory or actuation, emphasizing the role of spontaneous curvature in driving and steering locomotion.
4.1.7 Bend Miscellaneous
Rods might exhibit interesting and mutable behaviors. Armanini et al. (2017) studied a soft robotic arm clamped at its base with a revolute joint, discovering that when the load is smaller than the buckling critical load, the rod behaves as an elastica compass. The closed curve can be computed analytically by the EBRT. Conversely, when the load exceeds the critical buckling value, the rod behaves like an elastica catapult, causing the snap phenomenon, where the rod releases very quickly elastic potential energy. They computed the dynamic motion using a simplified massless rod model and FEM and validated the behaviors by experimenting with an elastic rod.
Instead, Cacace et al. (2020) present a model for a simulated octopus tentacle based on the dynamics of an inextensible string with curvature constraints and curvature controls. They derive the EoM and appropriate boundary conditions characterizing the corresponding equilibria. The model gives fourth-order evolutive nonlinear controlled PDE, generalizing the classic Euler’s dynamic Elastica equation. They explore the reachability of the model and introduce and compare two optimal control algorithms based on the augmented Lagrangian method solved with adjoint-based gradient descent.
Subclass | References | Rod Model | Contribution |
Cable | Jones et al. (2009) | CRT | Static model for real-time shape estimation of CSM. |
Dehghani and Moosavian (2014b) | CRT | Dynamic model of a continuum finger for planar grasping. | |
Gao et al. (2016) | CRT | Model of notched continuum manipulator for MIS. | |
Pneumatic | De Payrebrune and O’Reilly (2017) | EBRT | Derivation of constitutive law of a planar bending actuator. |
Sadati et al. (2019) | EBRT | ROM and rod models with absolute and relative states validated on STIFF-FLOP. | |
Smart | Daily-Diamond et al. (2017) | EBRT | Dynamic model inspired by caterpillar motion to analyse ondulation mechanics. |
Pawlowski and Zhao (2018) | CRT | Modeling the coupling between CSM body and thermal actuation. | |
Parallel | Li et al. (2018) | EBRT | Shape modeling of a PCM made of soft panels. |
Tubular | Pattanshetti and Ryu (2019) | CRT | Estimation of stiffness variation on a monolithic tube for MIS. |
4.2 All Deformations
The second most popular deformation class concerns models accounting for all the primary deformations of continuum and soft robots. Below, we review the literature on rod models exhibiting all strains in cable-driven, pneumatic, parallel continuum, and biomechanical structures. Table 2 summarizes the class with the notable works.
4.2.1 Deformations by Cables
Renda et al. (2012) present a geometrically exact static Cosserat model of a conical tendon-driven CSM inspired by the octopus. Cable tension is modeled as a pure tension element, neglecting friction between the cable and the silicone body. After a FEM validation of the algebraic equations, the Cosserat model is validated experimentally for a static planar motion. The following work investigated the dynamic interaction of the silicone arm with a dense medium and the coupled tendon condition, addressing the external forces such as gravity, buoyancy, drag, added mass, and cable load (Renda et al., 2014). The dynamic experimental validation concerned the replication of characteristic octopus movements. In addition, an ablation study suggests that minimal internal viscosity guarantees solution stability and minimizes computational costs. Then, Renda et al. (2018b) introduced a unified dynamic Cosserat rod framework for underwater soft robots, which considers the nonlinearities of soft bodies, offering generality that extends to shells. The framework, which could aid the evaluation of potential soft robot designs, was validated experimentally using an autonomous vehicle with four cable-driven octopus-inspired CSM. At the same time, Renda et al. (2018a) presented an approach called piecewise constant strain (PCS) that discretizes continuous Cosserat rods in pieces, assuming constant strain for each piece. The PCS generalizes the PCC (Webster III and Jones, 2010) providing (i) numerical robustness given by the exponential formulation, (ii) inclusion of shear and torsion due to the CRT, and (iii) possible extension to rigid kinematic chains. The PCS is validated in real-world experiments with a cable-driven siliconic CSM.
Concerning discretization approaches, it was proposed a real-time discretization technique to solve the PDE of a Cosserat rod (Till et al., 2019). The method discretizes the time derivatives and solves the ODEs for the boundary value problem (BVP) in at each timestep. The authors validate the efficient method by comparing the simulation to high-speed footage of a cantilever rod.
For MIS applications, (Ibister et al., 2021) presents a dynamic CRT model for a continuum robot composed of a backbone and disks to support actuating tendons. The model includes the gravity (7) and tip wrenches produced by the tendons. The comparison between simulation and experiments shows a mean tip position error.
Recently, a cable-driven continuum robotic arm was integrated within an aerial manipulation system (Samadikhoshkho et al., 2020). The arm dynamics are modeled with a Cosserat rod with a floating base , incorporating both the actuation and quadrotor effects. They proposed a robust adaptive position controller and proved its stability using the direct Lyapunov stability theorem. The scheme is validated in simulation for quasi-static, passive, and active arm conditions.
4.2.2 Deformations by Pneumatic Actuators
Trivedi et al. (2007, 2008b) present the first dynamic Cosserat rod model for pneumatic CSM based on a fiber-reinforced model of the air muscle actuators. The model includes the effect of material nonlinearities, distributed forces, and payloads and is geometrically exact for large curvature, shear, torsion, and extension. The model is experimentally validated statically on the OctArm V for planar movements with vertical and horizontal base orientation, achieving results about ten times more accurate than the constant curvature model. Subsequently, the model is used within a model-based design optimization study for pneumatic and hydraulic CSM (Trivedi et al., 2008a). Recently, the PCS discretization was combined with Screw Theory, providing a new model of actuation sources for pneumatic actuators (Renda et al., 2017). This approach enables the routing of fluidic actuators (even tendons) along any path in the soft arm. They define a soft geometric Jacobian for the PCS formulation and a technique to identify the active DoFs, applying it to the study of manipulable and force-closure conditions. Finally, they validate the theory on a STIFF-FLOP arm (Cianchetti et al., 2013). Additionally, Roshanfar et al. (2021) described a pneumatic soft robot with a Cosserat rod model, which was solved as an initial value problem with homogeneous Neumann and Dirichlet boundary conditions for a given tip force. The experimental validation measured a tip position error of .
Analytical solutions and detailed FEM simulations are useful means for comparison and validation. For example, analytic solutions for the statics of a cantilever quasi-static Cosserat rod under external loads are derived (Grazioso et al., 2018). The position and orientation fields of the arm are defined on a Lie group, while the velocity and deformation fields are obtained from the time and space derivatives, respectively. A constitutive law between internal forces and strains enables the definition of the arm’s internal strain energy. Therefore, the principle of virtual work provides the static equilibrium equations. Finally, quasi-static rod configurations are compared to analytical solutions for pure bending and shear loads. Conversely, a comparative study for a Cosserat model and a FEM model of a soft pneumatic actuator was presented (Berthold et al., 2021), providing a preliminary parameter identification process. The FEM model was also validated experimentally. Moreover, an operational point-dependent static Cosserat model is obtained from a detailed volume FEM of a fiber-reinforced pneumatic silicone soft robot (Wiese et al., 2022). FEM simulations determine the actuation-dependent curvature, while the rod stiffness properties are identified by applying external forces to the FEM. The validation against FEM and experimental data, under various conditions with and without loads, show that characterizing a single module is sufficient to predict the behavior of a two-module version of the soft robot.
To tackle large deformations of pneumatic soft robots, Mishra et al. (2022) suggest using an explicit fractional-order Bouc-Wen model that considers material nonlinearities, such as hysteresis. The constitutive law incorporated in a Cosserat rod and tested on a Robotino-XT demonstrated superior accuracy to other models (Wang et al., 2022a).
4.2.3 Deformations with Parallel Manipulators
Till and Rucker (2017) addresses the problem of elastic stability in PCM, introducing a condition to evaluate the elastic stability of PCM solutions based on Kirchhoff and Cosserat rods. A numerical test evaluates this condition and a heuristic stability metric. After verifying that the stability test matches other classical results, they experimentally validate the approach on a six-link PCM. The same authors propose a generalized kinetostatics Cosserat rod model (Black et al., 2017). Model linearization provides the manipulator Jacobian, end-effector compliance, input stiffness, and wrench reflectivity matrices. This enabled the analysis of the PCM manipulability and force-sensing accuracy. Using a similar PCM, they achieve accurate actuation-based force sensing and low orientation error.
Conversely, Armanini et al. (2021) extended the PCS parametrization to soft PCM. They present a mathematical tool for modeling closed-chain geometries composed of soft and rigid elements. The formulation is applied to modeling and analyzing the Fin-Ray finger, especially for studying bending in the opposite direction of the external force. Furthermore, this formulation allowed the development of a novel design of the Fin-Ray finger, finding the optimal number of rigid ribs and their optimal orientation in terms of maximum payload and bending capability.
The elegant work of Lilge and Burgner-Kahrs (2022) present a CRT-based kinetostatic modeling framework for tendon-driven PCM, which enabled to efficiently solve the forward, inverse, and velocity kinetostatic problems. In addition, it facilitates the study of kinematic properties like reachability, singularities, manipulability, and compliance. Experiments on a PCM made of four tendon-driven backbones attached via spherical joints to a platform achieved median pose accuracy of and .
4.2.4 Deformations in Bioinspired Modeling
Cosserat rods can be useful tools for biomechanical studies. With application to flapping flight and passive swimming, Boyer et al. (2017) provide a general unified framework for modeling bio-inspired locomotion robots using soft organs. They model soft parts in two ways: (i) the floating frame approach and (ii) the geometrically exact approach. They implemented the two models in a unified general algorithm based on recursive Newton-Euler. The comparison of the two approaches revealed that the former needs fewer coordinates and stronger deformation approximations. Conversely, the latter is more accurate and manages better the external load. The validation of the unified framework simulated two bio-inspired examples of soft locomotion.
Recently, Zhang et al. (2019) realistically simulates musculoskeletal architectures in PyElastica. They assembled heterogeneous, active, and passive Cosserat rods into dynamic structures that mimic bones, tendons, ligaments, fibers, and muscle connectivity. The simulations involve biological and soft robotic case studies, including the human elbow joint, the design and fabrication of swimming and walking millimeter-long bio-hybrid bots, the synthesis of slithering snakes with intricate musculature layouts, and the replication of feathered bird wings - using up to rods. In these applications, the CRT demonstrated versatility to advance forward and inverse bioengineering designs and fundamental discoveries in the functioning of living organisms.
Boyer et al. (2022) investigates the relationship between CRT and optimal control theory, deriving the statics and dynamics of a continuum robot as solutions to a minimization problem. With this different formulation, the authors revealed significant theoretical insights, such as the intrinsic singularity of the dynamic simulation of continuum robots. Furthermore, the iterative solution of the optimal control problem corresponds to the continuous version of the Newton-Euler algorithm for rigid kinematic chains. Finally, the presented approach is extended to locomotors and applied to a bio-inspired continuum swimmer, validating the theoretical work through simulations.
Subclass | References | Rod Model | Contribution |
Cable | Renda et al. (2014) | CRT | Studied the dynamic interaction of an octopus-inspired CSM with a dense medium. |
Pneumatic | Trivedi et al. (2008b) | CRT | Model validation of CSM with fiber-reinforced actuators varying base orientation. |
Parallel | Till and Rucker (2017) | CRT | Studied and experimentally validated the elastic stability of PCM. |
Armanini et al. (2021) | CRT | Modeled closed-chain soft PCM with validation on a soft-rigid finger. | |
Lilge and Burgner-Kahrs (2022) | CRT | Studied reachability, singularities, manipulability, and compliance of PCM. | |
Bioinspired | Zhang et al. (2019) | CRT | Simulates musculoskeletal systems assembling heterogeneous active/passive rods. |
Boyer et al. (2022) | CRT | Studied the link between CRT and optimal control simulating locomotors and swimmers. |
4.3 Bend & Twist
Like bending, various robot types exhibit the combination of bending and twisting induced by diverse actuation types (Fig. 5). Table 3 summarizes the Bend & Twist deformation class.
4.3.1 Bend & Twist by Cables
Bending and twisting are usually induced by cables and tendons through specific routing paths (Fig. 5). Rucker and Webster III (2011) present Cosserat rod models for the forward kinematics, statics, and dynamics of a tendon-driven continuum manipulator. The curved tendon paths deform the body in various shapes. Starting from the classical CRT, they derive the expression of distributed loads applied by tendons on the backbone . The model validation considers straight and non-straight tendon routings, point loads, and distributed external loads . The simulated matched experimental tip positions with a 1.7%L average error. Another experimental work presented a continuum arm with a Nitinol backbone and three tendons turning around the centerline with a constant radial distance (Dehghani and Moosavian, 2013), which extended a previous setup lacking an actuation system (Dehghani and Moosavian, 2011). They modified differential terms of an existing CRT (Rucker and Webster III, 2011), retaining the original accuracy with reduced numerical effort. The model was extended to real-time control using a stable moment-based algorithm to solve the BVP considering the physical characteristics of the rod (Dehghani and Moosavian, 2014a). They also propose a method to reduce the cumulative numerical integration errors that could make the rotation matrices non-orthogonal. Finally, they use the model to develop a Jacobian-based quasi-static tip position controller for a multi-segment continuum arm.
Cosserat rods can simulate the mechanics of several robotic systems, including robotic tails and octopus-inspired arms. The work in (Rone and Ben-Tzvi, 2014) used a Cosserat rod to study a cable-driven multi-section continuum robotic tail envisioned for onboard mobile robots. Prescribed tail configurations or trajectories are analyzed, while the governing equations compute the loading at the tail base transmitted to the mobile robot. This study investigates the impact of trajectory and design factors on the resulting loading profiles, considering factors like mode shape, speed, bending magnitude, and bending plane angle. Key findings include that a non-straight tail reference configuration increases the moment magnitude, short proximal segments provide a larger tip motion for a given bending angle, and higher mass distribution towards the final nodes enhances bending moments and control fidelity. Xu et al. (2017) propose a Cosserat rod model to describe the bending and torsion behavior of a bionic tendon-driven octopus tentacle. Given the location of the four cables, the model maps cable tension to expected bending. Exploiting the derived mapping, they propose an open-loop controller for the robot. Similarly, Niu et al. (2019) derive closed-form equations of a soft robot modeled through the CRT. The equations are validated with a cable-driven soft arm inspired by an octopus tentacle, reaching accurate predictions of the end-effector positions. Chikhaoui et al. (2019) investigated the kinematic modeling of a magnetically extensible tendon actuated continuum robot (METAbot) (Nguyen and Burgner-Kahrs, 2015). The comparison between a PCC beam mechanics approach and a Cosserat rod shows that the latter has higher position accuracy, although computationally heavier. Recently, Janabi-Sharifi et al. (2021) presented a tutorial on the dynamic Cosserat rod modeling of tendon-driven continuum robots. The tutorial explains the core assemblies of tendon-driven continuum robots. Then, it provides a unified approach for the dynamic analysis of tendon-driven continuum robots that bend and twist, including multi-section, concentric tubes, and parallel cooperative robots.
4.3.2 Bend & Twist in Soft Pneumatic Actuators
Actuation-induced torsion is not common in fluidic actuators. However, a Cosserat rod model for a CSM with two asymmetric pneumatic actuators that bend and rotate was introduced (Uppalapati et al., 2018). They optimize the rigidity parameters and the curvatures for quasi-static shapes. The model reproduced the experimental static workspace with a tip load attached. The robot’s dexterity was enhanced integrating an additional soft rotating actuator (Uppalapati and krishnan, 2021). The CRT predicts the deformed robot configuration subjected to external loads, and they experimentally determine elasticity and actuation parameters.
Another study introduced a multi-segment Euler-Bernoulli rod model to predict the bending deformation of micro-tube pneumatic actuators (Ji et al., 2019). They first employ the line-segment model and then extend it to a multi-segment version of Euler-Bernoulli’s rod model for a rapid design, analysis, and optimization of the bending deformation of such actuators. Finally, an analytical model of a generalized PneuNet soft actuator combines the minimum potential energy method and an Euler-Bernoulli rod (Gu et al., 2021). The study compared the rod model to FEM simulations and experimental results, achieving accurate deformed shape predictions, given the geometry and the applied pressures.
4.3.3 Bend & Twist in Tubular Robots
Biomedical applications could leverage CTR and eccentric tube robots that bend and twist. A geometrically exact KRT solved the statics problem for CTR, describing the shape of a collection of concentric pre-curved tubes under general wrench distributions (Rucker et al., 2010a). The model is validated experimentally for an active cannula with three concentric, telescopic, pre-curved Nitinol tubes in several configurations with varying tip loads. The results establish the basis for model-based position and force control of active cannulas and their use as force sensors under external loads. Also, they tackle the integration of unmodelled effects such as friction, elongation, and transverse shear (Rucker et al., 2010b). Subsequently, Till et al. (2020) present the first dynamic model for CTR. The model describes the coupled inertial dynamics of pre-curved concentric tubes considering the dynamic effects of linear and rotational tube inertia, material damping, Coulomb and viscous friction, and the inertia of a rigid body at the robot tip. The experimental validation measures the dynamic behavior during tissue grasping and snapping.
The structure of the ETR extends the CTR design. A kinematic model of a two-channel ETR with two continuum arms delivered through the channels is presented (Wang et al., 2019). Pre-curved superelastic tubes lining the arm’s working channels produce sheath deflection. Cosseat rods model the pre-curved tubes and the central backbone, achieving promising tip pose predictions for pure rotation actuation. Similarly, a continuum ETR sheath for MIS such as neuroendoscopy, requiring delivery of multiple robotic arms, was presented (Wang et al., 2021a). Super-elastic tubes lining the arm channels actuate the sheath, while distributed disks are attached to the backbone. The tube kinematics and the robot backbone are modeled using the CRT. Three experiments with a two-arm sheath validate the tip position accuracy. Another work proposes a kinematic Cosserat rod model of a snake-like continuum robot for MIS (Wang et al., 2020). The robot geometry comprises three identical super-elastic Nitinol tubes held eccentrically by a surrounding sheath. Assuming sheat inextensibility, they show that different actuations virtually yield the same robot shape and tip pose. Simulations and experiments validate the model for a single-stage ETR.
Safety is crucial in surgical procedures. A method to compute compliance of continuum robots, modeled as quasi-static Cosserat rods under diverse loads, was devised (Smoljkic et al., 2014). The robot compliance is described by ODEs and integrated with the forward kinematics. The experiments involve a non-actuated flexible rod composed of a super-elastic Nitinol tube surrounding a braided Pebax tube. The rod mounted on a force sensor with position sensors embedded in the tip is subject to external tip forces. A least square fitting of the measured tip trajectory versus the direct kinematics identifies material parameters and tube configuration. Finally, they compared the velocities obtained through position differentiation against those obtained multiplying the differentiated forces and the computed compliance matrices.
4.3.4 Bend & Twist with Smart Actuators
The KRT modeled TCP emphasizing critical factors like coil load and twist and incorporating them into the actuation equations for TCP muscles (Wu and Zheng, 2020). The model captures the quantitative relationship between tensile actuation and fabrication load, serving as a predictor for tensile stroke. Also, it was validated through TCP muscles made of silver-coated Nylon fiber under different fabrication loads, measuring recovery stress and maximum stroke during heating. Thermo-mechanical tension tests conducted at various room temperatures identified the elastic modulus of the fibers. To achieve variable stiffness in continuum robots, Wang et al. (2022b) propose to use winding-styled ropes inspired by snakes’ behavior. Controlling the temperature of an SMA spring increased their stiffness up to . The experimental validation of a kinetic-static Cosserat rod model obtained an error below .
4.3.5 Bend & Twist in Parallel Manipulators
4.3.6 Bend & Twist with Buckling Instability
Buckling instability is an important factor to consider in soft robots. With applications in soft robotics, Sipos and Várkonyi (2020) explores the interplay between spatial buckling and arbitrary intrinsic curvature on a simulated clamped Cosserat rod with a vertical tip load. The aim is to tune the control parameters to find the highest load for a given free-end position or the maximum reaching of the cantilever in a direction for a given load. The numerical investigation of the spatial instability of in-plane configurations of soft elastic cantilevers subject to an endpoint load demonstrates the importance of out-of-plane instability. Moreover, the capability of controlled intrinsic curvatures to stabilize the rod in configurations of high reaching is shown. Significant improvements occur for flattish cross-sections where the rigidities associated with out-of-plane bending and torsion are larger than the in-plane stiffness. At the same time, arms with high and narrow cross-sections are slightly more performant. Buckling instability can be triggered. Forghani et al. (2021) analyze the inverse dynamics and control of a bacteria-inspired uniflagellar robot in a fluid medium at a low Reynolds number. The robot, mimicking the locomotion of flagellated bacteria, features a flexible helical filament (flagellum) connected to a spherical head. A Kirchhoff elastic rod describes the flagellum, which rotates about the head at a controlled angular velocity and generates propulsive forces for the forward motion. The results demonstrate that by triggering the buckling instability of the flagellum, the robot follows a prescribed path in space. The control scheme only involves the flagellum angular velocity, and a deep neural network captures the dynamics predicting the robot’s trajectory. However, higher maneuverability tasks like drug delivery require multiple flagella, advanced controllers, and control inputs.
4.3.7 Bend & Twist in Insect-like Wings
The applicability of rod theories goes beyond manipulators. A study presented a hybrid dynamic Cosserat rod model of a 3D aerial insect-like robot with two soft wings (Belkhiri et al., 2012). In particular, Cosserat rods described the leading edge of the wings, which bend and twist based on continuous Newton-Euler dynamics. The wings, connected to the insect thorax through revolute joints, are treated as rigid bodies. The hybrid model combines discrete and continuous models considering the curvature deformations and external aerodynamic forces. The model replicates the kinematics and internal deformations of the wings and the rigid net insect motion produced by a flapping flight.
Subclass | References | Rod Model | Contribution |
Cable | Rucker and Webster III (2011) | CRT | General dynamic model of a continuum manipulator with curvilinear actuators. |
Pneumatic | Uppalapati and krishnan (2021) | CRT | Design and modeling of a CSM with parallel asymmetric fiber-reinforced elastomers. |
Tubular | Till et al. (2020) | CRT | Dynamic model of CTR validated for tissue grasping and snapping. |
Wang et al. (2021a) | CRT | Modeling of a superelastic ETR sheath for biomedical procedures. | |
Smart | Wu and Zheng (2020) | KRT | TCP muscles incorporating coil load and twist into actuation model. |
Buckling | Sipos and Várkonyi (2020) | CRT | Study relationship between buckling and on simulated cantilever rods. |
4.4 Bend & Stretch
The simultaneous bending and stretching emerge in many cable-driven and pneumatic CSM (Fig. 5). Table 4 reports illustrative examples of the Bend & Stretch class.
4.4.1 Bend & Stretch by Cables
Deflection and stiffness estimation are crucial for continuum robots actuated by cables. For instance, robot deflection and stiffness were predicted considering environmental loads with constant axial tendon displacements (Oliver-Butler et al., 2019). Preliminary experiments with an Euler–Bernoulli rod evaluate how tendon stretch and routing path affect the robot stiffness at any point. The analytical results enable stiffness analysis of candidate robot designs with light simulations and suggest the ability to increase robot stiffness by using converging tendon paths. Finally, they extend the study to a non-linear Cosserat rod model with large deflections, handling prescribed tendon displacements, tendon stretch, pretension, and slack. Experimental validation on a prototype shows good agreement. Also, a novel stiffness modulation approach using a Cosserat rod model was proposed for a tendon-driven soft CTR (Xiao et al., 2022). The experimental validation shows an accurate shape estimation. They also demonstrate that the robot stiffness changes with a Nitinol backbone. In particular, the tip compliance can be halved, with potential applications in manipulation tasks.
Concerning workspace estimation, the CRT is also applied to investigate equilibrium positions reachable by the end-effector of a planar tendon-driven CSM with a discrete actuation space (Walid et al., 2022). Experiments on a physical prototype efficiently estimated the robot workspace. Another example predicts the soft arm shape combining the EBRT, which considers the influence of external forces in bending, with the PCC model, which efficiently derives a relationship between cable length and arm curvature (Chen et al., 2022). Experiments on a cable-driven soft arm subject to external forces, or , yield a maximum error in the range .
4.4.2 Bend & Stretch in Soft Pneumatic Actuators
A study compares five models, including the Cosserat rod, a dynamic lumped system, and approaches with constant or variable curvature (Sadati et al., 2017b). The comparison made in simulation and on the STIFF-FLOP soft arm (Cianchetti et al., 2013) demonstrated that the lumped mass model better predicts the dynamics, while constant curvature methods are less accurate but are suitable for design optimization. Additionally, they modeled CSM actuated by braided pneumatic actuators (Sadati et al., 2017a). This specific design causes cross-sectional deformation, which was captured analytically through geometry deformation and CRT. This approach outperformed constant curvature approaches by and variable curvature by . A subsequent work presents a semi-analytical model for the dynamics of multi-segment CSM (Sadati et al., 2017c). The Cosserat rod static and Lagrangian dynamic models are derived and combined with the Ritz-Galerking method to reduce the infinite state space of CSM to a small finite number of states.
Moreover, a static Kirchhoff rod model was proposed for a hybrid CSM characterized by pneumatic actuators with embedded elastic rods, computing its deformations using the minimum energy principle (Sun et al., 2019). In addition, the classical CRT modeled a variable-length pneumatic CSM (Gilbert and Godage, 2019). They optimize the cross-sectional area of the actuators and a parameter relating stiffness and damping in the constitutive law. The model is validated experimentally on dynamic bending deformations obtained by pressurizing the actuators with alternating square pulses. A study characterized the bending behavior of a cantilevered extensile fluidic artificial muscle (Garbulinski and Wereley, 2022). An Euler-Bernoulli rod determined static properties like the bending stiffness, which increases more than four times in the pressure interval psi. Experimental data are applied to assess the dynamic properties, such as natural frequency and damping. Also, it was developed an analytical model of a PneuNet actuator based on the Euler-Bernoulli finite strain rod (Sachin et al., 2022). They describe the actuator behavior in free motion and for tip contacts and accurately predict the actuator deformation and the force characteristics. Another work showcases a parallel extensible soft robot with three pneumatic actuators, each with a silicone body, inner air chamber, winding wiring, and sealing head (Wang et al., 2022c). The dynamic model of a single actuator uses the CRT, and an interactive constraint topological structure solves the constraint force and moment of the parallel robot. A linear mapping between the pressure and the extension force is established for each actuator. Validation measures the static step response of a single actuator and conducts dynamic experiments with the complete soft robot. The model accurately captures the actuator dynamics and the parallel robot constraints. Rod models met traditional robotics models where the EBRT augmented the kinematic model of a pneumatic soft bending finger treated as a planar serial rigid robot parameterized with the Denavit-Hartenberg convention (Flores-Martínez et al., 2022). The experimental validation compares the actuator trajectory versus the kinematic model and the flexion-extension motion trajectory of average human fingers.
Manufacturing uncertainty could be significant in fluidic soft actuators. For instance, Eugster et al. (2022) described the kinematics of a soft arm using a CRT with a nonlinear pressure-dependent constitutive law, using the principle of virtual work and modeled the actuator with strain energy functions. Manufacturing imperfections are considered by scaling the relations of extensional and bending stiffness and by a constant shift of the actuator position . Assuming the pressure dilates the chambers, the model also considers cross-section deformations with a pressure-dependent diameter . The model is validated experimentally for static stretching and bending motions in vertical and horizontal mountings. They compare models with constant or pressure-dependent chamber radius, with or without chamber repositioning on the cross-section, and with a linear or Ogden material law. The strain energy for an Ogden material (Odgen et al., 2013) is
Here, is a material parameter, and is the principal axial stretch. Similarly, Alessi et al. (2023a) presented a dynamic Cosserat rod model for a 3D-printed pneumatic CSM. The model considers the pressure transient with a parameterized smooth step. Also, it captures the manufacturing variability of the actuators using strain gains that tune the pressure-strain relation for each actuator
where is the interior actuator area. An extensive experimental validation includes quasi-static and dynamic motion patterns at various actuation frequencies. Finally, an ablation study evaluates the contribution of different model components, uncovering that neglecting gravity and manufacturing uncertainties could degrade performance up to , depending on the motion.
Concerning fiber-reinforced actuators, Berthold et al. (2022) introduces a Cosserat rod model to deepen the understanding of the pressure-induced deformation behavior of a fiber-reinforced soft pneumatic actuator with three DoFs. By converting the inner chamber pressure into an equivalent force, they consider the compression effects of lateral pneumatic chamber walls. The validation compared the results to FEM simulations, obtaining a maximum mean tip position error of . Differently, Hanza and Ghafarirad (2023) introduces an inhomogeneous Cosserat rod model for a CSM with pneumatic fiber-reinforced actuators, using a nonlinear strain-force relationship to account for the effect of radial pressure. The proposed method outperformed the standard CRT by about .
Alongside, Liu et al. (2023) discusses another modelling approach of a pneumatic fiber-reinforced soft actuator with a semicircular cross section. Here, the EBRT is employed to explore the effects of self-gravity and external loads on the actuator configuration, to predict its shape.
Dou et al. (2023) introduce a hybrid CSM with pneumatic chambers and an antagonistic rod-driven compliant mechanism enabling elongation, contraction, bending, and stiffness regulation. They derive a kinematic model by adapting the EBRT and demonstrate the benefit of the compliant mechanism with an experimental validation comprising way-point tracking.
Recently, Lamping and de Payrebrune (2023) presented a study on soft robotic actuators that utilize pneumatic actuation with multiple chambers. The modeling approach uses the EBRT to characterize the deformation, providing a practical method for accurately determining the rod parameters essential for predicting the performance of these actuators, enhancing their design and control.
Subclass | References | Rod Model | Contribution |
Cable | Oliver-Butler et al. (2019) | EBRT/CRT | Stiffness analysis for various tendon stretch and routing paths . |
Chen et al. (2022) | EBRT | Augments the PCC with a rod model to quickly solve the CSM deformation under . | |
Pneumatic | Sadati et al. (2017c) | CRT | Utilizes the Ritz-Galerkin method to reduce the continuous state space of rod models. |
Alessi et al. (2023a) | CRT | Ablation study of model addressing manufacturing uncertainties in a 3D-printed CSM. | |
Dou et al. (2023) | EBRT | Design and modeling of a hybrid CSM incorporating an antagonistic compliant mechanism. | |
Smart | Goldberg et al. (2019) | CRT | Simulated locomotion of a caterpillar-inspired soft robot actuated by SMA. |
4.4.3 Bend & Stretch with Smart Actuators
Regarding locomotion with smart actuators, Goldberg et al. (2019) presents a model for a caterpillar-inspired soft robot actuated by SMA, leveraging the work by (Bergou et al., 2008) related to the elastic rod theory and the Lagrange motion equation for systems of particles. A planar version of Bergou’s formulation is employed here to model the dynamics of a single SMA actuator. The actuator stiffness and damping parameters are first estimated and subsequently used to simulate the locomotion of the caterpillar robot fabricated from a series of SMA actuators. Interestingly, increasing the actuator stiffness by yields simulation displacements higher than the prototype by .
4.4.4 Bend & Stretch in Parallel Manipulators
The CRT also aided the design, statics modeling, and workspace analysis of an extensible rod-driven PCM with five DoFs at the tip (Wu and Shi, 2022a). A variant of the Levenberg-Marquardt method makes the PCM model suitable for real-time applications. Open-loop errors on the tip position in tracking tasks are lower than in PCC models. Also, they propose a numerical procedure for efficiently estimating partial workspace using a ROM.
4.5 Bend & Twist & Stretch
Five contributions explored bending, twisting, and stretching combined. For cable-driven actuation, it was presented a general geometrically exact static CRT for CSM (Renda and Laschi, 2012). The work highlights the role of torsion deformation. The variety of the design parameters allows the simulation of a wide range of continuum manipulators and the coupled tendon drive for multiple sections. Similarly, a cable-driven conic CSM made of silicone was modeled by combining a geometrically exact dynamic Cosserat rod and the Kelvin model of viscoelastic materials (Wang et al., 2016). An experiment with the arm fixed at the base and oriented horizontally, subject to gravity, without actuation (), validates the viscoelastic model, which matches the experimental data better than a pure elastic model. Finally, the validation executed 2D and 3D motions actuating different cable combinations (). Also, a discrete Cosserat rod combined with a FEM described the dynamics of cable-driven soft robots (Adagolodjo et al., 2021). In particular, they use FEM to model the robot structure and Cosserat for the robot actuation. The implicit backward Euler time stepper integrated the differential equations for numerical stability. Direct and inverse simulations demonstrated the suitability of the approach for modeling robots actuated by cables or rods.
Aimed at colonoscopy, the CRT was employed to reconstruct the shape of a continuum endoscope given the force exerted on the distal end (Ryu et al., 2018). The experimental validation shows that longer endoscopes produce higher tip errors.
Also, Cosserat rods modeled a six-link PCM (Ghafoori and Khalaji, 2020). The model, initially validated for a single rod obtaining an average tip position error of , is applied to a six-link robot with good performance.
In a recent work (Roshanfar et al., 2023), the authors modelled and experimentally validated a soft robot with combined pneumatic and cable actuation. The robot is modelled using CRT where the tangent elastic modulus is expressed as a function of the internal pressure. In addition, the hyperelastic constitutive model was integrated into the Cosserat framework.
4.6 Bend & Stretch & Shear
Two works combine bending, stretching and shearing. For multi-backbone continuum robots, it was proposed a variable curvature statics-kinematic Cosserat rod model incorporating multi-backbone structural constraints (Chen et al., 2021). The model is validated on a robot with two multi-backbone continuum segments, each with four nitinol rods. The push-pull actuation bends the continuum robot. The experiments evaluate the tip positioning accuracy for circular trajectories and the shape discrepancy when the robot is subject to an end-point load in two configurations. They also implement the inverse kinematics in a real-time open-loop controller for a rectangular path following, with and without tip load. The proposed model outperforms the PCC.
Differently, a hybrid model combining the classical CRT with a data-driven stiffness estimation was developed (Bartholdt et al., 2021). Rather than solving numerically nonlinear constitutive equations, the model relies on a work point-dependent linear stress-strain relationship for computational efficiency. Experimental validation is carried out by identifying the stretching and bending stiffness of a soft pneumatic actuator and learning the actuation to stiffness maps using artificial neural networks.
4.7 Bend & Twist & Shear
The only work that performed an experimental validation for the compound class of bending, torsion, and shearing is by Grazioso et al. (2019), who proposes a novel FEM-inspired spatial discretization technique. In particular, the rod is divided into several nodes, defining a helicoidal shape function for the interpolation. Thanks to the specific choice of the shape function, the Cosserat rod dynamics was written with Lie Groups. The method, implemented in a new simulator called SimSOFT, is validated on the pure bending of a cantilever soft arm and the pure in-plane rotation of a soft arm with varying external conditions. In addition, they performed an experimental validation on the Princeton arm to benchmark the coupling of bending, torsion, and shear. The validation results report excellent performance in terms of accuracy and computational efficiency.
Bentley et al. (2023) employ a Cosserat string formulation to derive a tissue and needle force model. It describes normal and frictional forces along the shaft as a function of the planned needle path, friction model and parameters, and piercing force. The proposed force model and associated cost metric are safer and more clinically relevant for motion planning. They fit and validate the model through physical needle robot experiments in a gel phantom. The force model defines a bottleneck cost function for motion planning, which is evaluated against a path-length cost function in random environments, achieving a 62% peak force reduction.
4.8 Stretch & Shear
Combined stretching and shearing deformations are characteristic of peristaltic locomotion (Fig. 5). The peristaltic locomotion of earthworms was studied with a continuous model of compressible and incompressible slender bodies represented as Cosserat rods (Hemingway and O’Reilly, 2021). Incompressibility is enforced as an internal constraint using Green and Naghdi’s theory of a directed rod. Two linearly elastic isotropic material models are assumed. The material parameters are identified from experiments for small deformations. Motion effects from actuators or muscle contraction are modeled as external compressive loads using a doublet function for an assigned centerline force and a uniform pressure for a pair of assigned director forces. The method is showcased in a simulation of a soft robotic device.
4.9 Stretch
The stretching deformation successfully simulated peristaltic locomotion. The forward motion results from wave propagation through contraction and extension of the robot segments along the movement direction (Fig. 5). With bioinspired robotics applications, Nguewou-Hyousse et al. (2019) modeled a caterpillar soft robot using a network of linear and torsional springs and massless rods, modeled as planar discrete elastic rods (Goldberg et al., 2019).
4.10 Discussion on rod models
The analyzed works demonstrate a steady interest in rod models for continuum and soft robots. As shown in Fig. 6, the number of published papers on rod-based models increases.
4.10.1 On the Deformations
Researchers developed numerous rod-based models reformulating and adapting classical rod theories in the soft robotic field. Herein, we reviewed diverse models, which described different combinations of strain modes, giving rise to nine deformation classes (Fig. 4).
The most common class in literature is Bend, with the of the examined papers. This popularity is due to the propensity of slender bodies to bend. In fact, due to the assumption of , with cross-section diameter, the bending stiffness turns out smaller than the stretching or shear ones, easing the bending deformation. In addition, most used actuator paths are linear, which excite predominantly bending modes. These reasons justify the approximation of considering only the bending mode. Furthermore, we can observe that most models in the Bend class leverage the EBRT. However, as summarized in Table 1, several models employed the general CRT. While this represents a computational surplus for trivial bending structures, the generalization of Cosserat models pays off if the model gets deployed in physical interaction tasks.
The next class is All Deformation, with the of the examined papers. As expected, only models based on the CRT achieve all deformations (Table 2). This group of models received a deep study in the last decade, thanks to the rising interest in physical interaction with unstructured environments where all strain modes are significant. Indeed, these models are useful to describe accurately the deformations of slender robots caused by friction (e.g., anisotropic friction for locomotors) or by different mediums (e.g., water). In addition, the most common actuation sources (i.e., pneumatic and cables) excite elongation and compression, causing a significant change in the dimension workspace.
The third class accounting for of the investigated studies is Bend & Twist. These models mainly describe robots in which the unshearability and the unstretchability approximations hold, as summarized in Table 3. Including the twisting deformation allows the representation of backbones with generic 3D curves. However, the twisting deformation can be excited only by specific actuator paths (e.g., helicoidal cables) or asymmetric actuators.
Another significant class, summarized in Table 4, is Bend & Stretch () for the wide diffusion of linear extensible soft pneumatic actuators and tendon-driven robots. Such actuators extend or compress upon solicitation, which induces the continuum body to stretch and bend thanks to their radial disposition on the cross-section.
The remaining five categories are less explored. While the Bend, Twist & Stretch () is a particular case of the All Deformations class, the other classes comprehend few works. The lack of work is due to the difficulty of exciting these specific strain modes with the current actuation technologies or the small number of applications in which the assumptions are valid. In addition, fewer studies that explored shearing per se could be due to the tendency to consider shearing an undesirable or negligible deformation. However, shearing could be significant when there is a prominent environment interaction, as observed in peristaltic locomotion (Hemingway and O’Reilly, 2021) or complex manipulation (Grazioso et al., 2019).
4.10.2 On the Spatial Discretization
One characterizing feature of rod models is the spatial discretization method, which significantly impacts the model’s accuracy and complexity. For simulation and control purposes, the discretized model must be computationally efficient and as accurate as possible. In the analyzed literature, exist two main discretization techniques: Strain Parameterization (Sec. 2.2.3) and Discrete Elastic Rod (DER) (Sec. 2.2.4). The former is based on a discretization directly in the configuration space, allowing the user to choose the degrees of accuracy. In addition, it is possible also to neglect specific strain modes to reduce the computational time. The latter discretizes the length of the rod, considering also the external forces applied to the nodes. Furthermore, most of the existing spatial discretization algorithms are already implemented in dedicated simulators, such as SoRoSim (Mathew et al., 2022) for SP and PyElastica (Naughton et al., 2021) for DER. Straddling geometric and mechanical models, some examined works proposed to fit discrete points with functions, such as PH or Euler’s spirals, which is useful, especially for Shape Estimation.
4.10.3 On the Identification and Validation
Identification and validation of rod models play a crucial role in accurately reproducing the experimental robot behaviors. Despite the high computational cost, it emerged that detailed FEM simulations served a dual purpose:
to assess the correctness of the rod model,
to estimate the unknown parameters.
In the absence of precise FEM models, lumped-mass models with a high number of DoFs were employed. Otherwise, researchers resort to model optimization to match experimental data. In particular, when the robot morphology is not uniform, researchers usually estimate geometrical properties (e.g., cross-sectional area ) and material characteristics (e.g., elastic modulus , damping coefficient ). Experiments typically conducted are with the robot at different base orientations subject to external loads or at the tip. Most rod models are identified and validated through motion data collected with predefined actuation profiles . However, the literature needs rigorous mechanical tests to yield more accurate models, especially for unstructured physical interactions with the environment. Concerning performance metrics, rod models predicted experimental tip positions with Euclidean errors in the range . However, only a few studies assessed the angular errors, which play a role in dexterous manipulators. In summary, the wide use of experimental data suggests an increasing need to reduce the reality gap to develop more accurate controllers.
4.10.4 On the Impact of Materials and Manufacturing
Most of the analyzed rod models employed continuum and soft robots built with diverse materials, including traditional elastomers (e.g., silicone, rubber) or other polymers (e.g., SMA, hydrogels). Each material offers distinct properties, influencing the robot’s compliance, flexibility, and functionality. However, effectively modeling the nonlinear behavior inherent in soft materials using rod theories presents challenges. For instance, some combinations of materials and actuation mechanisms cause notable changes in cross-section areas or volume, which may render the rod-like assumption overly restrictive. Nonetheless, an extension of the CRT considering geometrical rescaling (15) could approximate complex dynamics (16).
The viscoelastic constitutive law of the material impacts significantly on the accuracy. In the classic rod theories, this relation is supposed linear, as already shown in (6). However, especially in the case of pneumatic actuation, the linear constitutive law is no longer valid. To tackle this issue, many analyzed works presented models with a nonlinear constitutive law, using hyperelastic model such as Odgen.
Manufacturing techniques provided continuum and soft robots with unprecedented dexterity (Wallin et al., 2018). However, manufacturing uncertainties may impact model accuracy. Recent rod models addressed manufacturing uncertainties considering variations in material (Alessi et al., 2023a) and geometric properties (Eugster et al., 2022). Despite these advancements, some gaps persist, particularly in capturing hysteresis effects or integrating the dynamic restoring process of self-healing materials into rod-based models. Hysteresis, a prevalent phenomenon in soft materials, strongly impacts the accuracy and predictability of the robot’s behavior. Furthermore, incorporating self-healing materials can potentially enhance the durability and longevity of soft robots but remains relatively unexplored in rod theories (Terryn et al., 2017). Addressing these challenges could advance the capabilities of continuum and soft robots, enabling their integration into diverse applications and real-world tasks.
5 Controllers using rod models
Controlling continuum robots can be very challenging, due to the intrinsic features of this type of robots. Material nonlinearities (e.g., hysteresis), the infinite-dimensional configuration space, and the under-actuated nature of the system make the direct implementation of well-known controllers for rigid manipulators difficult or impossible.
The literature on the control of rigid robots primarily focuses on controlling the end-effector in terms of pose, velocity, or force. In the realm of continuum robots, the natural equivalent of this task is tip control. To exploit the continuum nature of soft mechanisms and achieve tasks that are impossible for their rigid counterparts, the researchers have proposed two other tasks: Multi-Point (MP) control and shape control. The goal of MP control is to control the pose of a set of cross-sections, while shape control aims to regulate the backbone of the robot, fully utilizing the potential of the continuum structures.
To tackle the inherent properties of continuum robots, rod theories have been widely used for control design, resulting in an increasing trend of works in the last decades, as shown in Fig. 6.
The developed controllers fall into two different categories: model-based and learning-based.
5.1 Model-based Controllers
The Model-Based (MB) control strategy involves leveraging prior knowledge of the robot to guide it toward achieving a specific task. The motivation behind developing MB controllers is the possibility of mathematically guaranteeing the stability and performance of the controlled system. However, implementing controllers in physical prototypes can be challenging, often requiring an extensive identification procedure. This context sees the spatial discretization technique as crucial to balance accuracy and computational efficiency. Therefore, the model selection and the discretization technique are integral parts of the control design process.
5.1.1 Formalization
To formalize the control problem, let us consider a generic configuration vector, whatever discretization algorithm is used. According to Della Santina et al. (2023), the shape control problem can be particularized in two sub-problems: (i) regulation and (ii) trajectory tracking (TT). In the last case, the goal of shape control is to find an input such that
where are the desired configuration and its time derivative, respectively. The condition (48) can be particularized in the regulation case, imposing .
Similarly to rigid robots, it is possible to formalize the tip control problem, rewriting (48) in the task space, such as
where and are the desired pose and velocity twist of the tip. Similarly, the tip pose regulation condition can be obtained, particularizing (49) with . Lastly, the MP control problem can be formalized by extending (49) in a set of cross-sections’ desired poses and velocity twists.
To solve these control problems, it is possible to represent the MB control framework in a general scheme, shown in Fig. 7. The block “Sensors” refers to the sensors mounted on the robotic platform, mapping in the measurements . The “Estimator” block implements the Shape Estimation/Reconstruction algorithms, providing an estimation of , , defined as , . Finally, the MB controller block computes the input , processing the error between the estimated configuration and the desired one. More in-depth, the control input is composed of two actions (Fig. 7). The former is the feedforward action , depending on the chosen model, the estimated state and the desired trajectory. The latter is the feedback action that depends on the error w.r.t. the desired configuration and its time derivative. Finally, the specific control algorithm computes the control output combining the feedforward and feedback terms.
5.1.2 Inverse Kinematics
One of the first approaches proposed in the MB literature is the Inverse Kinematics control. The main idea is to use the inverse kinematic model to find the actuation input such that the tip reaches a desired pose or velocity twist. In the above case, the rod-based kinematic models enable the design of controllers that account for the continuum robots’ deformations.
In the planar case, the EBRT provides a simple and efficient kinematic model. Ataka et al. (2020a) proposed a kinematic pose control for an inflatable eversion CSM for trajectory tracking in a planar environment. Assuming a pure bending robot, they used EBRT to model the effect of internal inflation pressure on each bending segment of the prototype. A further advancement in the modeling is the characterization of the variable structural stiffness caused by changes in inflation pressure. Additionally, an observer based on the Extended Kalman Filter estimates the unknown model parameters online from chamber pressure and bending angle measurements. These parameters approximate the actuator-space Jacobian matrix online, which controls the CSM’s tip pose under various stiffness conditions. Simulation and experimental validation using an inflatable CSM made of unstretchable fabric demonstrated that the proposed method achieved position control in a planar environment despite different robot inflation pressures (i.e., stiffness values).
Mauzé et al. (2021) achieve micrometer positioning accuracy with a PCM made of three planar kinematic chains linked to a rigid moving platform. Each chain has a prismatic actuator and a flexible rod that deforms continuously, with bending and twisting deformations modeled with the CRT. A regular Cartesian controller based on inverse kinematics is implemented using the parameters identified for the forward kinematic model. The controller is validated for a circular and square trajectory. For both, large orientation displacements of the platform increase the angular error.
A closed-loop multi-point orientation control strategy for discretely magnetized continuum manipulators was developed (Richter et al., 2021). A non-homogeneous magnetic field model conforming to Maxwell’s equations was formulated and applied to the actuation system to predict the exerted wrenches on the manipulator. The approach is demonstrated on a manipulator containing two permanent magnets actuated inside a non-homogeneous magnetic field. Here, a quasi-static CRT models the manipulator deformation under external wrenches, while a Jacobian provides the actuation inputs. Experiments in simulation and on a real robot demonstrate that the algorithm independently controlled the orientation of two interconnected magnets in a non-uniform magnetic field. However, the control frequency is limited to Hz due to the computational cost of shape reconstruction and the Jacobian.
Campisano et al. (2021) propose a controller for CSM under tip follower actuation. In particular, the strategy is to couple a quasi-static regime and sensor feedback of the actuation sources. This approach achieves (i) rapid computation of the Cosserat rod model, (ii) actuator nonlinearities compensation, and (iii) oscillation reduction. In conclusion, the closed-loop control algorithm is validated experimentally and compared with an open-loop one, to show the effectiveness of the design process.
Renda et al. (2022) proposed an inverse kinematics controller based on the GVS model, using the actuators-strain mapping in (39). The control algorithm is designed for trajectory tracking of a set of points in . The kinematic model is extended, including a map between the length of threadlike actuators and the configuration of the CSM. This formulation covers the general case of actuators’ routing, such as discontinuous or overlapping routing. The control algorithm is validated through simulations on a CSM with parallel and helical actuators’ routing.
5.1.3 Inverse Dynamics
To exploit the mechanical impedance of continuum robots, the controller requires a knowledge of the dynamics of the elastic rod. Furthermore, the dynamic models allow us to consider the influence of external forces, such as gravity or contacts. Therefore, an inverse dynamic model can be used, compensating undesired nonlinear effects and disturbances.
In the case of shape control, Mbakop et al. (2021a) proposed an inverse dynamics controller for a soft finger. The control algorithm is based on a mix of a geometrical and mechanical approach, combining the PH and the Euler-Bernoulli model. While the former allows the reconstruction of the finger’s shape, the latter maps actuators’ magnitude to the positions of PH control points. They experimentally validated the control law with two prototypes: (i) a planar Fluidic Elastomeric Actuator soft finger and a pneumatic CSM. Applying an external tip load to the CSM proved that combining the PH and EBRT improves accuracy and the computational time, w.r.t. using only the PH model.
The subsequent simulation study proposed a new decentralized Inverse Dynamic controller based on the CRT (Doroudchi and Berman, 2021). The controller considers all strain modes and incorporates the system dynamics inside different gain matrices. Each piece or unit of the soft arm is active and independent from the other pieces. This approach increases the computational efficiency and successfully tracks configurations requiring shear deformations with average normalized root-mean-square errors below .
5.1.4 Feedback Linearization
Continuum robots have an inherent under-actuation, which makes control design more difficult. Feedback Linearization (FL) is a widely used method for controlling under-actuated mechanisms (Spong, 1994). The main idea of this control algorithm is to linearize the system with a feedback control action, compensating all the nonlinear terms. The FL controller can be combined with a linear controller (e.g. Linear Quadratic Regulator), in order to achieve the desired performances.
For a task-space control, Rucker et al. (2022) proposed a controller for continuum robots based on a highly under-actuated planar Kirchhoff rod (i.e., the EBRT). The main idea is to show that a maximal coordinate form of the Kirchhoff model is suitable for control purposes, exploiting the under-actuation of the system. The controller is derived by solving a constrained Lagrangian problem, discretizing the length of the Kirchhoff rod. The control algorithm is based on an input-output Feedback Linearization combined with an Sliding Mode Control (SMC) in an outer loop. An observer based on the constrained model formulation provides control feedback, adding corrective virtual external forces based on the estimation error. Lastly, simulations of an actual two-DoFs soft prototype evaluate the control framework and demonstrate the consistency of the algorithms during tip position trajectory tracking. The numerical examples of the SMC also demonstrate robustness to parameter variations.
Caradonna et al. (2024) presented a 2D soft segment with a revolute joint at its base controlled by a FL controller. By analyzing the equilibrium map of the system’s zero dynamics, they proposed a method to employ the system for the pick-and-place task, accounting for the added mass from the object. They utilize the FL controller with the Curvature Parameterization (27) and Strain Parameterization (11) and compare the two discretization techniques emphasizing numerical robustness, computational cost, and accuracy. The control algorithm and pick-and-place strategy validated through simulations demonstrated consistency.
5.1.5 Energy-Shaping
Another well-known control algorithm designed for under-actuated systems is the Energy-Shaping Control (ESC) (Spong, 1996). The main idea is to find an input such that the system reaches a desired energy state. The rod models are particularly effective in this case, thanks to the Hamiltonian formulation of the elastic rods.
Caasenbrood et al. (2021, 2022) proposed an ESC for CSM for tip pose regulation. To apply the control algorithm, the Authors reformulated the GVS model in a Port-Hamiltonian form, showing the passivity condition of the system and enabling the design of the ESC. The control law is validated through numerical examples, proving the capability of the controlled system to reach the desired tip pose. Furthermore, thanks to the properties of the specific control algorithm, both under-actuation and hyper-redundancy are effectively managed. The controlled system demonstrates smooth convergence to the target, and a pseudo-joint strategy similar to that observed in nature.
Chang et al. (2023) introduced an ESC for a CSM inspired by an octopus tentacle employing the CRT with the DER discretization. They modeled three main types of muscles: transverse, longitudinal, and oblique. The transverse muscles necessitate the DER approach due to the invalid assumption of a rigid cross-section. Leveraging the system’s passivity properties, they designed an ESC by solving an optimal control problem based on Hamiltonian principles. Additionally, they developed an efficient method to design a task-specific equilibrium configuration by solving an optimization problem in the group. The model and ESC are validated on the PyElastica simulator, controlling the CSM for reaching and grasping a static object.
Tiwari and Banavar (2023) proposed a ESC for a 2D flexible single-link manipulator, based on CRT. In particular, the model used is geometrically exact, i.e. the manipulator’s configuration space lies on an infinite-dimensional Lie Group. To implement the controller in the discrete-time domain, the Authors used a lie group variational integrator, derived from the Lagrangian variational principle of the flexible single-link manipulator (Tiwari and Banavar, 2023). The results in silico indicate that the discrete closed-loop system achieves shape regulation during point-to-point rotational maneuvers.
5.1.6 Sliding Mode
An alternative to the above controllers for under-actuated systems is the SMC (Xu and Ümit Özgüner, 2008), due to the inherent robustness. The main idea of this strategy is to control the nonlinear system with a discontinuous control signal, forcing it to a specific curve on the phase space (e.g., a line).
An adaptive SMC based on CRT is proposed by Alqumsan et al. (2019). The controller is designed for planar tip position control, which considers the saturation of the actuation sources. Furthermore, the marginal stability of the proposed adaptive control law is provided. The validation is conducted in simulation, where the PDE of the Cosserat rod are solved by the Generalized- method (Chung and Hulbert, 1993), an implicit solver that assures unconditional stability and numerical dissipation for high-frequency noise.
Subsequently, the high nonlinearity of a large-scale and ultralight two-section inflatable CSM made of TPU-coated Nylon is captured by a dynamic Cosserat rod recursively to predict the behavior of the thin-walled and incomplete continuum structure (Li et al., 2021). An extended inverse dynamic model compensates for the error in a feed-forward manner. Moreover, stiffness and damping matrices are calibrated on deflection data and a weight release experiment, respectively. The study also compares the rod model versus a rigid-body and a PCC model. Finally, three closed-loop model-based controllers for trajectory tracking are compared to further demonstrate the viability of the CRT. Here, an adaptive SMC generated the desired pressure differences in the presence of model uncertainties. The chattering phenomenon caused by the conventional SMC is attenuated by integrating a fuzzy logic control, using continuous membership functions to replace the discontinuous .
Mishra et al. (2023) proposed a Terminal SMC (Venkataraman and Gulati, 1993) for tip position TT control. The model on which the controller is based consists of a combination of CRT and fractional Bounc-Wen model (Mishra et al., 2022), taking into account the hysteresis effect. In this work, the Terminal SMC is robust, fast, non-singular, and with exponential convergence. In addition, an obstacle avoidance algorithm is presented, using a combination of cumulative danger field and damped least square methods. The model and the controller are validated in a simulated version of Robotino-XT (Wang et al., 2022a).
5.1.7 Model Predictive Control
The Model Predictive Control (MPC) (Rawlings et al., 2017) is a well-known controller based on iterative and finite-horizon optimization. Furthermore, it is possible to optimize while satisfying constraints, such as actuators’ saturation.
Towards MIS, Soltani et al. (2017) introduce a 3D, stable, non-linear model-based hybrid force/position controller to apply a constant force to the cardiac tissue while tracking the desired trajectory. A quasi-static Cosserat rod models the distal shaft of tendon-driven catheters, reformulated for controller design. The position controller is based on a non-linear MPC tracker satisfying input constraints and local stability conditions. The position control operates in an inner loop, while the force control is in an outer control loop. Simulations conducted with a bidirectional ablation catheter predict the control performance applied in catheterization. Finally, experimental validation is carried out on a real pig heart, though the authors suggest the necessity of more computational efficiency and incorporation of imaging modalities with position tracking systems.
Recently, a novel motion planning method of fish-like soft robots actuated by macro-fiber composite pairs was developed (Barbosa et al., 2023). The structure should mimic oscillatory and undulation movements, accomplished if the amplitude of the tail motion is greater than that of the head motion. The motion planning adopts MPC for generating the actuation signals. This model-based strategy requires a small-sized model. Therefore, they extracted a linear modal state-space model from a free-free dynamic EBRT considering the electro-mechanical coupling of the actuator pairs. Moreover, design strategies like the addition of concentrated and distributed masses were addressed to mimic fish-like motion, guaranteeing suitable mode shapes for the structure. Four numerical case studies demonstrated the capability of the proposal for deriving bounded input signals that generate oscillatory and undulation movements even with disturbances.
5.1.8 Miscellaneous
An open-loop optimal control method to control an octopus-inspired planar arm modeled with CRT was proposed (Wang et al., 2021b). The main goal is to use the optimal control approach to replicate two stereotypical motions: (i) reaching (i.e., bending propagation from the base to the tip) and (ii) fetching (i.e., creation of pseudo-joints). The functional cost of the optimal problem is the elastic potential and actuation energy of the arm, with a terminal cost related to the specific task. An iterative forward-backward algorithm solves the optimization. The proposal validated in silico demonstrated the emergence of propagating waves, as observed in stereotypical strategies.
Conversely, Doroudchi et al. (2018) designed three controllers based on the dynamics of a linear Euler-Bernoulli rod. Thanks to the linearized form of the EBRT, the authors can apply the optimal control theory easily. The novelty here is that the optimal controllers are decentralized, with a modular soft robotic arm where each piece has independent actuation and sensorization. They observe that ill-conditioned system matrices, common in discretized beam models, may hinder control convergence. The iterative algorithm is also very sensitive to the initial condition. Furthermore, a third controller based on the gradient descent method provides a bounded solution for the controller gains and reduces the sensitivity to initial conditions. Finally, simulations validated the approach.
To control aerial continuum manipulation systems, Hashemi et al. (2023) proposed a terminal synergetic controller to accomplish tip position TT. After presenting a novel platform that integrates a quadrotor with a tendon-bent CTR, they designed a two-loop control architecture. In particular, the first loop implements a terminal synergetic controller to ensure the continuum robot’s tip follows the desired trajectory. The second loop’s controller manages the quadrotor’s position and velocity. In addition, this last control algorithm is based on a combination of synergetic control and a hybrid feedback law. For the first time, a synergetic controller is applied in the literature, showing robustness against disturbances and model uncertainties. Finally, they validate the control framework simulating an aerial continuum manipulation system.
Controller | References | Rod Model | Task | Validation | ||||||||||||||||||||
Inverse Kinematics |
Inverse Dynamics |
Feedback Linearization |
Energy-Shaping |
Sliding Mode |
Model Predictive Control |
Open-Loop Optimal Controller | Wang et al. (2021b) | CRT (DER) | Reaching, Fetching | Numerical | ||||||||||||||||||||
Doroudchi et al. (2018) | EBRT | Shape Regulation | Numerical | |||||||||||||||||||||
Synergetic Control | Hashemi et al. (2023) | CRT | Tip Position TT | Numerical |
5.1.9 Shape/State Estimation
As seen in Sec. 5.1.1 and in Fig. 7, a general MB shape controller requires feedback in terms of configuration vector and its time derivative . Depending on the spatial discretization, the vector could be difficult or impossible to measure. To tackle this issue, a general MB control framework requires an Estimator, which maps the measures in the estimated state . For continuum and soft robots, this problem is referred to as Shape/State Estimation or Shape/State Reconstruction. Currently, most Shape Estimation methods use geometrical models or Computer Vision algorithms (e.g., (Bezawada et al., 2022; AlBeladi et al., 2021)). However, incorporating prior knowledge of rod-based models can significantly enhance performance.
In the pioneering work of Trivedi and Rahn (2014), the Authors suggest three approaches for Shape Estimation, based on CRT: (i) mounting load cells at the base, (ii) employing cable encoders, and (iii) mounting an inclinometer at the end of each piece. A maximum tip position error of is observed in simulations. Lastly, they use the OctArm VI (Grissom et al., 2006) to experimentally validate all three approaches. Another study proposes a real-time shape-estimation method based on the force-torque for CTR measured at the tubes’ basis (Donat et al., 2021). It extends a shape estimation algorithm for elastic Kirchhoff rods. They modeled a CTR combining planar PCC segments lying on different equilibrium planes. The approach is evaluated with single and two combined additively manufactured tubes at high frequency, achieving a mean deviation of mm along the tube. Differently, Lilge et al. (2022) develop a model for shape and strain estimation of a continuum cable-driven robot using a Gaussian Process regression, a mathematical tool for estimating continuous-time trajectories in . The idea involves substituting (i) time with arc-length and (ii) kinematic and dynamic laws based on CRT. This method efficiently estimates the robot’s shape using noisy measurements from sensors such as strain gauges, tracking coils, and an external camera. The real-world experiments provide excellent performances, with an average tip pose error of mm and . Finally, Yousefi et al. (2023) propose an optimization-based method that simultaneously estimates the shape and the forces acting on a continuum robot by employing the quasi-static CRT. Magnetic localization determines the position of multiple robot points. The method estimates the robot’s shape and force in a wide range of conditions validated experimentally, with and without knowledge of the contact positions.
5.2 Learning-based Controllers
Learning-based controllers leverage the physics of rod models and utilize Machine Learning (ML) to derive parameterized control policies , which provide and end-to-end mapping between desired poses and sensor measurements to motor commands :
The vector of policy parameters are the weights of ANN trained via Supervised Learning (SL) or RL.
We report learning-based controllers in chronological order to highlight the evolution of the literature.
5.2.1 Supervised Learning
SL is viable to address the soft robotic control problem for simple tasks. It relies on pseudo-random motion data of actuations and corresponding task space to train an inverse model (50).
Initial learning strategies used rod models to rapidly prototype controllers. For instance, Thuruthel et al. (2017) developed dynamic models of CSM using recurrent neural networks and presented a trajectory optimization method for task space control. They validated the controller on a Cosserat rod with PCS parameterization and a pneumatic CSM. Recently, Wang and Rojas (2022) developed a closed-loop controller of continuum robots, employing RNN to approximate the forward and inverse dynamics. They extended the CRT in (Till et al., 2019) to account for spine compression and collect motion data for training their forward models. A non-parametric Gaussian process regression compensates for the error between the real robot and the RNN. A hybrid controller alternates between two stand-alone policies. They validated the control architecture on reaching and tracking tasks using a tendon-driven continuum robot with 3D-printed compressible spines. The results show that combining simulated and experimental data improves the control accuracy.
5.2.2 Reinforcement Learning
RL is one of the main ML paradigms and a versatile framework suited for decision-making and control (Sutton and Barto, 2018). The two fundamental objects of an RL problem are the agent and the environment. In soft robotic control, the agent is a learning-based controller implementing an ANN, whereas the environment is everything outside the agent, comprising the soft robot and other objects. Fig. 8 depicts the agent-environment interaction highlighting the role of ANN and rod models. The agent and the environment interact at discrete time steps . At every interaction step, the agent observes the (possibly partial) state of the environment including sensor readings and task information. The agent acts according to a policy to select actions taking the form of low-level controls like torques or pressures. A policy can be a deterministic mapping from states to actions, . Alternatively, a policy can be stochastic, mapping a state to a distribution over actions . The environment transitions into another state according to the dynamics of the rod model. The agent also perceives a scalar reward signal from the environment, indicating the goodness of the current state specifying the control task. The goal of the agent is to maximize its return, the cumulative sum of rewards:
where is a discount factor. Therefore, the goal of RL is to find an optimal policy that maximizes the expected return
The following deep RL contributions are summarized chronologically in Table 6. Satheeshbabu et al. (2019) applied deep Q-learning (DQL) with experience replay for the quasi-static position control of a pneumatic CSM capable of bending and twisting. An open-loop policy was trained in simulation utilizing a static Cosserat rod. They validate it in simulation and on the physical platform, subject to various external loads. To cope with the high-dimensional state-action space in soft robotics, DQL (Mnih et al., 2013) employs an ANN to approximate the value function of state-action pairs, defined as
A deterministic policy can leverage to select the best action in a state as .
The following work increased the dexterity of the CSM and trained a closed-loop controller for precise quasi-static positioning via deep deterministic policy gradient (DDPG) (Satheeshbabu et al., 2020). Experiments in simulation investigated the robustness of the control policy for effects of loading, reachability, and workspace discontinuity. Finally, the controller was deployed on the prototype. The DDPG algorithm also derived a closed-loop controller for the quasi-static positioning of a pneumatic CSM modeled with the KRT (Uppalapati et al., 2020). Interestingly, the CSM was integrated into a mobile platform including a rigid arm and a sensorized gripper. The system was teleoperated and validated in the agricultural task of picking berries using different maneuvering strategies.
The potential of continuum mechanical models is fully exploited in physical interaction tasks. Naughton et al. (2021) effectively applied several deep RL algorithms to train various control policies. The control tasks included point reaching, trajectory tracking, and maneuvering through structured and unstructured obstacles. The controllers were learned and tested in simulation using a synthetic CSM based on a dynamic CRT. The rod was actuated by applying distributed internal torques, modeled via splines characterized by control points and vanishing values at the rod ends. The work showcased a captivating interaction between the rod and the environment. In particular, the maneuvering task required the free end of the rod to reach a position behind fixed obstacles. The task was solved with a reward function that only considered the distance to the target and a penalty term for the occurrence of numerical errors
Interestingly, the rod learned to navigate and interact with the obstacles without an explicit reward term. Finally, Alessi et al. (2023b) applied proximal policy optimization (PPO) to train a closed-loop position control policy for dynamic trajectory tracking with a pneumatic CSM. The controller was trained in simulation leveraging a dynamic Cosserat rod model of the CSM (Alessi et al., 2023a). After training, simulation tests evaluated the generalization capabilities of the policy to new observations, dynamics, and tasks. The experiments included tracking new trajectories subject to unknown external forces or using different material properties (e.g., ). In addition, the learned controller was transferred without retraining to intercept a moving object. The policy has shown good generalization capabilities for all four tests. Subsequently, Alessi et al. (2024) elaborated on the previous simulated environment to account for the dynamics of a dexterous CSM and its interaction with objects. They developed a learning-based control architecture for pose/force control, which enabled dynamic pushing with CSM in the real world. To mitigate the significant sim-to-real gap in soft robotics, they originally applied domain randomization (DR), an effective sim-to-real technique (Zhao et al., 2020b), emphasizing soft materials.
Controller | References |
Task |
Open loop ANN | Satheeshbabu et al. (2019) | CRT () | DQL | Tip Position TT (way-point) | (None) | ||||||
Closed loop ANN | Satheeshbabu et al. (2020) | CRT () | DDPG | Tip Position TT (way-point) | (None) | ||||||
Uppalapati et al. (2020) | KRT () | DDPG | Tip Position TT (way-point) | (None) | |||||||
Naughton et al. (2021) | CRT () | Several | Tip Pose TT, Obstacle maneuvering | ||||||||
Alessi et al. (2023b) | CRT () | PPO | Tip Position TT, Object interception | ||||||||
Alessi et al. (2024) | CRT () | PPO | Tip Pose/Force for Pushing | (DR) |
5.3 Discussion on Rod-based Controllers
5.3.1 On the Model-based Controllers
The emergence of computationally efficient and accurate models enables researchers to develop model-based controllers that can accurately predict and react to the nonlinear nature of the elastic rods. As shown in Table 5, the controllers based on Inverse Kinematics are the most popular choice, achieving tip trajectory tracking tasks in real-world prototypes. Despite the simplified model, the feedback action is able to reject the unmodelled effects, showing the usability of that strategy.
However, a dynamic model is essential to exploit the properties of continuum robots. Most of the proposed controllers adopt well-known approaches from the under-actuation literature, such as FL, ESC, SMC, or MPC. However, only a few proposed controllers presented an experimental validation. The reasons behind this sim-to-real gap could be the lack of robustness or physical limits of the prototypes. While the Inverse Dynamics, FL, and ESC are sensitive to unmodelled effects and parametric uncertainties, the discontinuous output of the SMC is difficult to implement with the most popular actuation technologies (e.g., tendons, pneumatic). Specifically, the actuators’ internal dynamics may limit the effective bandwidth of the system, hence filtering the discontinuous output of the controller. The MPC is a promising but not widely explored controller in the continuum robots literature.
Another relevant aspect of the analyzed works is the discretization technique employed. Two main aspects eased the control design: (i) the advancement of discretization techniques and (ii) the tunability of the accuracy. The former allows the researchers to design the controller with a finite number of state variables and compute efficiently the robot’s kinematics or dynamics. Selecting appropriate discretization techniques is crucial, particularly for determining the minimum frequency necessary to ensure numerical stability. In addition, the DER technique takes into account also cross-section deformations, especially useful for modeling and controlling muscles in the case of bio-inspired robotic systems (Chang et al., 2023). The latter aspect allows for the choice of accuracy levels and the number of strain modes the model can predict. A common approach simplifies the model and designs a controller that treats unmodeled strains as external disturbances. This strategy can enhance the speed of the control loop and its robustness.
Regarding research gaps, it is possible to note the lack of experimental validation for MB controllers. Furthermore, there are few studies involving external forces due to interactions with environments, particularly because of the complexity involved in modeling and estimating contact and friction forces. Further developments on model-based controllers could benefit from using physics-inspired ANN to lessen the computational burden of complex expressions (e.g., the Coriolis matrix) or use them to compute the EoM of the elastic rod, as already explored in (Lutter and Peters, 2023; Liu et al., 2024). Finally, the literature on rod-based planning (Mishra et al., 2023; Bentley et al., 2023) is sparse, likely due to the current maturity of the controllers. Recently, some planning algorithms using only geometrical approaches have been proposed (Rao et al., 2024).
5.3.2 On the Learning-based Controllers
Learning-based controllers utilize the physics of rod models in various applications, with a clear distinction between SL and RL.
Rod models have a marginal role in SL strategies, which used rod models as proxies for motion data collection under predefined open-loop policies. The motion data were employed to train ANN, which efficiently emulate forward or inverse robot models, and replaced rods to facilitate the preliminary validation of control hypotheses. Although SL was widely used in early soft robotic controllers, it limits the robot tasks to simple reaching and tracking due to the challenges in collecting supervised data for physical interaction tasks. Therefore, this approach could fade away in the future.
Conversely, the role of rod models is central for developing complex deep RL control policies, especially to achieve unstructured physical interaction with CSM. Deep RL agents exploited computationally feasible rod models embedded in simulated environments to learn optimal iteratively. Herein, RL not only replaces the burden of collecting and labeling datasets offline but also unlocks a new range of possibilities not fully explored yet. The literature review highlighted the role of rod models in RL policies. Indeed, as reported in Table 6, control tasks evolved from position control for quasi-static tracking to simulated pose control for obstacle maneuvering and pose/force control for dynamic pushing. For instance, initial policies limited the physical interaction to static payloads without addressing the sim-to-real gap (Satheeshbabu et al., 2020). On the other hand, efforts in developing simulated environments (Naughton et al., 2021) combined with careful reward engineering and sim-to-real techniques (Alessi et al., 2024) facilitated recent advancements in physical interaction with CSM. In summary, rod models combined with recent deep RL methods and sim-to-real transfer techniques (Zhao et al., 2020b) advanced the state of the art of soft robot control.
Future developments could explore model-based RL. Here, the term model-based refers to the agent not only learning a policy but also learning a predictive model of the environment used for planning (Moerland et al., 2023). Alternative research lines can be borrowed from rigid robotics and tailored to the nonlinearities of continuum and soft robots (Ibarz et al., 2021).
6 Conclusion
This paper provided an overview of the modeling and control of continuum and soft robots using rod theories. First, we formulated the four main rod theories used in robotics under a unified mathematical framework. Subsequently, we provided mathematical tools for modeling the actuation in soft robotics. This was followed by a classification and summary of the rod-based models for continuum and soft robots across nine classes of deformations, offering a unique perspective on the evolution of the state of the art, together with a discussion on research trends and gaps that emerged. Furthermore, we reviewed recent control strategies leveraging rod models to perform various tasks, followed by a discussion on the encountered approaches as well as our insights. Through an in-depth exploration of various rod-based models and control strategies, the diverse range of applications of rod theories demonstrates their capabilities to capture the complex mechanical behaviors of continuum and soft robots. Despite the remarkable advancements in incorporating rod theories into soft robot models and controllers, there are still areas requiring further research. For instance, accurate modeling of processes inherent in soft materials, such as hysteresis, is still a challenge. Related to the control, addressing efficient sim-to-real policy transfer in learning-based approaches is essential for real-world robot deployment.
This work was supported by the Future and Emerging Technologies (FET) program of the European Commission, through the PROBOSCIS project (G.A. ID 863212), by the US Office of Naval Research Global under Grant N62909-21-1-2033, and in part by Khalifa University under Awards No. RIG-2023-048, RC1-2018-KUCARS
Appendix A Query for models and controllers
The two Scopus advanced search queries used for the models and controllers are respectively:
TITLE-ABS-KEY (“soft” OR “continuum”) AND TITLE-ABS-KEY (robot* OR “arm” OR “manipulator”) AND TITLE-ABS-KEY (“Cosserat” OR “Kirchhoff” OR “Timoshenko” OR “Euler-Bernoulli” OR “Geometrically exact” OR “rod” OR “beam”) AND (theor* OR model*).
TITLE-ABS-KEY (“soft” OR “continuum”) AND TITLE-ABS-KEY (robot* OR “arm” OR “manipulator”) AND TITLE-ABS-KEY (“Cosserat” OR ”Kirchhoff” OR “Timoshenko” OR “Euler-Bernoulli” OR “Geometrically exact” OR “rod” OR “beam”) AND (control*).
Appendix B Lie Algebra
The Tilde Operator is defined as
(55) For a vector , its application gives an skew-symmetric matrix
(56) -
The Hat operator is defined as
(57) Let be . The Hat operator gives
(58) -
The Vee operator is the inverse of the Hat operator
(59) Consequentially, it is possible to write .
The Adjoint Representations consist in two maps: and . The former is defined as
(60) where . The latter is defined as
(61) Furthermore, it is possible to define also the co-adjoint operator , that is

### v2

@@
 - piecewise constant strain
+- GVS
+- Geometrical Variable Strain
+- ISP
+- Implicit Strain Parameterization
 - DER
 - Discrete Elastic Rod
+- POD
+- Proper Orthogonal Decomposition
+- SVD
+- Singular Value Decomposition
 - CSM
@@
 - trajectory tracking
+- IK
+- Inverse Kinematics
+- ID
+- Inverse Dynamics
 - MPC
@@
 - minimally invasive surgery
-Carlo Alessi & Egidio Falotico, Brain-Inspired Robotics Laboratory, The BioRobotics Institute, Scuola Superiore Sant’Anna, Pisa, Italy.
+Carlo Alessi, Istituto Italiano di Tecnologia, Genoa, Italy (carlo.alessi@iit.it).
+Egidio Falotico, The BioRobotics Institute, Scuola Superiore Sant’Anna, Pisa, Italy (egidio.falotico@santannapisa.it).
 Rod models in continuum and soft robot control: a review
 Abstract
-Continuum and soft robots can positively impact diverse sectors, from biomedical applications to marine and space exploration, thanks to their potential to adaptively interact with unstructured environments.
-However, the complex mechanics exhibited by these robots pose diverse challenges in modeling and control.
-Reduced order continuum mechanical models based on rod theories have emerged as a promising framework, striking a balance between accurately capturing deformations of slender bodies and computational efficiency.
-This review paper explores rod-based models and control strategies for continuum and soft robots.
-In particular, it summarizes the mathematical background underlying the four main rod theories applied in soft robotics.
-Then, it categorizes the literature on rod models applied to continuum and soft robots based on deformation classes, actuation technology, or robot type.
-Finally, it reviews recent model-based and learning-based control strategies leveraging rod models.
-The comprehensive review includes a critical discussion of the trends, advantages, limits, and possible future developments of rod models.
-This paper could guide researchers intending to simulate and control new soft robots and provide feedback to the design and manufacturing community.
+Continuum and soft robots can transform diverse sectors, including healthcare, agriculture, marine, and space, thanks to their potential to adaptively interact with unstructured environments.
+These robots exhibit complex mechanics that pose diverse challenges in modeling and control.
+Among various models, continuum mechanical models based on rod theories can effectively capture the deformations of slender bodies in contact-rich scenarios.
+This structured review paper focuses on the role of rod models in continuum and soft robot control with a vertical approach.
+We provide a comprehensive summary of the mathematical background underlying the four main rod theories applied in soft robotics and their variants.
+Then, we review the literature on rod models applied to continuum and soft robots, providing a novel categorization in deformation classes.
+Finally, we survey recent model-based and learning-based control strategies leveraging rod models, highlighting their potential in real-world manipulation.
+We critically discuss the trends, advantages, limitations, research gaps, and possible future developments of rod models.
+This paper aims to guide researchers who intend to simulate and control new soft robots while providing feedback to the design and manufacturing community.
 keywords:
 Soft Robotics; Modeling, Control, and Learning for Soft Robots; Biologically-Inspired Robots1 Introduction
-Continuum robots are defined as robots with distributed deformations throughout their structure, resulting in infinite degrees of freedom (DoFs).
-This characteristic gives rise to a hyper-redundant configuration space, allowing the robot tip to reach any point in the three-dimensional (3D) workspace with a virtually infinite number of configurations (Trivedi et al., 2008c).
-Soft robots, built with soft materials or deformable structures (Laschi et al., 2016), are continuum robots.
-Generally speaking, soft robots assume a variety of morphologies.
-In this paper, we refer to both classes with their partial overlap.
-Continuum and soft robots deform elastically while safely interacting with unstructured environments thanks to the inherent compliance of soft materials (Laschi and Cianchetti, 2014; Laschi et al., 2016).
-Novel design and manufacturing technologies are accelerating the prototyping of highly dexterous robotic systems (Rus and Tolley, 2015; Wang et al., 2024).
-However, the modeling and control of continuum and soft robots are still challenging problems.
-Indeed, soft materials cause extreme hyper-redundancy and exhibit nonlinear properties such as hysteresis and stress softening.
-These challenges inspired the development of diverse modeling techniques (Armanini et al., 2023; Gilbert, 2021), each characterized by underlying assumptions, mathematical framework, and a trade-off between accuracy and computational cost.
-Soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical (Fig.
-1).
-Concurrently, innovative control strategies emerged to leverage these models (George Thuruthel et al., 2018; Della Santina et al., 2023).
-Popular modeling methods are inherently data-driven (Laschi et al., 2023).
-Despite not requiring geometric and physical expertise, they rely on pseudo-random motion data to train artificial neural networks mapping actuation to task space (Chin et al., 2020).
-This robot-independent method effectively derived efficient forward models employed in learned policies for tracking (Centurelli et al., 2022; Piqué et al., 2022) and throwing (Bianchi et al., 2023).
-However, they are black-box, could suffer overfitting, and the burden of collecting interaction datasets limits their applicability.
-Other approaches, like discrete material models, discretize continuous bodies a priori.
-For example, pseudo-rigid models represent them with a chain of rigid links connected by joints.
-While these provide satisfactory results for hyper-redundant arms (Venkiteswaran et al., 2019), they poorly approximate elastic structures.
-In addition, lumped-mass models employ an array of masses, springs, and dampers (Habibi et al., 2020).
-This modularity can model complex phenomena but could require intensive system identification.
-Alternatively, geometrical models utilize geometrical assumptions on the deformed shape of the soft body.
-For instance, functional models describe the deformations using polynomials (Della Santina and Rus, 2019).
-In addition, piecewise constant curvature (PCC) models discretize continuous bodies in circular arcs with constant curvature (Xie et al., 2022).
-These models were effectively used within proportional-derivative (PD) task space controllers for various dynamic tasks (Fischer et al., 2022).
-Since most continuum robots have actuators parallel to the centerline producing constant-curvature deformations, PCC is viable for uniform and lightweight robots (Webster III and Jones, 2010).
-However, the suitability of these approximations could degrade when robots are subject to significant external forces or interactions with the environment.
-Finally, continuum mechanical models characterize soft robots with continuous configuration spaces and define deformations in physical terms.
-They enable the simulation of physical interactions and the study of robot mechanics.
-The classical finite element method (FEM) accurately represents complex 3D geometries to the extent of high computational costs and involved mathematical formulations (Coevoet et al., 2017).
-Since such models would impede real-time control applications, their use is often limited to designing and simulating soft robotic components (Xavier et al., 2021) or as a benchmark.
-Nonetheless, recent model order reduction (MOR) techniques are making 3D mechanical models more affordable (Allard et al., 2007; Dubied et al., 2022).
-Since continuum and soft robots are slender, other methods employ reduced order models of elastic rods, one-dimensional (1D) objects with a dimension (the length) much larger than the others.
-Elastic rods effectively describe the deformations of slender bodies (i.e., stretching, shearing, bending, twisting), balancing computational efficiency and an accurate representation of complex mechanics.
-Therefore, rod theories are promising candidates for soft robotics modeling and control, which are the focus of this paper.
-1.1 Article Contribution
-In this study, we investigate the role of rod theories in the modeling, simulation, and control of continuum and soft robots. The main contributions are:
-1.
-review of the fundamentals of four major rod theories with a common formulation, which facilitates the comparison and highlights connections to other modeling strategies;
-2.
-review and analysis of rod models applied to continuum and soft robots, originally classified on deformation classes; and
-3.
-review and analysis of recent model-based and learning-based controllers leveraging rod models.
-First, we summarize the mathematical fundamentals underlying the four main rod theories applied in soft robotics: Cosserat-Reissner, Kirchhoff-Love, Timoshenko, and Euler-Bernoulli (Sec.
+Continuum robots are defined as robots with distributed deformations throughout their structure, resulting in an infinite number of degrees of freedom (DoFs).
+This property creates a hyper-redundant configuration space, allowing the robot tip to reach any point in the three-dimensional (3D) workspace with a virtually infinite number of configurations (trivedi2008SoftRob).
+Soft robots, built with soft materials or deformable structures (laschi2014soft), are a subset of continuum robots.
+Generally speaking, soft robots can assume a variety of morphologies.
+In this paper, we refer to both classes, considering their partial overlap.
+Continuum and soft robots can deform elastically while safely interacting with unstructured environments thanks to the inherent compliance of soft materials (laschi2016soft).
+Advances in design and manufacturing technologies are rapidly accelerating the development of highly dexterous robotic systems (rus2015design; WangYanmei2024Advancements).
+Despite this progress, modeling and controlling these robots remain challenging problems.
+Indeed, the use of soft materials introduces extreme hyper-redundancy and complex nonlinear behaviors, including hysteresis and stress softening.
+One of the earliest continuum formulations for hyper-redundant robots represented a manipulator as a smooth backbone curve with distributed parameters and derived partial differential equations for its dynamics (chirikjian1993continuum; chirikjian1994hyper).
+Since then, diverse modeling techniques were developed (armanini2023soft; gilbert2021mathematical), each characterized by underlying assumptions, mathematical frameworks, and trade-offs between accuracy and computational cost.
+As summarized in Figure 1, soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical.
+Concurrently, innovative control strategies emerged to leverage these models (george2018control; della2023model; falotico2024learning).
+In this paper, we focus on the achievements of rod models.
+Rod models were widely employed in continuum and soft robotics, representing a fundamental framework for modeling, simulation, and control.
+Despite their extensive use, the literature lacks a comprehensive, structured review that consolidates key advancements and critically examines the contributions of these models.
+To address this gap, this paper investigates the role of rod theories in continuum and soft robotic systems from a vertical perspective.
+We guide the reader from the mathematical formalization of rod theories, through their application in modeling continuum and soft robots, to recent developments in model-based and learning-based control.
+By structuring the literature on deformation modes, we aim to provide not only a synthesis of the state of the art but also a practical reference for researchers who want to model continuum and soft robots for control applications.
+First, we highlight related surveys in continuum and soft robotics that provide valuable complementary perspectives, where rod models are treated marginally (Sec.
 2).
-The same notation facilitates finding similarities and differences between the rod theories and supports the analysis of deformation classes.
-Further, it highlights theoretical insights on spatial discretization (i.e., positions vs configurations), connections to other models (e.g., the polynomial curvature), and implications on the control problem.
-Also, we consider the general modeling of actuation-induced distributed loads with different actuator routings (Sec.
+Second, we characterize the major modeling techniques to identify their strengths and limitations with respect to rod models (Sec.
 3).
-Then, we review and discuss rod models applied to continuum and soft robots classified by deformation, and eventually robot types or actuation systems (Sec.
+Then, the main contributions of this paper are as follows:
+A comparative review of four major rod theories and discretization techniques, including Cosserat-Reissner, Kirchhoff-Love, Timoshenko, and Euler-Bernoulli.
+We present them within a consistent mathematical framework to facilitate comparison and highlight connections to alternative modeling approaches (Sec.
 4).
-The analyzed models could guide researchers intending to simulate new soft robots in selecting the most suited rod theory.
-In addition, it could provide feedback to the design and manufacturing community.
-Finally, we review and discuss recent model-based and learning-based controllers that leverage rod models (Sec.
+For completeness, we consider the general modeling of actuation-induced distributed loads with different actuator routings (Sec.
 5).
-Along the sections, we critically discuss trends, research gaps, advantages, and limitations of rod models.
-The paper concludes with an overview of the main remarks (Sec.
+A structured review of rod-based models for continuum and soft robots.
+We organize rod models into nine deformation classes, providing a solid reference for the development of novel soft robot models (Sec.
 6).
-1.2 Related Surveys
+Deformation modes are the core behaviors that enable soft robots to move and interact with their environment.
+Classifying the literature by these modes highlights the mechanical effects of actuation, independently of the technology used, and clarifies the modeling and control approaches adopted.
+A comprehensive review of model-based control for continuum soft manipulators using rod models.
+We cover a range of approaches from foundational methods, such as inverse kinematics, inverse dynamics, and feedback linearization, to advanced techniques like model predictive control and optimal control (Sec.
+7).
+A review of learning-based controllers leveraging rod models. We emphasize Reinforcement Learning (RL) as a promising key enabler to achieve contact-rich manipulation tasks with CSM (Sec. 8).
+We provide our insights and critiques along the sections, and finally consolidate the discussion of the main trends, research gaps, and emerging challenges (Sec.
+9).
+While this review focuses on rod theories applied in modeling and control, we believe our insights could inspire new research directions in the wider field of soft robotics.
+The paper concludes with a summary of the lessons learned (Sec.
+10).
+2 Related Work
 Soft robotics has been the subject of extensive surveys, reviews, and perspectives, which represent a wealth of knowledge and a growing interest in a thriving research field.
-The seminal paper by Webster III and Jones (2010) reviews PCC kinematic models of continuum robots, emphasizing the robot-specific and robot-independent mappings and providing a perspective on further developments.
-Conversely, our work focuses on rod models, which generalize PCC approaches in case of significant external forces.
-Recently, Armanini et al.
-(2023) presented a structured overview of all modeling approaches in soft robotics.
-They propose a classification in four model classes (continuum mechanics, geometrical, discrete material, and surrogate), providing the fundamental theoretical grounds and analyzing their applicability.
+The seminal work of webster2010design reviews kinematic models of piecewise constant curvature (PCC) for continuum robots, highlighting both robot-specific and robot-independent mappings, and outlining directions for future development.
+In contrast, our work provides a comprehensive treatment of rod models that extend PCC formulations to scenarios involving significant interaction forces, such as during manipulation.
+Lately, gilbert2021mathematical provides a unified mechanics-based framework for modeling slender biomedical continuum robots, comparing Cosserat rod and general continuum formulations.
+It offers guidelines for selecting model DoFs to balance accuracy and computational cost, addressing challenges such as under-actuation, actuator–structure coupling, and contact-rich environments.
+Similarly, dupont2022continuum provides a unified overview on modeling and control of continuum robot architectures—including tendon-actuated, multibackbone, concentric tube, magnetic, and soft designs—focusing on their specific clinical applications while outlining future research directions for widespread clinical use.
+In parallel, zhang2022survey covers a long soft robotics pipeline, spanning design, actuation methods, modeling, and control.
+They focus on major technical concerns and provide perspectives on the development of continuum robots.
+Conversely, the perspective by mengaldo2022concise provides a unified formulation for modeling the embodied intelligence of soft robots.
+They review the theories explaining the practical significance of embodied intelligence.
+Conversely, our paper is a vertical investigation of rod theories for the modeling and control of continuum and soft robots.
+Another work reviews methods for soft robot modeling and control, emphasizing the use of real-time finite element method (FEM) (schegg2022review).
+Recently, sadati2023reduced focuses on model order reduction (MOR) techniques and reduced order models, clarifying their disentangled usage in soft robotics.
+Within this framework, they treat modeling approaches, constitutive laws, and solution strategies.
+While they survey the theoretical aspects of modeling, we also provide practical insights into the experimental validation of rod models.
+Finally, armanini2023soft presented a structured overview of all modeling approaches in soft robotics.
+They propose a classification into four model classes (continuum mechanics, geometrical, discrete material, and surrogate), analyzing the fundamental theoretical grounds and their applicability.
 However, the control problem is only touched marginally.
-The perspective by Mengaldo et al.
-(2022) provide a unified formulation for modeling the embodied intelligence of soft robots.
-It broadly overviews the theories explaining the practical significance of embodied intelligence.
-Conversely, this paper is a review/survey article specific to rod-based models and controllers.
-Another work reviews state-of-the-art methods for soft robot modeling and control, emphasizing the use of real-time FEM (Schegg and Duriez, 2022).
-Also, Zhang et al.
-(2022) broadly covers a long soft robotics pipeline: design, actuation methods, modeling, and control.
-It focuses on major technical concerns, providing a perspective on the development of continuum robots.
-Finally, Sadati et al.
-(2023) survey MOR techniques and ROM, clarifying their disentangled usage in soft robotics.
-Within this framework, they broadly overview modeling approaches, constitutive laws, and solution strategies.
-While they survey the theoretical aspects of modeling, we also provide insights into the experimental validation of rod models.
-Regarding control, George Thuruthel et al.
-(2018) provide a broad overview of model-based and model-free control strategies for continuum soft manipulators, without concern for the modeling.
-Also, Wang and Chortos (2022) provide a broad review of actuation mechanisms and control strategies, including open-loop, closed-loop, and autonomous methods.
-They discuss emergent directions in the control-actuator interface, underactuation, and utilization of artificial intelligence.
-Della Santina et al.
-(2023) describe the control problem of CSM specifically for the model-based view.
+Regarding control, george2018control provide a broad overview of model-based and model-free control strategies for CSM, however, without concern for the modeling.
+Building on this, wang2021survey provides a systematic review of Machine Learning (ML) methods for continuum robot control, highlighting model-free approaches for inverse kinematics and closed-loop control, and outlining trends to integrate data-driven methods to handle model uncertainties.
+Similarly, wang2022control reviews actuation mechanisms and controls, including open-loop, closed-loop, and autonomous methods.
+They discuss emergent directions in the control-actuator interface, underactuation, and the adoption of artificial intelligence.
+Subsequently, della2023model describe the problem of model-based control of CSM.
 They present a unified formulation of soft robot dynamics independent of the modeling technique.
-Then, they discuss the shape control and tracking problems and explore open challenges (e.g., under-actuation, environment interactions, actuator dynamics, task space control) while surveying the literature.
-Conversely, we review and discuss rod-based controllers utilizing model-based strategies and deep Reinforcement Learning (RL).
-In summary, this review on rod models complements related works with minor intersections.
+Then, they discuss shape control and tracking problems and explore open challenges, such as underactuation, environmental interaction, actuator dynamics, and task-space control, while surveying the literature.
+Finally, falotico2024learning explore how different models influence the design of learning-based controllers.
+In contrast to prior work, the present paper walks through the evolution of the literature of controllers for rod models, using both model-based strategies and deep RL.
+In summary, our review paper effectively complements related work with minor overlap.
+Indeed, most existing works mainly provide broad overviews of modeling techniques.
+In contrast, our vertical focus on rod theories aims to guide the reader from their mathematical formulations to rod-based models of continuum and soft robots, and further to recent control strategies, both model-based and learning-based.
+We aim to provide new perspectives by structuring the literature around deformation modes.
+This also serves as a practical reference for modeling continuum robots for control-oriented applications.
+3 Features of Soft Robot Models
+The soft robotics community proposed various techniques to describe continuum and soft robots.
+According to armanini2023soft, soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical (Figure 1).
+Clearly, each strategy has advantages and limitations that depend on the purpose and objectives.
+Also, some approaches could partially overlap or be combined to derive hybrid models (falotico2024learning).
+In the context of modeling and control, in this section, we answer the following question:
+What are the advantages and disadvantages of rod models with respect to other techniques?
+To this end, we provide a concise characterization of the features of the existing modeling techniques and their role in the control problem.
+In addition, we report the main features in Table 1.
+In summary, rod models strike a good balance between usability, efficiency, and accuracy.
+3.1 Machine Learning-based Models
+ML is a powerful data-driven approach to derive the forward model of CSM (laschi2023learning).
+Unlike analytical methods, which require explicit geometrical and physical descriptions of the robot mechanics, ML only requires training artificial neural networks on descriptive motion data to approximate the map between actuation and task space (chin2020machine).
+The advantages of this approach include independence from the robot, the computational efficiency of the forward model, and the potential implicit representation of complex physical phenomena, such as hysteresis (falotico2024learning).
+Indeed, learned forward models were effectively employed within control policies for tracking (centurelli2022closed; pique2022controlling) and throwing (bianchi2023softoss).
+However, they might become unsuitable for complex manipulation tasks due to the challenges of collecting and labeling representative interaction data (alessi2024pushing).
+Other limitations include the lack of interpretability of the forward model and the tendency to overfit the training data.
+3.2 Geometrical Models
+Geometrical models describe the deformed backbone of the CSM with a specific mathematical curve.
+This idea allows us to define the kinematics of CSM using a finite number of configuration variables that describe the curve and have clear geometrical meaning (e.g., curvature).
+The most widely used approach, the PCC, discretizes continuous bodies in circular arcs with constant curvature and null torsion (webster2010design).
+This model is viable for lightweight uniform CSM with actuators parallel to the centerline producing constant curvature.
+Also, refined methods proposed variable (mahl2014variable) or polynomial curvature (della2019control).
+Naturally, these methods allow for trade-offs between accuracy and efficiency, and their formulation eases the application of classical rigid-robot controllers.
+Indeed, geometrical models proved effective for various tracking tasks within model-based (fischer2022dynamic) and learning-based (lou2024controlling) controllers.
+A limitation of the geometrical assumption, however, is that it can degrade when soft robots are subject to significant external forces from interaction with unstructured environments.
+Unless augmented with other theories (xie2022simplified), they lack specific parameters to describe materials.
+| Features | ML | Geometrical | Discrete | Rod | FEM |
+| Computational Efficiency | ✓ | ✓ | ✗ | ||
+| Accurate for Complex Geometries | ✗ | ✓ | |||
+| Usable in Real-time Control | ✓ | ✓ | ✗ | ||
+| Data-driven Adaptability | ✓ | ||||
+| Handles Large Deformations | ✓ | ✓ | |||
+| Incorporates Material Models | ✗ | ✓ | ✓ | ||
+| Ease of Formulation/Implementation | ✓ | ✓ | ✓ | ✗ | |
+| Generalization to New Tasks | ✗ | ✓ | ✓ | ||
+| Model Interpretability | ✗ | ✓ | ✓ | ✓ | ✓ |
+| Handles Physical Interactions | ✓ | ✓ | |||
+| (✓) = Highly Suitable or Advantageous | |||||
+| () = Moderately Suitable or Partially Effective | |||||
+| (✗) = Less Suitable or Challenging | |||||
+3.3 Discrete Models
+The concept behind discrete models is to discretize the system body at the beginning of the modeling process. Two main methods under this category are the lumped-mass and the pseudo-rigid models.
+The lumped-mass method approximates the CSM as a set of lumped masses, springs, and dampers (habibi2020lumped).
+In this case, the configuration variables represent the spatial displacements of each lumped mass.
+The modularity of this approach can be adapted to represent arbitrarily complex phenomena (e.g., nonlinear friction) and morphologies (e.g., hybrid kinematic chains).
+On the other hand, to describe CSM with the same fidelity of continuum mechanics, they require a high number of lumped masses that entail computational effort in addition to data-intensive system identification.
+Nonetheless, the adoption of this method is facilitated by open-source simulators like SoMo (graule2021somo) and Titan (austin2020titan).
+Conversely, the pseudo-rigid approach represents CSM with a chain of rigid links connected by joints, finding an equivalent hyper-redundant rigid robot.
+Their formulation allows us to exploit the standard controllers from rigid robotics directly (bruno1994robotics).
+Therefore, pseudo-rigid models provide satisfactory results for hyper-redundant arms (venkiteswaran2019shape).
+However, they do not perform well in spatial accuracy and would require several DoFs to represent large deformations.
+Therefore, pseudo-rigid methods are less suitable for modeling the continuous elastic structures found in soft robotics.
+Nonetheless, morimoto2021model; morimoto2022characterization successfully adapted the MuJoCo simulator (todorov2012mujoco) to describe CSM and train learning-based controllers.
+3.4 Continuum Mechanical Models
+Finally, continuum mechanical models apply the principles of continuum mechanics to characterize soft robots with continuous configuration spaces and define deformations in physical terms.
+This formulation enables the simulation of physical interactions (e.g., contact with objects or fluids) and the in-depth study of robot mechanics, particularly for large deformations.
+These properties make them promising towards CSM interacting with unstructured environments.
+Depending on the dimensionality and approximation level, we can consider 3D models based on the FEM and low-dimensional formulations such as rod models.
+The FEM is one of the most widely used numerical approaches for solving PDE, particularly for modeling 3D elastic bodies.
+FEM discretizes the soft robot volume into a mesh, which is composed of interconnected finite elements (e.g., line segments, triangles, tetrahedra).
+The points defining these elements are called nodes, where physical quantities like displacement or stress are computed.
+Then, the continuous solution can be approximated by interpolating values between the nodes.
+This approach enables accurate simulation of the complex, nonlinear deformations to the extent of an involved mathematical formulation and high computational costs.
+Indeed, the solutions provided by standard commercial software (e.g., Abaqus, Ansys, COMSOL) would impede real-time control.
+Therefore, they were often limited to designing and simulating soft robotic components (xavier2021finite) or as a benchmark.
+Nonetheless, recent MOR techniques implemented in the SOFA simulator (allard2007sofa; coevoet2017software; dubied2022sim) are making 3D mechanical models more affordable even for learning-based control (agabiti2023whole; menager2023toward).
+When the robot is slender, it can be efficiently represented using elastic rod formulations, which approximate the body as a one-dimensional (1D) continuum.
+In particular, the Cosserat rod theory (CRT) can effectively describe all modes of deformations in CSM (renda2014dynamic; gazzola2018forward; till2019real).
+The CRT associates a reference system to each cross-section along the CSM.
+Each cross-section can translate and rotate relative to the adjacent cross-sections to simulate bending, twisting, shearing, and stretching.
+Thanks to the low-dimensional yet physics-based formulation, rod models retain all the benefits of continuum mechanics at a fraction of the computational cost of FEM.
+We should note that one limitation of rod models is that they are limited to describing inherently slender structures, although heterogeneous rod assemblies can describe complex systems (renda2018unified; zhang2019modeling).
+Beyond modeling, rod models are also gaining popularity in model-based control (caradonna2024model) and learning-based manipulation (alessi2024pushing).
+To sum up, rod theories are promising candidates for soft robotics modeling and control, which are the focus of the remainder of this paper.
 4 Background on Rod Theories
@@
 An elastic rod might exhibit (i) bending, (ii) torsion, (iii) stretching (i.e., elongation or compression), and (iv) shear.
-Fig.
-2 shows the four rod theories and the modeled strain modes.
-In this work, we adopted frequently the Lie Algebra notation.
-The reader can find in Appendix B the definition of the employed operators.
+Figure 2 shows the rod theories and the modeled strains.
+In this work, we frequently adopted the Lie Algebra notation.
+Its mathematical operators are defined in Appendix B.
 4.1 Problem Statement
-An Elastic Rod can be kinematically described through a time-variant 3D curve, usually called backbone.
+An Elastic Rod can be kinematically described through a time-variant curve, usually called backbone.
 The length of this curve can be parameterized by a material curvilinear abscissa .
@@
 Usually, in literature, the point of the backbone at is called base and at is called tip of the robot.
-Furthermore, every cross-section has associated geometrical and material quantities: the cross-section area , the second moment of area about the local axes ; the mass density of the rod , the Young Modulus of the material , and the shear Modulus , where is the Poisson coefficient.
+Every cross-section has associated geometrical and material quantities: the cross-section area , the second moment of area about the local axes , the mass density of the rod , the Young Modulus of the material , and the shear Modulus , where is the Poisson coefficient.
 4.2 Cosserat Rod Theory
-The Cosserat rod theory (CRT) (Cosserat and Cosserat, 1896) describes all strain modes of the rod.
+The CRT (cosserat1896) describes all strain modes of the rod.
 This model is useful for describing CSM that interact with cluttered environments, in which shear and elongation deformations play a fundamental role.
@@
 Definition (5) indicates the infinite DoFs of an elastic rod.
-The Configuration Space results in a functional space of curves in (Boyer et al., 2020).
+The Configuration Space results in a functional space of curves in (boyer2020dynamics).
 A Configuration Space with an infinite dimension is a unique feature of soft robots, making control very challenging.
@@
 The Actuation Forces Wrench expresses the internal active forces exerted by the actuators. Refer to Sec. 5 for more details.
-After defining these quantities, it is possible to compute the Dynamics of an elastic rod using the Poincaré equations (Renda et al., 2018a)
-The EoM (8) is a set of partial differential equations of a Cosserat Rod in the local frame.
-It is common also to find the same set of PDE in an explicit form and expressed in the inertial frame, such as
+For clarity, the distributed wrenches (, , ) represent the forces and torques acting on a particular cross-section (), rather than force or torque densities distributed along the body’s length.
+After defining these quantities, it is possible to compute the Dynamics of an elastic rod using the Poincaré equations (renda2018discrete)
+The EoM (8) is a set of PDE of a Cosserat Rod in the local frame. It is common also to find the same set of PDE in an explicit form and expressed in the inertial frame, such as (gazzola2018forward)
 where denotes a variable expressed in the inertial frame, is the second moment of the area expressed in the inertial frame, and and are the contributions of the internal and actuation forces.
@@
 From a control and numerical implementation perspective, it is convenient to find a method to discretize the continuous nature of the soft segment.
-Renda et al.
-(2020) and Boyer et al.
-(2020) proposed a Strain Parameterization, expressing the Kinematics and the Dynamics of the elastic rod in terms of strain twist .
+renda2020geometric and boyer2020dynamics proposed a Strain Parameterization, also known as Geometrical Variable Strain (GVS) mathew2025reduced, expressing the Kinematics and the Dynamics of the elastic rod in terms of strain twist .
 Let us again consider the definition of the strain twist (2), which is easily rewritten as .
@@
 This approach allows the user to choose the degree of approximation or neglect certain strain modes.
-To solve the PDE, the proposed method uses the Magnus Expansion (Haier et al., 2006), resulting in a convenient Product of Exponentials, which is widely used in classical robotics (Murray et al., 1994)
+To solve the PDE, the proposed method uses the Magnus Expansion (harier2006geometric), resulting in a convenient Product of Exponentials, which is widely used in classical robotics (murray)
 where denotes the Magnus expansion of the strain twist. Thanks to that, it is possible to rewrite the Differential Kinematics with a well-known form, such as
 where denotes the Soft Geometric Jacobian. Similarly, the Dynamics can be rewritten in a classical Lagrangian form
-The EoM (14) can be solved using a standard time solver, such as Runge-Kutta or Explicit Euler, as implemented in the SoRoSim simulator (Mathew et al., 2022).
+The EoM (14) can be solved using a standard time solver, such as Runge-Kutta or Explicit Euler, as implemented in the SoRoSim simulator (mathew2022sorosim).
+Recently, the Authors showed how to compute the analytical derivatives of the GVS approach mathew2024analytical, evolving SoRoSim as a differentiable simulator.
+In this work, the Authors also showed the benefits of the implicit time-integration schemes, such as the Implicit Euler and the Newmark- methods.
+Finally, to fully describe soft robots interacting with the environment, xun2024cosserat proposed an extension of (14), including frictional contacts with rigid and soft bodies.
 4.2.4 Discrete Elastic Rod
-Another method to discretize the continuum nature of elastic rods was introduced in the pioneering work of Gazzola et al.
-(2018).
+Another method to discretize the continuum nature of elastic rods was introduced in the pioneering work of gazzola2018forward.
 First, the Authors derived the EoM (9), including an elongation/compression ratio defined as , where is the curvilinear abscissa in the rest configuration.
@@
 Furthermore, scaling the geometrical quantities partially relaxes the assumption of rigid cross-sections.
-To numerically resolve the EoM (16), they extended for a Cosserat rod the spatial discretization algorithm proposed in a previous study (Bergou et al., 2008), which discretizes the rod in a sequence of rigid segments connecting nodes.
+To numerically resolve the EoM (16), they extended for a Cosserat rod the spatial discretization algorithm proposed in a previous study (bergou2008discrete), which discretizes the rod in a sequence of rigid segments connecting nodes.
 For each node, the EoM consider the interactions with the other nodes and external forces.
 In addition, it is possible to associate kinematic and dynamic quantities to each node and segment to solve (16).
-In the discrete domain, some quantities, such as curvature, must be expressed in an integrated form over the domain (Gazzola et al., 2018).
+In the discrete domain, some quantities, such as curvature, must be expressed in an integrated form over the domain (gazzola2018forward).
 This domain corresponds to the Voronoi region associated with the interior nodes ,
@@
 with . Finally, the discretized EoM of the rod can be rewritten in algorithmic form as follows:
-where is the point-wise mass associated with the node, is a domain dilation factor, and are the discrete difference and the averaging operator defined in (Gazzola et al., 2018).
+where is the point-wise mass associated with the node, is a domain dilation factor, and are the discrete difference and the averaging operator defined in (gazzola2018forward).
 In the first equation of (19), , and .
@@
 This distinction arises from the definition of discrete curvature and bending stiffness within the interior nodes set.
-For the time integration of the discretized EoM (19), the authors proposed the use of a Second-Order Position Verlet time integrator, which exhibits a good balance between numerical accuracy and computational cost (Gazzola et al., 2018).
-To ensure numerical stability, they proposed an empirical law to choose the integration frequency as .
-The discrete EoM (19) were implemented in the PyElastica simulator (Naughton et al., 2021).
+For the time integration of the discretized EoM (19), the authors proposed the use of a Second-Order Position Verlet time integrator, which exhibits a good balance between numerical accuracy and computational cost (gazzola2018forward).
+To ensure numerical stability, instead of enforcing rigorous Courant-Friedrichs-Levy stability conditions (courant1967partial), they proposed an empirical law to choose the integration time step as .
+The discrete EoM (19) were implemented in the PyElastica simulator (naughton2021elastica).
 4.3 Kirchhoff Rod Theory
-The Kirchhoff rod theory (KRT) (Love, 1906) is a special case of CRT that considers an elastic rod unstretchable and unshearable.
+The Kirchhoff rod theory (KRT) (Love1906) is a special case of CRT that considers an elastic rod unstretchable and unshearable o2017modeling; gazzola2018forward; coleman1993dynamics.
 It is particularly suitable for CSM that bend around any axis and twist.
@@
 In particular, the constraint of an unstretchable and unshearable rod can be written as
-This constraint can be added in the EoM (16), resulting in
+The constraint (20) enforces , which implies (gazzola2018forward). Incorporating this constraint into the EoM leads to the following modified equations of motion:
 The linear internal force serves as a Lagrangian multiplier and it is a virtual internal force that constrains the rod from stretching or shearing.
+As stated in gazzola2018forward, inserting the constraint of an unshearable and unstretchable rod can increase the computational time.
 In the case of Strain Parameterization, it is sufficient to apply the constraint (20) in the strain twist
 Therefore, the basis function matrix only generates the angular strain mode vector .
+Unlike the previous approach, (22) does not apply as a constraint, and the EoM can be solved without considering any Lagrangian multipliers.
 4.4 Euler-Bernoulli Rod Theory
-The Euler-Bernoulli rod theory (EBRT) (Timoshenko, 1983) is one of the simplest rod theories in which the rod can only bend around one axis.
+The Euler-Bernoulli rod theory (EBRT) (timoshenko1983history) is one of the simplest rod theories in which the rod can only bend around one axis.
 It can be considered the 2D case of the KRT without the twist.
@@
 where is the distributed external moments around the -axis. The EoM (24) is linear, assuming that the constitutive law of the bending moment is
-where is the -component of the internal moment . For control purposes, the linearity of (24) facilitates the use of efficient controllers from the classic control theory (Doroudchi et al., 2018).
+where is the -component of the internal moment . For control purposes, the linearity of (24) facilitates the use of efficient controllers from the classic control theory (doroudchi2018decentralized).
 4.4.2 Nonlinear Euler-Bernoulli
 In the nonlinear EBRT (or Euler’s Elastica), the rod length is parameterized by the arclength .
-Recalling the Fundamental Theorem of the local theory of Curves in the 2D case, every regular curve can be determined by the curvature (do Carmo, 1976).
-From this concept, Della Santina and Rus (2019) reformulated Elastica in a classical robotic formulation.
+Recalling the Fundamental Theorem of the local theory of Curves in the 2D case, every regular curve can be determined by the curvature (carmo_differential_geometry).
+From this concept, della2019control reformulated Elastica in a classical robotic formulation.
 Let be the -component of the curvature twist .
@@
 Similarly to the Strain Parameterization, they propose to consider the curvature as an infinite sum of monomials in , i.e.
-The geometrical meaning of the polynomial curvature is to constrain the shape of the backbone curve to be a Generalized Cornu Spiral, i.e., a curve with a polynomial curvature (Della Santina and Rus, 2019).
-It is worth highlighting that the (27) is equivalent to write (11) with a polynomial basis function, related only to the curvature (Caradonna et al., 2024).
+The geometrical meaning of the polynomial curvature is to constrain the shape of the backbone curve to be a Generalized Cornu Spiral, i.e., a curve with a polynomial curvature (della2019control).
+It is worth highlighting that the (27) is equivalent to write (11) with a polynomial basis function, related only to the curvature (caradonna2024model).
 Defining as joint variables, it is possible to write the EoM as
@@
 4.5 Timoshenko Rod Theory
-The Timoshenko-Ehrenfest rod theory (Timoshenko and Goodier, 1951) extends the EBRT, relaxing the unshearability constraint.
+The Timoshenko-Ehrenfest rod theory (Timoshenko1951) extends the EBRT, relaxing the unshearability constraint.
 Therefore, the equality between the tangent angle and the cross-sections’ angle is no longer valid (i.e., ).
@@
 Consider a Cosserat rod with actuators.
-Let be the distance from the center of the cross-section and the -th actuator (Fig.
-3).
+Let be the distance from the center of the cross-section and the -th actuator (Figure 3).
 Each actuator applies an internal active wrench that depends on the actuator routing.
 We can express the internal active wrench applied by the actuator as
-where is the tangent versor to the actuator path and is the magnitude of the i-th actuator. For example , can be
-where is the cable tension, is the pneumatic pressure and is the internal cross-section area of the fluidic chamber. Moreover, the expression of the tangent versor is derived as
+where is the tangent unit vector to the actuator path and is the magnitude of the i-th actuator. For example, can be
+where is the cable tension, is the pneumatic pressure and is the internal cross-section area of the fluidic chamber. The expression of the tangent unit vector is derived as
 where is expressed in homogeneous coordinates and the operator extracts the first three rows of a homogeneous vector. The resultant of the contributions of actuators is
@@
 From the definition of actuation matrix (35), it is possible to map the actuation magnitude vector to the distributed active load .
-Renda et al.
-(2020) propose a statics-based method to relate the actuators with excited strain modes.
+renda2020geometric propose a statics-based method to relate the actuators with excited strain modes.
 In particular, let be the statics of a Cosserat Rod as
@@
 Finally, substituting the constitutive relation (6) and the definition of actuation forces (34), we derive
-From this equality, Renda et al. (2024) propose a trivial form of the static equation, choosing the implicit parametrization
+From this equality, renda2024dynamics propose a trivial form of the static equation, choosing the implicit parametrization
 Consequentially, the static equation degenerates in a trivial form, such as
@@
 Furthermore, in (35) considers also the geometrical and the material information of the robot, contained in .
-Finally, (39) can be a useful starting point to build the optimal functional basis matrix , avoiding redundant or inefficient shape functions, preserving accuracy and computational efficiency.
+Finally, (39) provides a useful starting point for constructing the optimal functional basis matrix , avoiding redundant or inefficient shape functions while preserving accuracy and computational efficiency.
 6 Rod Models for Continuum and Soft Robots
-The rod models were classified based on combinations of the four deformation modes.
-Out of the fifteen possible combinations, only nine classes emerged.
-Fig.
-4 shows the distribution of rod models over deformations classes.
-Within each deformation class, we divided the papers mainly by actuation, robot type, or topic.
-In each division, the articles follow a chronological or logical order.
-An advanced Scopus query aided the selection of the articles (see Appendix A).
-Fig.
-5 shows how different robots might exhibit these deformations induced by actuators or the interaction with the environment.
-We survey rod models ordering the deformation classes from most to least popular (see Fig.
-4).
-We summarize the four major deformation classes (i.e., Bend, All Deformation, Bend & Twist, Bend & Stretch) respectively in Table 1-4 to facilitate the literature comprehension.
-Each table arbitrarily reports representative articles to distill the deformation class and highlight the versatility of rod models.
+We categorize rod models based on the deformation modes they represent, independently of the underlying actuation technology.
+From a control perspective, the specific actuation method (e.g., tendons, pneumatic chambers) is secondary to the type of strain it produces, such as elongation or contraction, torsion, or bending.
+The admissible deformation modes also implicitly determine the suitable rod theory.
+For example, a robot limited to planar bending is consistent with the Euler-Bernoulli framework, while more complex deformation behaviors might require Cosserat theory.
+While several studies formally employ the Cosserat rod formulation, in practice, they restrict the model to a limited subset of strain modes, effectively reducing it to simpler formulations.
+By explicitly classifying models according to the combination of strain modes they allow, we highlight such modeling choices.
+Out of the fifteen possible combinations of deformation modes, only nine classes emerged.
+We summarize the deformation classes in Table 2 to facilitate the literature comprehension, where we arbitrarily report representative articles to distill each class and highlight the versatility of rod models.
+An advanced Scopus query aided the initial selection of the articles (see Appendix A).
+We also discussed articles beyond the query.
 6.1 Bend
-Bending is a prevalent deformation thanks to the slender morphology of continuum and soft robots (Fig.
-5).
-Herein, we review the evolution of rod models undergoing bending through various mechanisms.
-Table 1 distills the deformation class reporting representative works.
-4.1.1 Bend by Cables
-The literature presents several solutions to predict the bending of soft robots.
-The CRT computed in real-time the kinematics of tendon-actuated robots subject to external loads , modeling the tendon actuation as a single point moment applied to the backbone where each tendon is attached, achieving position errors below on a Nitinol rod (Jones et al., 2009).
-Direct and inverse kinematic models were presented drawing inspiration from the structure of octopus arms (Giorelli et al., 2012).
-The direct kinematics exploited a Cosserat rod model.
-Conversely, the inverse kinematics computes the Jacobian matrix iteratively from the task space.
-The kinematic models overcome the slack problem as they work directly on the cable tensions instead of controlling the cable extension.
-Actuations validated the model on a prototype immersed in water.
-The challenge of predicting the bending angle of a planar manipulator driven by four wires and with triangular notches was addressed using a kinematic Timoshenko rod model (Wenlong et al., 2013).
-A two-node Timoshenko rod element modeled each triangular unit, accurately predicting the bending angles measured on the tensioned manipulator.
-Concerning planar grasping, it was proposed a dynamic Cosserat rod model for a tendon-driven continuum robotic finger (Dehghani and Moosavian, 2014b).
-The finger grasping is simulated as an external point force acting on the finger backbone , assuming that the objects do not slip.
-Also, a singularity-free version of the model equations is derived using the Taylor expansion.
-A preliminary experimental validation is conducted by attaching different weights on the tip of a superelastic NiTinol rod.
-Also, the EBRT was derived using the Lagrangian method to predict the shape of a multi-backbone continuum robot (He et al., 2013).
-The experimental validation obtained position errors lower than 4%L.
-Moreover, Gao et al.
-(2016) predicted manipulator deformations due to external forces, tendon forces, or friction.
-They provide a planar Cosserat rod model of dexterous continuum manipulators, achieving sub-millimeter tip error predictions in the experimental validation.
-A dynamic model of a cable-driven soft tentacle was developed to explore the relationship between the strain field and the rope length by formulating a rope-driven actuator model (Ma et al., 2021).
-A kinematic model based on CRT was presented.
-Then, a Newton-Euler inverse dynamics algorithm for an equivalent CSM computed the matrices of the Lagrangian inverse dynamics model.
-A computationally efficient analytical tip force estimator was proposed for tendon-driven catheters (Hooshiar et al., 2021).
-First, they used a Bezier-based shape interpolation assuming constant curvature.
-Then, they proposed an optimal problem to invert the system statics and compute the tip force .
-The experimental validation achieved a N mean absolute error at high frequency.
-Concerning the interaction with objects, a general CRT-based framework investigated the coupled dynamics of tendon-driven cooperative continuum robots interacting with flexible objects (Jalali and Janabi-Sharifi, 2021).
-The framework is validated experimentally using two cable-driven spring steel arms performing trajectories with a Nitinol object.
-Workspace investigations are crucial for dexterous manipulation tasks.
-A study on estimating the workspace of CSM was conducted by developing a quasi-static Cosserat rod model and introducing a continuation algorithm for a tendon-driven CSM (Amehri et al., 2021).
-The algorithm maps the exterior and interior workspace boundaries, addressing challenges like actuator saturation, physical limitations, self-collision, or elastic instabilities.
-The study highlights that complex configurations of CSM can lead to isolated boundaries and voids.
-Validation through numerical simulations on planar and spatial configurations demonstrated the effectiveness of the proposal.
-Other works focus on the stiffness analysis.
-A Cosserat rod model for a continuum manipulator with variable stiffness was used to compute the tip stiffness (Zhao et al., 2020a).
-Experimental results demonstrate the ability to adjust the tip stiffness in various directions with an increase of about ten times the minimum stiffness.
-Also, a novel cable-driven continuum robot with layer jamming sheaths, which control the stiffness in the transverse and axial directions, is designed, modeled, and validated (Fan et al., 2022).
-An Euler-Bernoulli rod and the Maxwell-Mohr method derived the effective robot stiffnesses, expressing the rod deflection as
-where is the moment due to the unit load.
-Extensive static force-deflection tests are conducted with different layer jamming sheaths, vacuum pressures, and bending angles.
-A novel approach proposed additional cables to selectively stiffen planar cable-driven robots with limited coupling with the actuating cables (Molaei et al., 2022).
-This approach increases the control of shape and stiffness of continuum manipulators without significantly affecting the basic kinematics and addressing uncontrolled deformations like buckling.
-A planar Cosserat rod model based on (Rucker and Webster III, 2011) identified the design conditions to route the additional cables.
-The simulations show a contained coupling between the length changes of the actuating and stiffening cables.
-The proposal is validated experimentally on a prototype with two bending cables and two stiffening cables, demonstrating that the secondary tendons increase the stiffness by ten times.
-Stiffness is also investigated for a simple rod-shaped silicone soft robot modeled with a geometrically exact CRT (Grube and Seifried, 2022).
-The stiffness properties are investigated by applying horizontal pulling forces up to 3 N to the rod and robot free-ends.
-They analyzed experimentally a linear viscoelastic material model and the nonlinear Saint Venant-Kirchhoff material model.
-The analysis of the damping mechanism has shown that linear damping can be assumed for oscillations with small to medium amplitudes since nonlinear material laws improve the model accuracy only slightly.
-Therefore, nonlinear effects have to be considered for oscillations of higher amplitudes.
-Xiao et al.
-(2023) present an on-demand stiffness modulation approach that retains the intrinsic compliance of elastomeric bodies.
-They use a Nitinol tube as the backbone, which slides through a cable-driven soft body to achieve robot pose or stiffness modulation.
-They develop and validate a CRT model in different scenarios by varying and the contact force .
-Experimental results indicate average shape and tip errors below 1%L, while simulations demonstrate that backbone insertion enhances the workspace and reduces compliance by 57.7%.
-Their design was further validated by object manipulation and minimally invasive surgery (MIS) case studies.
-A generalized kinetostatics CRT model with friction considered for multi-module and multi-segment continuum robots is proposed (Wang et al., 2023).
-The model is applied to a two-module rod-driven continuum robot with winding ropes to analyze its deformation and load characteristics.
-The experimental validation, with and without loads, achieves less than shape error.
-4.1.2 Bend in Soft Pneumatic Actuators
-Several works estimate the bending deflection in pneumatic soft robots.
-An Euler-Bernoulli rod described a bending pneumatic soft actuator (de Payrebrune and O’Reilly, 2016).
-Through experimental procedures with the robot clamped at one end, the curvature was expressed as a function of the applied pressure .
-The results show that the curvature field is a function of the applied pressure.
-FEM simulations often provide a benchmark.
-Using FEM and experiments, a five-parameter constitutive relation for a rod-based pneumatic soft arm model was explored (De Payrebrune and O’Reilly, 2017).
-The constitutive law’s parameters are determined for pneumatic soft arms through FEM simulations under different conditions.
-Furthermore, various actuator sections are investigated (e.g., rectangular, semicircular).
-The parameterization is validated by comparing FEM with the rod model for various actuator sections, subject to tip loads .
-Another work introduced interoperable dynamic modeling for fluidic elastomeric soft fingers (Mbakop et al., 2021b).
-The model employs an Euler-Bernoulli framework with Pythagorean Hodograph (PH) curves reconstructing the kinematic shape from virtual control points.
-The shape of the soft finger is recovered as a function of the applied pressure and control points as
-Here, is the height from the ventral part of the finger in the cross-section, and is the normalized curvilinear coordinate along the neutral axis of the soft finger.
-In this way, a relationship is established between control inputs and PH curve control points, enabling dynamic inversion.
-FEM simulations validated the PH-EBRT approach for single and two-segment soft fingers, while the experimental validation measures cartesian error and bending curvature on 3D-printed soft fingers of varying stiffness.
-Interestingly, EBRT successfully modeled a compliant robotic fish tail with fluidic actuators (Wu and Shi, 2022b).
-The model predicts actuator deflections and nonlinear strains.
-Nonlinear FEM validated the approach by comparing the results in terms of the first, second, and third obtained natural frequencies, showing good agreement.
-Finally, the bending of a beam-shaped soft actuator with an embedded fluidic network was analyzed by leveraging a dynamic Euler-Bernoulli rod with large deflections and nonlinear strains (Janizadeh Haji and Bamdad, 2023).
-Its accuracy is compared against nonlinear FEM and 4th Runge–Kutta.
-Examining multi-mode and nonlinear strain effects revealed instabilities in some frequencies for the nonlinear model.
-Geometric parameters influence actuator deformations, with increased channel distribution density enhancing deflection.
-Moreover, larger fluid channel diameter and actuator length require fewer inlet pressures for the same bending, while higher actuator width and height demand higher pressures.
-Conversely, Sadati et al.
-(2019) introduced a ROM and a discretized model based on EBRT with absolute states for the STIFF-FLOP.
-In addition, they improve the model in (Renda et al., 2018a) with an Euler-Bernoulli discretized model with relative states.
-An experimental comparison revealed that the ROM is more accurate and numerically robust, while absolute states provide higher computational efficiency.
-In contrast, relative states show higher sensitivity and computational burden.
-Finally, a lumped-mass model provided a benchmark with the highest computational cost among the considered models.
-Bending structures also exhibit stiffness modulation and adaptation.
-For instance, CRT modeled a soft bending actuator with an added fabric layer, which changes the actuator stiffness, enhancing the bending and preventing longitudinal stretching (Nalkenani et al., 2021).
-The authors incorporate fabric effects as a boundary condition in the Cosserat model and validate experimentally across actuation frequencies and loading conditions, producing accurate behavior predictions.
-Also, it was proposed a nonlinear kinematic Cosserat rod model of a pneumatic surgical CSM with stiffness adaptation (Roshanfar et al., 2022).
-The model incorporates nonlinear material properties through the two-term Mooney-Rivlin hyperelastic constitutive law, whose strain energy density is
-Here, and are the deviatoric strains invariants while and are material constants.
-The simulation and experimental results for tip displacement and stiffness show similar trends with a maximum error of 8.25%.
-More recently, the work in (Namdar Ghalati et al., 2022) presented a method to obtain the static shape of a semicylindrical soft fiber-reinforced pneumatic bending actuator under external force constraints.
-The pressure-angle relationship in free motion is derived using the Neo-Hookean strain energy theory of hyper-elastic materials, i.e.,
-Here, is the first invariant of the right Cauchy-Green deformation tensor and , , are the principal stretches (Odgen et al., 2013).
-They investigated the effects of external forces using the EBRT, which modeled the arm as a cantilever beam under external load.
-4.1.3 Bending with Smart Actuators
-Bending is achieved by several innovative actuation technologies.
-Inspired by the forward crawling motion of the Manduca sexta caterpillar, a dynamic Euler-Bernoulli rod model for a soft robot with shape memory alloy (SMA) actuators was introduced to analyze undulation mechanics (Daily-Diamond et al., 2017).
-The model employs a motion pattern based on the Witch of Agnesi function
-Its shape is analogous to the probability density function of the Cauchy distribution.
-Lastly, experimental data on an SMA actuator’s flexural rigidity, curvature, and recorded caterpillar leg velocity enable studying ground contact forces and the energetics of successful forward locomotion.
-Differently, CRT solved direct and inverse kinematics of soft robots actuated by twisted and coiled polymers (Pawlowski and Zhao, 2018).
-Direct kinematics determines the manipulator shape based on TCP temperatures, and inverse kinematics uses an augmented Jacobian approach (Rucker and Webster, 2011).
-A subsequent work integrates a thermodynamic TCP model and a dynamic Cosserat rod model, with promising forward dynamics results (Pawlowski et al., 2018).
-Additionally, an Euler-Bernoulli rod model predicted the deflection of a micro soft robot made from a magnetic elastic composite material (Xiang et al., 2019).
-The 3D-printed robot, having a head and a tail for rolling and crawling, is controlled by two magnetic fields generated by a 3D Helmholtz coil.
-Experimental validation for rolling and crawling velocity under different input waveforms, amplitudes, and frequencies shows a linear relation of rolling speed to rotating magnetic field frequency.
-Moreover, crawling speed correlates positively with magnetic field amplitude and frequency, being faster with sinusoidal compared to triangular wave inputs.
-A recent work developed a dynamic model of a continuum robotic arm with a 3D-printed elastic backbone and orthogonal plates supporting distributed soft cubic actuators, composed of a novel temperature-responsive hydrogel, each with an embedded Joule heater (Doroudchi et al., 2020).
-Leveraging the CRT, they simplify the PDE into reduced ordinary differential equations by integrating semi-discretized distributed force and rotational inertia components.
-Assuming robot isotropy and homogeneity, they estimate the Young modulus and damping coefficients through vibrational tests.
-Experimental verification accurately predicts tip displacements with a maximum of error.
-Finally, a study proposes a model of CSM with magnetic actuation (Tariverdi et al., 2020).
-They use Lie group synchronous variational integrators to derive a geometrically exact model based on an extended dynamic Kirchhoff rod and Hamiltonian formulation.
-The model considers nonlinearities, magnetic field loads, gravity, and internal and external dissipation forces from friction.
-Experiments involve dynamic motions with metallic and polymeric soft rods.
-Polymer rods show higher errors due to fabrication uncertainties and nonlinear elastic properties.
-Increasing the number of nodes improves the system frequency accuracy.
-4.1.4 Bending in Parallel Manipulators
-A novel parallel continuum manipulator (PCM) constructed with soft panels and its shape modeling approach was presented (Li et al., 2018).
-They used the nonlinear EBRT to model the deformations of each soft panel.
-By solving a constrained optimization for the kinematics, they calculated the deformed shapes and obtained the relationship between the panels’ length and the pose of the moving platform.
-Experiments validated the model reaching good shape predictions.
-Li et al.
-(2023) present a kinematic model of a planar PCM made of two independent flexible panels.
-The formulation unifies the geometric constraints, the load balance of the moving platform, and the panel’s statics.
-For each panel, the relation between internal moment and deflection is described by the EBRT via elliptic integrals, which account for inflection points with large deformations.
-Simulation results compared against FEM reveal that, under large deflections, the panels exhibit two inflection points close to both ends, which affect the distribution of the internal moment along the panel, (25).
-Stress checking confirms that the deformations are within the elastic region.
-4.1.5 Bending in Tubular Robots
-Bending is the prominent deformation of tubular robots aimed for MIS.
-Pattanshetti and Ryu (2019) employ the CRT to model a laser-patterned monolithic tube with varying structural properties along the length.
-The combination of CRT and a vision-based identification algorithm enables the characterization of the non-linearities and the bending stiffness along the rod with increasing strain.
-Another work adjusts the piecewise variable-strain (Renda et al., 2020; Boyer et al., 2020) to simulate multi-section concentric tube robots by including the tubes sliding motion (Renda et al., 2021).
-Rotation motions of the tubes are included as generalized coordinates instead of boundary kinematic conditions.
-This procedure led to a minimum set of closed-form algebraic equations solvable for both the shape variables, the actuation forces, and the torques.
-The resulting equations facilitate control, design optimization, and stability assessment.
-The work tackles one of the main limitations of the sliding structure of standard CTR, the spaghetti problem.
-4.1.6 Bending in Simulated Locomotion
-Bending is also exploited to investigate in simulation locomotion strategies (Fig.
-5).
-Zhou et al.
-(2015b) describe a soft robotic limb attached to a mass with an Euler-Bernoulli rod to examine in quasi-static conditions the peristaltic locomotion controlling the intrinsic curvature on a rough horizontal surface.
-The paper also provides design recommendations for the optimal performance of soft robots with variable curvature and adhesion.
-Interestingly, they also studied grasping strategies using a gripper with two fingers attached to a palm modeled using two rods (Zhou et al., 2015a).
-Concerning snake-like locomotion, especially undulatory swimming, is examined by studying a planar inextensible elastic Cosserat rod with adjustable spontaneous curvature, providing an internal actuation mechanism resembling muscular action in a snake (Cicconofri and DeSimone, 2015).
-Equations of motion are derived for (a) the rod moving along a prescribed curve and (b) the rod sliding longitudinally without fixed lateral slipping.
-Non-standard boundary conditions arising from constraints enable solving the equations.
-The one-dimensional equation captures system kinematics and dynamics without restrictive assumptions on trajectory or actuation, emphasizing the role of spontaneous curvature in driving and steering locomotion.
-4.1.7 Bend Miscellaneous
-Rods might exhibit interesting and mutable behaviors.
-Armanini et al.
-(2017) studied a soft robotic arm clamped at its base with a revolute joint, discovering that when the load is smaller than the buckling critical load, the rod behaves as an elastica compass.
-The closed curve can be computed analytically by the EBRT.
-Conversely, when the load exceeds the critical buckling value, the rod behaves like an elastica catapult, causing the snap phenomenon, where the rod releases very quickly elastic potential energy.
-They computed the dynamic motion using a simplified massless rod model and FEM and validated the behaviors by experimenting with an elastic rod.
-Instead, Cacace et al.
-(2020) present a model for a simulated octopus tentacle based on the dynamics of an inextensible string with curvature constraints and curvature controls.
-They derive the EoM and appropriate boundary conditions characterizing the corresponding equilibria.
-The model gives fourth-order evolutive nonlinear controlled PDE, generalizing the classic Euler’s dynamic Elastica equation.
-They explore the reachability of the model and introduce and compare two optimal control algorithms based on the augmented Lagrangian method solved with adjoint-based gradient descent.
-Subclass | References | Rod Model | Contribution |
-Cable | Jones et al. (2009) | CRT | Static model for real-time shape estimation of CSM. |
-Dehghani and Moosavian (2014b) | CRT | Dynamic model of a continuum finger for planar grasping. | |
-Gao et al. (2016) | CRT | Model of notched continuum manipulator for MIS. | |
-Pneumatic | De Payrebrune and O’Reilly (2017) | EBRT | Derivation of constitutive law of a planar bending actuator. |
-Sadati et al. (2019) | EBRT | ROM and rod models with absolute and relative states validated on STIFF-FLOP. | |
-Smart | Daily-Diamond et al. (2017) | EBRT | Dynamic model inspired by caterpillar motion to analyse ondulation mechanics. |
-Pawlowski and Zhao (2018) | CRT | Modeling the coupling between CSM body and thermal actuation. | |
-Parallel | Li et al. (2018) | EBRT | Shape modeling of a PCM made of soft panels. |
-Tubular | Pattanshetti and Ryu (2019) | CRT | Estimation of stiffness variation on a monolithic tube for MIS. |
+Bending is a prevalent deformation in continuum and soft robots due to the natural propensity of these slender structures to bend (Figure 4).
+Various rod models were employed to predict the bending motion under different actuation mechanisms, stiffness modulation strategies, or dynamic interactions with the environment.
+To clarify the different modeling approaches, we distinguish between planar (2D) bending and spatial (3D) bending.
+6.1.1 Planar bending (2D)
+Planar bending assumes that deformation occurs in a single plane, which significantly simplifies the mathematical formulation.
+Among the approaches, the EBRT is particularly suitable for modeling simple soft robots that bend due to its simplicity.
+Indeed, EBRT modeled various robotic architectures, achieving sub-millimeter tip position accuracy in multi-backbone continuum robots (he2013analytic) and enabling shape modeling in parallel continuum manipulators via its non-linear formulation (li2018shape).
+The Timoshenko rod theory, which is suitable for planar deformations (Figure 2), was used to model planar bending in cable-driven soft robots with complex geometries, such as triangular notched structures (wenlong2013mechanics).
+Pneumatically actuated soft robots also benefited from rod models.
+Among these, EBRT effectively captured the relationship between internal pressure and bending deformation (de2016constitutive).
+Later, this approach was extended to handle large deflections and non-linear strains (janizadeh2023steady; wu2022design).
+sadati2019reduced introduced a ROM and a discretized formulation of the STIFF-FLOP robot using EBRT, demonstrating high accuracy, numerical robustness, and near real-time performances.
+Further extending EBRT applications, li2023analytic developed a kinematic model for a planar PCM with two independent flexible panels.
+This formulation integrates geometric constraints, load balance considerations for the moving platform, and panel statics.
+Beyond kinematic descriptions, gao2016mechanical developed a planar Cosserat rod model for dexterous continuum manipulators, incorporating external forces, tendon forces, and friction, and achieving sub-millimeter tip error predictions in experimental validation.
+Dynamic planar bending has also been addressed through variational approaches.
+In particular, tariverdi2020dynamic proposed a Lie group variational integration framework experimentally validated on magnetically actuated soft continuum manipulators, achieving sub-millimeter tip accuracy.
+Planar bending models were also applied to investigate locomotion strategies.
+For instance, zhou2015flexing modeled peristaltic locomotion with an Euler–Bernoulli rod, and cicconofri2015study used a planar Cosserat rod with adjustable curvature to reproduce snake-like swimming.
+As a further extension, daily2017dynamical developed a dynamic Euler–Bernoulli model of a caterpillar-inspired robot with shape memory alloy (SMA) actuators, linking undulation mechanics to ground contact forces.
+Similarly, xiang2019study applied an Euler–Bernoulli formulation to predict rolling and crawling in a magnetically actuated soft robot.
+These works prove the versatility of rod models beyond manipulation.
+6.1.2 Spatial bending (3D)
+Spatial bending occurs when deformation is not confined to a single plane but involves curvature around multiple axes, eventually accompanied by torsion.
+These cases typically require Cosserat or Timoshenko formulations to accurately capture the complete range of strain modes.
+CRT, despite its higher complexity, has also extensively contributed to bending modeling in cable-driven robots (jones2009three_dim).
+An early example of CRT-based bending modeling is provided by giorelli2012two, who developed a kinematic model for a cable-driven soft arm with non-uniform cross-sections and derived its inverse model using the Jacobian matrix.
+In pneumatic actuators, CRT was coupled with nonlinear constitutive laws (e.g., Mooney–Rivlin hyperelasticity) to improve accuracy (Roshanfar2023Hyperelastic).
+Similarly, namdar2022static computed the static shape of a semicylindrical soft fiber-reinforced pneumatic bending actuator under external force constraints, using a Neo-Hookean model to derive the pressure-angle relationship in free motion (ogden2013).
+CRT was also applied to soft robots embedding smart actuators (Pawlowski2018modeling; Pawlowski2018dynam_model; doroudchi2020dynamic).
+Beyond kinematic descriptions, dynamic models mainly based on CRT, provide more suitability for dynamic tasks and allow the integration of actuation into its formulation (Ma2021dynamicsModel).
+Further refinements in CRT-based cable-driven robot models included friction effects arising from interactions between cables and the robot structure (wang2023general).
+CRT was also leveraged to model tasks involving interaction with external objects.
+jalali2021dynamic introduced a CRT-based framework to analyze the coupled dynamics of tendon-driven cooperative continuum robots interacting with flexible objects.
+Likewise, dehghani2014dynamics proposed a dynamic Cosserat rod model for a tendon-driven continuum robotic finger, where the grasping forces are modeled as external forces acting on the backbone.
+These studies highlight the effectiveness of CRT in dynamic interaction tasks, including force estimation in cable-driven catheters, as demonstrated by hooshiar2021analytical, who used a Bézier-based shape interpolation and an optimization approach to estimate tip forces with high precision.
+Another important aspect of dynamic interactions concerns workspace estimation, namely the determination of the spatial region that the soft robot’s end-effector is able to reach given its actuation and mechanical constraints.
+amehri2021workspace developed a quasi-static CRT model with a continuation algorithm to map workspace boundaries in tendon-driven manipulators, addressing self-collision, elastic instabilities, and actuator saturation.
+The study found that complex CSM configurations could lead to isolated boundaries and voids, which are essential to consider for accurate workspace predictions.
+Beyond kinematic and dynamic predictions, stiffness modulation also plays an important role in predicting soft robots behavior (xiao2023kinematics) as well as mitigating undesired effects such as buckling, which can hinder precise movements (molaei2022cable).
+Notable advancements involve variable stiffness continuum manipulators, which, leveraging CRT-based modeling can increase stiffness up to ten times, resulting in improved control precision (zhao2020sim).
+Other techniques include the addition of layer jamming sheaths to control stiffness both along the axial and transverse directions (fan2022novel).
+In parallel, research on a simple rod-shaped silicone soft robot, modeled using CRT, investigated stiffness properties by applying horizontal pulling forces, up to 3 N, at the free ends of the rod (grube2022simulation).
+Two material models were examined: a linear viscoelastic model and a nonlinear Saint Venant-Kirchhoff material model.
+The damping mechanism was analyzed, revealing that linear damping can be assumed for small to medium oscillations.
+However, non-linear effects become relevant when oscillations reach higher amplitudes, enhancing the model’s accuracy.
+Nalkenani2021modeling presented a CRT model of a soft bending actuator with an additional fabric layer to modify the actuator stiffness, modeled as boundary condition of the Cosserat model.
+This adjustment improved bending behavior while preventing longitudinal stretching.
+FEM was widely used to augment and validate rod-based models, particularly for pneumatic actuators.
+For instance, de2017development proposed a five-parameter constitutive relation for a pneumatic soft arm, with parameters identified via FEM simulations under different loading conditions.
+Conversely, mbakop2021inverse integrated Pythagorean Hodograph (PH) curves with an Euler–Bernoulli framework to reconstruct actuator shapes from virtual control points, establishing a direct link between control inputs and geometry.
+Both approaches, as well as more recent formulations (li2023analytic), were validated through FEM and experimental results, confirming their accuracy.
+In parallel, rone2013continuum developed a high-fidelity lumped-parameter dynamic model for cable-driven continuum robots, based on the principle of virtual power.
+It captured curvature variations while accounting for inertial, actuation, friction, elastic, and gravitational effects.
 6.2 All Deformations
-The second most popular deformation class concerns models accounting for all the primary deformations of continuum and soft robots.
-Below, we review the literature on rod models exhibiting all strains in cable-driven, pneumatic, parallel continuum, and biomechanical structures.
-Table 2 summarizes the class with the notable works.
-4.2.1 Deformations by Cables
-Renda et al.
-(2012) present a geometrically exact static Cosserat model of a conical tendon-driven CSM inspired by the octopus.
-Cable tension is modeled as a pure tension element, neglecting friction between the cable and the silicone body.
-After a FEM validation of the algebraic equations, the Cosserat model is validated experimentally for a static planar motion.
-The following work investigated the dynamic interaction of the silicone arm with a dense medium and the coupled tendon condition, addressing the external forces such as gravity, buoyancy, drag, added mass, and cable load (Renda et al., 2014).
-The dynamic experimental validation concerned the replication of characteristic octopus movements.
-In addition, an ablation study suggests that minimal internal viscosity guarantees solution stability and minimizes computational costs.
-Then, Renda et al.
-(2018b) introduced a unified dynamic Cosserat rod framework for underwater soft robots, which considers the nonlinearities of soft bodies, offering generality that extends to shells.
-The framework, which could aid the evaluation of potential soft robot designs, was validated experimentally using an autonomous vehicle with four cable-driven octopus-inspired CSM.
-At the same time, Renda et al.
-(2018a) presented an approach called piecewise constant strain (PCS) that discretizes continuous Cosserat rods in pieces, assuming constant strain for each piece.
-The PCS generalizes the PCC (Webster III and Jones, 2010) providing (i) numerical robustness given by the exponential formulation, (ii) inclusion of shear and torsion due to the CRT, and (iii) possible extension to rigid kinematic chains.
-The PCS is validated in real-world experiments with a cable-driven siliconic CSM.
-Concerning discretization approaches, it was proposed a real-time discretization technique to solve the PDE of a Cosserat rod (Till et al., 2019).
-The method discretizes the time derivatives and solves the ODEs for the boundary value problem (BVP) in at each timestep.
-The authors validate the efficient method by comparing the simulation to high-speed footage of a cantilever rod.
-For MIS applications, (Ibister et al., 2021) presents a dynamic CRT model for a continuum robot composed of a backbone and disks to support actuating tendons.
-The model includes the gravity (7) and tip wrenches produced by the tendons.
-The comparison between simulation and experiments shows a mean tip position error.
-Recently, a cable-driven continuum robotic arm was integrated within an aerial manipulation system (Samadikhoshkho et al., 2020).
-The arm dynamics are modeled with a Cosserat rod with a floating base , incorporating both the actuation and quadrotor effects.
-They proposed a robust adaptive position controller and proved its stability using the direct Lyapunov stability theorem.
-The scheme is validated in simulation for quasi-static, passive, and active arm conditions.
-4.2.2 Deformations by Pneumatic Actuators
-Trivedi et al.
-(2007, 2008b) present the first dynamic Cosserat rod model for pneumatic CSM based on a fiber-reinforced model of the air muscle actuators.
-The model includes the effect of material nonlinearities, distributed forces, and payloads and is geometrically exact for large curvature, shear, torsion, and extension.
-The model is experimentally validated statically on the OctArm V for planar movements with vertical and horizontal base orientation, achieving results about ten times more accurate than the constant curvature model.
-Subsequently, the model is used within a model-based design optimization study for pneumatic and hydraulic CSM (Trivedi et al., 2008a).
-Recently, the PCS discretization was combined with Screw Theory, providing a new model of actuation sources for pneumatic actuators (Renda et al., 2017).
-This approach enables the routing of fluidic actuators (even tendons) along any path in the soft arm.
-They define a soft geometric Jacobian for the PCS formulation and a technique to identify the active DoFs, applying it to the study of manipulable and force-closure conditions.
-Finally, they validate the theory on a STIFF-FLOP arm (Cianchetti et al., 2013).
-Additionally, Roshanfar et al.
-(2021) described a pneumatic soft robot with a Cosserat rod model, which was solved as an initial value problem with homogeneous Neumann and Dirichlet boundary conditions for a given tip force.
-The experimental validation measured a tip position error of .
-Analytical solutions and detailed FEM simulations are useful means for comparison and validation.
-For example, analytic solutions for the statics of a cantilever quasi-static Cosserat rod under external loads are derived (Grazioso et al., 2018).
-The position and orientation fields of the arm are defined on a Lie group, while the velocity and deformation fields are obtained from the time and space derivatives, respectively.
-A constitutive law between internal forces and strains enables the definition of the arm’s internal strain energy.
-Therefore, the principle of virtual work provides the static equilibrium equations.
-Finally, quasi-static rod configurations are compared to analytical solutions for pure bending and shear loads.
-Conversely, a comparative study for a Cosserat model and a FEM model of a soft pneumatic actuator was presented (Berthold et al., 2021), providing a preliminary parameter identification process.
-The FEM model was also validated experimentally.
-Moreover, an operational point-dependent static Cosserat model is obtained from a detailed volume FEM of a fiber-reinforced pneumatic silicone soft robot (Wiese et al., 2022).
-FEM simulations determine the actuation-dependent curvature, while the rod stiffness properties are identified by applying external forces to the FEM.
-The validation against FEM and experimental data, under various conditions with and without loads, show that characterizing a single module is sufficient to predict the behavior of a two-module version of the soft robot.
-To tackle large deformations of pneumatic soft robots, Mishra et al.
-(2022) suggest using an explicit fractional-order Bouc-Wen model that considers material nonlinearities, such as hysteresis.
-The constitutive law incorporated in a Cosserat rod and tested on a Robotino-XT demonstrated superior accuracy to other models (Wang et al., 2022a).
-4.2.3 Deformations with Parallel Manipulators
-Till and Rucker (2017) addresses the problem of elastic stability in PCM, introducing a condition to evaluate the elastic stability of PCM solutions based on Kirchhoff and Cosserat rods.
-A numerical test evaluates this condition and a heuristic stability metric.
-After verifying that the stability test matches other classical results, they experimentally validate the approach on a six-link PCM.
-The same authors propose a generalized kinetostatics Cosserat rod model (Black et al., 2017).
-Model linearization provides the manipulator Jacobian, end-effector compliance, input stiffness, and wrench reflectivity matrices.
-This enabled the analysis of the PCM manipulability and force-sensing accuracy.
-Using a similar PCM, they achieve accurate actuation-based force sensing and low orientation error.
-Conversely, Armanini et al.
-(2021) extended the PCS parametrization to soft PCM.
-They present a mathematical tool for modeling closed-chain geometries composed of soft and rigid elements.
-The formulation is applied to modeling and analyzing the Fin-Ray finger, especially for studying bending in the opposite direction of the external force.
-Furthermore, this formulation allowed the development of a novel design of the Fin-Ray finger, finding the optimal number of rigid ribs and their optimal orientation in terms of maximum payload and bending capability.
-The elegant work of Lilge and Burgner-Kahrs (2022) present a CRT-based kinetostatic modeling framework for tendon-driven PCM, which enabled to efficiently solve the forward, inverse, and velocity kinetostatic problems.
-In addition, it facilitates the study of kinematic properties like reachability, singularities, manipulability, and compliance.
-Experiments on a PCM made of four tendon-driven backbones attached via spherical joints to a platform achieved median pose accuracy of and .
-4.2.4 Deformations in Bioinspired Modeling
-Cosserat rods can be useful tools for biomechanical studies.
-With application to flapping flight and passive swimming, Boyer et al.
-(2017) provide a general unified framework for modeling bio-inspired locomotion robots using soft organs.
-They model soft parts in two ways: (i) the floating frame approach and (ii) the geometrically exact approach.
-They implemented the two models in a unified general algorithm based on recursive Newton-Euler.
-The comparison of the two approaches revealed that the former needs fewer coordinates and stronger deformation approximations.
-Conversely, the latter is more accurate and manages better the external load.
-The validation of the unified framework simulated two bio-inspired examples of soft locomotion.
-Recently, Zhang et al.
-(2019) realistically simulates musculoskeletal architectures in PyElastica.
-They assembled heterogeneous, active, and passive Cosserat rods into dynamic structures that mimic bones, tendons, ligaments, fibers, and muscle connectivity.
-The simulations involve biological and soft robotic case studies, including the human elbow joint, the design and fabrication of swimming and walking millimeter-long bio-hybrid bots, the synthesis of slithering snakes with intricate musculature layouts, and the replication of feathered bird wings - using up to rods.
-In these applications, the CRT demonstrated versatility to advance forward and inverse bioengineering designs and fundamental discoveries in the functioning of living organisms.
-Boyer et al.
-(2022) investigates the relationship between CRT and optimal control theory, deriving the statics and dynamics of a continuum robot as solutions to a minimization problem.
-With this different formulation, the authors revealed significant theoretical insights, such as the intrinsic singularity of the dynamic simulation of continuum robots.
-Furthermore, the iterative solution of the optimal control problem corresponds to the continuous version of the Newton-Euler algorithm for rigid kinematic chains.
-Finally, the presented approach is extended to locomotors and applied to a bio-inspired continuum swimmer, validating the theoretical work through simulations.
-Subclass | References | Rod Model | Contribution |
-Cable | Renda et al. (2014) | CRT | Studied the dynamic interaction of an octopus-inspired CSM with a dense medium. |
-Pneumatic | Trivedi et al. (2008b) | CRT | Model validation of CSM with fiber-reinforced actuators varying base orientation. |
-Parallel | Till and Rucker (2017) | CRT | Studied and experimentally validated the elastic stability of PCM. |
-Armanini et al. (2021) | CRT | Modeled closed-chain soft PCM with validation on a soft-rigid finger. | |
-Lilge and Burgner-Kahrs (2022) | CRT | Studied reachability, singularities, manipulability, and compliance of PCM. | |
-Bioinspired | Zhang et al. (2019) | CRT | Simulates musculoskeletal systems assembling heterogeneous active/passive rods. |
-Boyer et al. (2022) | CRT | Studied the link between CRT and optimal control simulating locomotors and swimmers. |
+With the advancement of soft robotics, the research community started exploring the use of continuum and soft robots in unstructured environments.
+In particular, when interacting with irregularly shaped objects, shear deformation cannot be neglected.
+Additionally, advancements in actuation technologies necessitate the consideration of stretch and torsion modes.
+Within this context, the CRT emerged as a fundamental tool for modeling the deformation of continuum and soft robots.
+A recent contribution by Tuummers2023Cosserat further consolidates CRT as a standard framework, by rigorously deriving both Newtonian and Lagrangian formulations for multi-segment tendon-driven continuum robots and demonstrating their equivalence, with practical guidelines for numerical implementation.
+As discussed in Sec.
+4, the CRT describes the motion of an elastic rod with an infinite number of DoFs.
+Discretization techniques play a key role in numerically solving the governing equations of Cosserat rods.
+Several methods were developed to enhance numerical robustness and computational efficiency.
+For example, renda2018discrete introduced the piecewise constant strain (PCS) method, which discretizes a continuous Cosserat rod into segments with constant strain.
+This method extends the PCC approach (webster2010design) by incorporating shear and torsion.
+In parallel, till2019real proposed a real-time discretization technique that transforms the PDE of a Cosserat rod into a boundary value problem (BVP) by discretizing time derivatives, significantly improving computational performance.
+Incorporating external forces such as buoyancy, drag, and added mass is essential for accurately modeling the interactions between soft robots and their surroundings.
+In this context, renda20123d; renda2014dynamic analyzed the static and dynamic behavior of a silicone arm in a dense medium, capturing its coupled tendon conditions and external loads.
+The model was further developed in a unified Cosserat rod framework applied to underwater robots (renda2018unified).
+By accounting for nonlinearities and hydrodynamic effects, this approach enabled the evaluation of soft robot designs under submerged conditions.
+For pneumatic actuators, renda2017screw extended the PCS method by integrating Screw Theory, allowing flexible routing of fluidic and tendon actuators along arbitrary paths.
+In PCM, armanini2021discrete adapted the PCS formulation to closed-chain soft robotic systems, facilitating the modeling and design optimization of structures like the Fin-Ray finger.
+Material nonlinearities, such as hysteresis, can significantly affect the performance of soft robots.
+mishra2022fractional tackled the challenge of large deformations in pneumatic soft robots by proposing a fractional-order Bouc-Wen model within a Cosserat rod framework.
+By incorporating nonlinear material effects, the model demonstrated superior accuracy in capturing the behavior of soft actuators compared to traditional approaches.
+Extensive validation of CRT models is essential for ensuring their predictive accuracy.
+Cable-driven soft robots were widely studied and experimentally validated.
+For example, renda20123d validated a static Cosserat model of a tendon-driven soft arm using FEM simulations and experimental comparisons.
+The dynamic extension of this model (renda2014dynamic) was experimentally verified by replicating characteristic octopus movements.
+Later, renda2018unified demonstrated the effectiveness of their unified dynamic framework using an underwater robotic system.
+Further validation of discretization-based approaches was performed in renda2018discrete and till2019real, confirming the numerical robustness of their methods.
+Conversely, butler2019continuum investigated deflection and stiffness in robots subject to constant axial tendon displacement and environmental loads, comparing parallel and converging actuator paths.
+They validated the Cosserat rod model on a CSM subject to tip loads in the kg range, providing actuations to reach bending angles up to .
+The simulation results aligned well with the experimental data.
+However, as the authors noted, limitations emerged for larger curvatures due to the increased role of friction, which was not included in the model.
+For pneumatic soft robots, trivedi2007geometrically; trivedi2008geometrically developed geometrically exact CRT models and validated them on fiber-reinforced pneumatic manipulators, achieving significantly higher accuracy than constant curvature models.
+The screw-theoretic PCS formulation (renda2017screw) was validated on the STIFF-FLOP arm, while Roshanfar2021towards tested their pneumatic Cosserat rod model experimentally, reporting an tip position error.
+Also, FEM was extensively used to compare and validate CRT models.
+Indeed, grazioso2018analytic derived analytical solutions for quasi-static Cosserat rods and verified them against FEM simulations under external loads.
+Similarly, berthold2021preliminary conducted a comparative study between FEM and CRT models for pneumatic actuators, proposing a parameter identification process.
+Further refinement was made by wiese2022towards, who obtained a static Cosserat model from detailed FEM simulations, validating it experimentally across various loading conditions.
+For parallel continuum robots, till2017elastic investigated elastic stability using Cosserat rod models and experimentally validated their stability conditions.
+In parallel, black2017parallel introduced a kinetostatic CRT framework for PCM, achieving high accuracy in force sensing and manipulability analysis.
+Then, lilge2022kinetostatic extended this approach, developing a comprehensive CRT-based framework for tendon-driven PCM, experimentally achieving a median pose accuracy of .
+Cosserat rod models found widespread applications in bioinspired robotics and biomechanical studies.
+boyer2017locomotion proposed a unified CRT-based framework for modeling the locomotion of bioinspired robots with soft appendages.
+They analyzed flapping flight and passive swimming, comparing two different modeling approaches based on Newton-Euler dynamics.
+For musculoskeletal modeling, zhang2019modeling implemented Cosserat rod simulations in PyElastica (naughton2021elastica), assembling heterogeneous structures to replicate musculoskeletal architectures (e.g., snakes, bird wings).
+Their work demonstrated the effectiveness of CRT in bioengineering design, simulating complex musculoskeletal interactions in biological and robotic applications.
+boyer2022statics explored the intersection of CRT and optimal control theory, revealing fundamental insights into the singularity of dynamic continuum robot simulations.
+By reformulating the problem as a minimization task, their approach provided a novel perspective on the statics and dynamics of continuum robots, validated through simulations of bioinspired continuum swimmers.
+Finally, to find optimal bases for the GVS approach, alkayas2024soft introduced a data-driven reduction method based on Proper Orthogonal Decomposition (POD).
+Their approach begins by creating a high-fidelity GVS digital twin from experimental data, which is then simulated to generate a comprehensive strain dataset.
+By applying Singular Value Decomposition (SVD) to this data and truncating the least significant singular values, the model’s dimensionality is reduced while preserving its accuracy.
+This process yields an optimal linear basis of the form .
+Building on this work, a subsequent study extended the search to the nonlinear case (i.e., ), employing autoencoders to identify a more expressive and compact representation (alkayas2025structure).
 6.3 Bend & Twist
-Like bending, various robot types exhibit the combination of bending and twisting induced by diverse actuation types (Fig. 5). Table 3 summarizes the Bend & Twist deformation class.
-4.3.1 Bend & Twist by Cables
-Bending and twisting are usually induced by cables and tendons through specific routing paths (Fig.
-5).
-Rucker and Webster III (2011) present Cosserat rod models for the forward kinematics, statics, and dynamics of a tendon-driven continuum manipulator.
-The curved tendon paths deform the body in various shapes.
-Starting from the classical CRT, they derive the expression of distributed loads applied by tendons on the backbone .
-The model validation considers straight and non-straight tendon routings, point loads, and distributed external loads .
-The simulated matched experimental tip positions with a 1.7%L average error.
-Another experimental work presented a continuum arm with a Nitinol backbone and three tendons turning around the centerline with a constant radial distance (Dehghani and Moosavian, 2013), which extended a previous setup lacking an actuation system (Dehghani and Moosavian, 2011).
-They modified differential terms of an existing CRT (Rucker and Webster III, 2011), retaining the original accuracy with reduced numerical effort.
-The model was extended to real-time control using a stable moment-based algorithm to solve the BVP considering the physical characteristics of the rod (Dehghani and Moosavian, 2014a).
-They also propose a method to reduce the cumulative numerical integration errors that could make the rotation matrices non-orthogonal.
-Finally, they use the model to develop a Jacobian-based quasi-static tip position controller for a multi-segment continuum arm.
-Cosserat rods can simulate the mechanics of several robotic systems, including robotic tails and octopus-inspired arms.
-The work in (Rone and Ben-Tzvi, 2014) used a Cosserat rod to study a cable-driven multi-section continuum robotic tail envisioned for onboard mobile robots.
-Prescribed tail configurations or trajectories are analyzed, while the governing equations compute the loading at the tail base transmitted to the mobile robot.
-This study investigates the impact of trajectory and design factors on the resulting loading profiles, considering factors like mode shape, speed, bending magnitude, and bending plane angle.
-Key findings include that a non-straight tail reference configuration increases the moment magnitude, short proximal segments provide a larger tip motion for a given bending angle, and higher mass distribution towards the final nodes enhances bending moments and control fidelity.
-Xu et al.
-(2017) propose a Cosserat rod model to describe the bending and torsion behavior of a bionic tendon-driven octopus tentacle.
-Given the location of the four cables, the model maps cable tension to expected bending.
-Exploiting the derived mapping, they propose an open-loop controller for the robot.
-Similarly, Niu et al.
-(2019) derive closed-form equations of a soft robot modeled through the CRT.
-The equations are validated with a cable-driven soft arm inspired by an octopus tentacle, reaching accurate predictions of the end-effector positions.
-Chikhaoui et al.
-(2019) investigated the kinematic modeling of a magnetically extensible tendon actuated continuum robot (METAbot) (Nguyen and Burgner-Kahrs, 2015).
-The comparison between a PCC beam mechanics approach and a Cosserat rod shows that the latter has higher position accuracy, although computationally heavier.
-Recently, Janabi-Sharifi et al.
-(2021) presented a tutorial on the dynamic Cosserat rod modeling of tendon-driven continuum robots.
-The tutorial explains the core assemblies of tendon-driven continuum robots.
-Then, it provides a unified approach for the dynamic analysis of tendon-driven continuum robots that bend and twist, including multi-section, concentric tubes, and parallel cooperative robots.
-4.3.2 Bend & Twist in Soft Pneumatic Actuators
-Actuation-induced torsion is not common in fluidic actuators.
-However, a Cosserat rod model for a CSM with two asymmetric pneumatic actuators that bend and rotate was introduced (Uppalapati et al., 2018).
-They optimize the rigidity parameters and the curvatures for quasi-static shapes.
-The model reproduced the experimental static workspace with a tip load attached.
-The robot’s dexterity was enhanced integrating an additional soft rotating actuator (Uppalapati and krishnan, 2021).
-The CRT predicts the deformed robot configuration subjected to external loads, and they experimentally determine elasticity and actuation parameters.
-Another study introduced a multi-segment Euler-Bernoulli rod model to predict the bending deformation of micro-tube pneumatic actuators (Ji et al., 2019).
-They first employ the line-segment model and then extend it to a multi-segment version of Euler-Bernoulli’s rod model for a rapid design, analysis, and optimization of the bending deformation of such actuators.
-Finally, an analytical model of a generalized PneuNet soft actuator combines the minimum potential energy method and an Euler-Bernoulli rod (Gu et al., 2021).
-The study compared the rod model to FEM simulations and experimental results, achieving accurate deformed shape predictions, given the geometry and the applied pressures.
-4.3.3 Bend & Twist in Tubular Robots
-Biomedical applications could leverage CTR and eccentric tube robots that bend and twist.
-A geometrically exact KRT solved the statics problem for CTR, describing the shape of a collection of concentric pre-curved tubes under general wrench distributions (Rucker et al., 2010a).
-The model is validated experimentally for an active cannula with three concentric, telescopic, pre-curved Nitinol tubes in several configurations with varying tip loads.
-The results establish the basis for model-based position and force control of active cannulas and their use as force sensors under external loads.
-Also, they tackle the integration of unmodelled effects such as friction, elongation, and transverse shear (Rucker et al., 2010b).
-Subsequently, Till et al.
-(2020) present the first dynamic model for CTR.
-The model describes the coupled inertial dynamics of pre-curved concentric tubes considering the dynamic effects of linear and rotational tube inertia, material damping, Coulomb and viscous friction, and the inertia of a rigid body at the robot tip.
-The experimental validation measures the dynamic behavior during tissue grasping and snapping.
-The structure of the ETR extends the CTR design.
-A kinematic model of a two-channel ETR with two continuum arms delivered through the channels is presented (Wang et al., 2019).
-Pre-curved superelastic tubes lining the arm’s working channels produce sheath deflection.
-Cosseat rods model the pre-curved tubes and the central backbone, achieving promising tip pose predictions for pure rotation actuation.
-Similarly, a continuum ETR sheath for MIS such as neuroendoscopy, requiring delivery of multiple robotic arms, was presented (Wang et al., 2021a).
-Super-elastic tubes lining the arm channels actuate the sheath, while distributed disks are attached to the backbone.
-The tube kinematics and the robot backbone are modeled using the CRT.
-Three experiments with a two-arm sheath validate the tip position accuracy.
-Another work proposes a kinematic Cosserat rod model of a snake-like continuum robot for MIS (Wang et al., 2020).
-The robot geometry comprises three identical super-elastic Nitinol tubes held eccentrically by a surrounding sheath.
-Assuming sheat inextensibility, they show that different actuations virtually yield the same robot shape and tip pose.
-Simulations and experiments validate the model for a single-stage ETR.
-Safety is crucial in surgical procedures.
-A method to compute compliance of continuum robots, modeled as quasi-static Cosserat rods under diverse loads, was devised (Smoljkic et al., 2014).
-The robot compliance is described by ODEs and integrated with the forward kinematics.
-The experiments involve a non-actuated flexible rod composed of a super-elastic Nitinol tube surrounding a braided Pebax tube.
-The rod mounted on a force sensor with position sensors embedded in the tip is subject to external tip forces.
-A least square fitting of the measured tip trajectory versus the direct kinematics identifies material parameters and tube configuration.
-Finally, they compared the velocities obtained through position differentiation against those obtained multiplying the differentiated forces and the computed compliance matrices.
-4.3.4 Bend & Twist with Smart Actuators
-The KRT modeled TCP emphasizing critical factors like coil load and twist and incorporating them into the actuation equations for TCP muscles (Wu and Zheng, 2020).
-The model captures the quantitative relationship between tensile actuation and fabrication load, serving as a predictor for tensile stroke.
-Also, it was validated through TCP muscles made of silver-coated Nylon fiber under different fabrication loads, measuring recovery stress and maximum stroke during heating.
-Thermo-mechanical tension tests conducted at various room temperatures identified the elastic modulus of the fibers.
-To achieve variable stiffness in continuum robots, Wang et al.
-(2022b) propose to use winding-styled ropes inspired by snakes’ behavior.
-Controlling the temperature of an SMA spring increased their stiffness up to .
-The experimental validation of a kinetic-static Cosserat rod model obtained an error below .
-4.3.5 Bend & Twist in Parallel Manipulators
-4.3.6 Bend & Twist with Buckling Instability
-Buckling instability is an important factor to consider in soft robots.
-With applications in soft robotics, Sipos and Várkonyi (2020) explores the interplay between spatial buckling and arbitrary intrinsic curvature on a simulated clamped Cosserat rod with a vertical tip load.
-The aim is to tune the control parameters to find the highest load for a given free-end position or the maximum reaching of the cantilever in a direction for a given load.
-The numerical investigation of the spatial instability of in-plane configurations of soft elastic cantilevers subject to an endpoint load demonstrates the importance of out-of-plane instability.
-Moreover, the capability of controlled intrinsic curvatures to stabilize the rod in configurations of high reaching is shown.
-Significant improvements occur for flattish cross-sections where the rigidities associated with out-of-plane bending and torsion are larger than the in-plane stiffness.
-At the same time, arms with high and narrow cross-sections are slightly more performant.
-Buckling instability can be triggered.
-Forghani et al.
-(2021) analyze the inverse dynamics and control of a bacteria-inspired uniflagellar robot in a fluid medium at a low Reynolds number.
-The robot, mimicking the locomotion of flagellated bacteria, features a flexible helical filament (flagellum) connected to a spherical head.
-A Kirchhoff elastic rod describes the flagellum, which rotates about the head at a controlled angular velocity and generates propulsive forces for the forward motion.
-The results demonstrate that by triggering the buckling instability of the flagellum, the robot follows a prescribed path in space.
-The control scheme only involves the flagellum angular velocity, and a deep neural network captures the dynamics predicting the robot’s trajectory.
-However, higher maneuverability tasks like drug delivery require multiple flagella, advanced controllers, and control inputs.
-4.3.7 Bend & Twist in Insect-like Wings
-The applicability of rod theories goes beyond manipulators.
-A study presented a hybrid dynamic Cosserat rod model of a 3D aerial insect-like robot with two soft wings (Belkhiri et al., 2012).
-In particular, Cosserat rods described the leading edge of the wings, which bend and twist based on continuous Newton-Euler dynamics.
-The wings, connected to the insect thorax through revolute joints, are treated as rigid bodies.
-The hybrid model combines discrete and continuous models considering the curvature deformations and external aerodynamic forces.
-The model replicates the kinematics and internal deformations of the wings and the rigid net insect motion produced by a flapping flight.
-Subclass | References | Rod Model | Contribution |
-Cable | Rucker and Webster III (2011) | CRT | General dynamic model of a continuum manipulator with curvilinear actuators. |
-Pneumatic | Uppalapati and krishnan (2021) | CRT | Design and modeling of a CSM with parallel asymmetric fiber-reinforced elastomers. |
-Tubular | Till et al. (2020) | CRT | Dynamic model of CTR validated for tissue grasping and snapping. |
-Wang et al. (2021a) | CRT | Modeling of a superelastic ETR sheath for biomedical procedures. | |
-Smart | Wu and Zheng (2020) | KRT | TCP muscles incorporating coil load and twist into actuation model. |
-Buckling | Sipos and Várkonyi (2020) | CRT | Study relationship between buckling and on simulated cantilever rods. |
+Bending and twisting are the fundamental deformation modes of CSM, enabling them to achieve complex 3D shapes and dexterous movements.
+Modeling these modes is particularly important in applications where interaction with the environment is not limited to planar motions but requires full spatial maneuverability.
+In many cases, incorporating twisting alongside bending provides the best trade-off between accuracy and computational efficiency.
+This allows continuum robots to follow arbitrary 3D curves in space, which is essential for tasks such as wrapping around an object, executing dexterous maneuvers, or navigating constrained environments.
+Biomedical applications especially benefit from bending-twisting modes, as in concentric tube robots for minimally invasive surgery (MIS) (rucker2010model; rucker2010geometrically; till2020dynamic) and eccentric tube robots for multi-arm sheaths (wang2019steering; wang2020eccentric; wang2021eccentric).
+Safety-critical domains also utilize bending-twisting compliance analysis to predict deformations under external loads (smoljkic2014compliance).
+Twisting deformations are usually induced by particular cable or tendon routing.
+Tendon-driven continuum robots employ non-straight tendon paths that wrap around the backbone to generate both bending and torsional modes.
+Early work by rucker2011statics established Cosserat rod formulations where curved tendon paths produce distributed loads along the backbone.
+Similarly, dehghani2013modeling developed a Nitinol-backbone continuum arm with three tendons spiraling around the centerline, extending their previous untwisted setup (dehghani2011mod_cont).
+They further refined this approach for real-time control using a stable moment-based algorithm (dehghani2014compact).
+Twisting was also demonstrated in robotic tails (rone2014continuum), octopus-inspired manipulators (xu2017modelling; niu2019closed), tendon-actuated extensible robots (chikhaoui2019comparison), and multi-section tendon robots (janabi2021cosserat).
+In fluidic actuation, torsional effects are less common but can be induced through asymmetric chambers or fiber reinforcement.
+For example, uppalapati2018parameter modeled a manipulator with two asymmetric pneumatic actuators capable of bending and rotating, which were later extended with an additional rotating actuator (uppalapati2021design).
+Other approaches include models based on Euler-Bernoulli for micro-tube pneumatics (ji2019rapid) and generalized PneuNet structures, combining bending and twisting predictions validated against FEM (gu2021analytical).
+Cosserat rods were extensively applied to tendon-driven systems, while Euler-Bernoulli or energy-based methods are often preferred for pneumatic networks due to their simpler kinematics.
+Examples include asymmetric pneumatic actuators capable of torsion (uppalapati2018parameter; uppalapati2021design), micro-tube pneumatics modeled with multi-segment Euler–Bernoulli rods (ji2019rapid), and PneuNet actuators modeled analytically with energy minimization (gu2021analytical).
+These models capture how chamber geometry, reinforcement, and pressurization induce coupled bending and twisting modes.
+However, the majority of bending-twisting continuum robots are tendon-driven, concentric-tube, or eccentric-tube systems.
+These architectures require a geometrically exact description of distributed loads and torsional effects, which is naturally expressed through the CRT.
+Concerning CTR, renda2021sliding adapted the piecewise variable-strain (mathew2025reduced) to simulate multi-section CTR by including the tubes’ sliding motion.
+Rotation motions of the tubes are included as generalized coordinates rather than boundary kinematic conditions.
+Notably, this procedure led to a minimal set of closed-form algebraic equations solvable for both the shape variables, the actuation forces, and the torques.
+The resulting equations also facilitate control, design optimization, and stability assessment.
+Unlike beam simplifications, Cosserat rods account for coupling between bending and torsion and handle spatially varying loads from tendon paths, external wrenches, and contact.
+This explains why most bending-twisting works in the literature adopt the Cosserat rod formalism (rucker2011statics; dehghani2013modeling; dehghani2014compact; rone2014continuum; xu2017modelling; niu2019closed; chikhaoui2019comparison; janabi2021cosserat; wang2019steering; wang2020eccentric; wang2021eccentric).
+Buckling instability is another important nonlinear phenomenon to consider.
+When modeling thermally activated twisted and coiled polymer (TCP) muscles, wu2020modeling employed KRT, since the small-strain assumptions and direct link to coil load-twist relations simplify the actuation model.
+Similarly, Kirchhoff rods were used in micrometer-scale parallel continuum robots (mauze2021micrometer) and bacteria-inspired flagellated swimmers exploiting buckling instabilities for propulsion (forghani2021control).
+Conversely, Cosserat rods were applied to study global buckling in soft robotic arms (sipos2020longest), where large deflections and intrinsic curvature control are essential.
+Thus, Kirchhoff rods often appear in actuation-centric or microscale studies where analytical tractability is crucial, while Cosserat rods dominate in large-deformation, shape-control analyses.
+Finally, workspace analysis requires accurate models of bending and twisting under actuation and loading.
+Tendon-driven Cosserat models enable forward kinematics and Jacobian-based control for workspace prediction (dehghani2014compact).
+For pneumatic robots, static workspace matching was validated experimentally with asymmetric actuators modeled by Cosserat rods (uppalapati2018parameter; uppalapati2021design).
+PneuNet and micro-tube actuator models (ji2019rapid; gu2021analytical) allow for workspace design and optimization by linking actuator geometry and pressure inputs to achievable curves.
+In biomedical contexts, concentric and eccentric tube models predict reachable poses in confined environments (rucker2010model; wang2021eccentric).
+| Deformation Class | Subclass | Reference | Contribution | Rod Model |
+| Bend | Cable | jones2009three_dim | Static model for real-time shape estimation of CSM. | CRT |
+| dehghani2014dynamics | Dynamic model of a continuum finger for planar grasping. | CRT | ||
+| gao2016mechanical | Model of notched continuum manipulator for MIS. | CRT | ||
+| Pneumatic | de2017development | Derivation of constitutive law of a planar bending actuator. | EBRT | |
+| sadati2019reduced | ROM and rod models with absolute and relative states validated on STIFF-FLOP. | EBRT | ||
+| Smart | daily2017dynamical | Dynamic model inspired by caterpillar motion to analyse ondulation mechanics. | EBRT | |
+| Pawlowski2018modeling | Modeling the coupling between CSM body and thermal actuation. | CRT | ||
+| Parallel | li2018shape | Shape modeling of a PCM made of soft panels. | EBRT | |
+| Tubular | pattanshetti2019kinematic | Estimation of stiffness variation on a monolithic tube for MIS. | CRT | |
+| All Deformations | Cable | rucker2011statics | Derived exact models for the forward kinematics, statics, and dynamics of CSM with general tendon routing. | CRT |
+| renda2014dynamic | Studied the dynamic interaction of an octopus-inspired CSM with a dense medium. | |||
+| butler2019continuum | Predicted deflection and stiffness in CSM subject to constant actuation and external loads. | |||
+| Pneumatic | trivedi2008geometrically | Model validation of CSM with fiber-reinforced actuators varying base orientation. | ||
+| Parallel | till2017elastic | Studied and experimentally validated the elastic stability of PCM. | ||
+| armanini2021discrete | Modeled closed-chain soft PCM with validation on a soft-rigid finger. | |||
+| lilge2022kinetostatic | Studied reachability, singularities, manipulability, and compliance of PCM. | |||
+| Bioinspired | zhang2019modeling | Simulates musculoskeletal systems assembling heterogeneous active/passive rods. | ||
+| boyer2022statics | Studied the link between CRT and optimal control simulating locomotors and swimmers. | |||
+| Bend & Twist | Cable | janabi2021cosserat | Tutorial on the dynamic modeling of tendon-driven continuum robots that bend and twist. | CRT |
+| Pneumatic | uppalapati2021design | Design and modeling of a CSM with parallel asymmetric fiber-reinforced elastomers. | CRT | |
+| Tubular | till2020dynamic | Dynamic model of CTR validated for tissue grasping and snapping. | CRT | |
+| wang2021eccentric | Modeling of a superelastic ETR sheath for biomedical procedures. | CRT | ||
+| renda2021sliding | Adapted the piecewise variable strain to CTR including the tubes’ sliding motion. | CRT | ||
+| Smart | wu2020modeling | TCP muscles incorporating coil load and twist into actuation model. | KRT | |
+| Buckling | sipos2020longest | Study relationship between buckling and on simulated cantilever rods. | CRT | |
+| Bend & Stretch | Cable | xiao2023kinematics | Addressed stiffness regulation for a soft arm with a sliding backbone, achieving +57.7% stiffness. | CRT |
+| chen2022model | Augments the PCC with a rod model to quickly solve the CSM deformation under . | EBRT | ||
+| Pneumatic | sadati2017control | Utilizes the Ritz-Galerkin method to reduce the continuous state space of rod models. | CRT | |
+| gilbert2019validation | Experimentally validated of a CSM model on bending motions generated by square pulse actuations. | CRT | ||
+| Garbulinski2022bending | Characterized bending, natural frequency, and damping of a cantilevered extensile artificial muscle. | EBRT | ||
+| alessi2023ablation | Ablation study of model addressing manufacturing uncertainties in a 3D-printed CSM. | CRT | ||
+| dou2023design | Design and modeling of a hybrid CSM incorporating an antagonistic compliant mechanism. | EBRT | ||
+| Smart | goldberg2019planar | Simulated locomotion of a caterpillar-inspired soft robot actuated by SMA. | CRT | |
+| Rod-driven | wu2022design | Design, statics modeling, and numerical workspace analysis of an extensible rod-driven PCM. CRT | ||
+| Bend & Twist & Stretch | Cable | renda2012general | Simulated the coupled tendon drive for multi-section CSM highlighting the role of twist. | CRT |
+| wang2016three | Compared the Kelvin viscoelastic material model vs a purely elastic model on a real siliconic arm. | |||
+| adagolodjo2021coupling | Combined FEM to model the robot body with CRT to model actuation. | |||
+| Roshanfar2023Hyperelastic | Investigated pneumatic stiffness regulation on a cable-driven CSM using the hyperelastic material model. | |||
+| Tubular | ryu2018application | Reconstructed the shape of a continuum endoscope given external forces exerted on the distal end . | CRT | |
+| Parallel | ghafoori2020modeling | Modeled a six-link PCM starting from an experimental validation of a single elastic rod. | CRT | |
+| Bend & Stretch & Shear | Tubular | chen2021variable | Experimentally validates a model for multi-backbone continuum robot. | CRT |
+| Pneumatic | bartholdt2021parameter | Models a soft pneumatic actuator combining rods and ANN | ||
+| Bend & Twist & Shear | — | grazioso2019geometrically | Proposed a novel FEM-inspired discretization technique and the SimSoft simulator. | CRT |
+| Tubular | bentley2023safer | Proposed a force model and associated cost metric for safer and clinically relevant for motion planning. | ||
+| Stretch & Shear | Bioinspired | hemingway2021continuous | Studied the peristaltic locomotion of the earthworm Lumbricus terrestris to model soft robots. | CRT |
+| Stretch | Bioinspired | Hyousse2019distributed | Studied the peristaltic locomotion of a caterpillar-inspired soft robot. | CRT |
 6.4 Bend & Stretch
-The simultaneous bending and stretching emerge in many cable-driven and pneumatic CSM (Fig. 5). Table 4 reports illustrative examples of the Bend & Stretch class.
-4.4.1 Bend & Stretch by Cables
-Deflection and stiffness estimation are crucial for continuum robots actuated by cables.
-For instance, robot deflection and stiffness were predicted considering environmental loads with constant axial tendon displacements (Oliver-Butler et al., 2019).
-Preliminary experiments with an Euler–Bernoulli rod evaluate how tendon stretch and routing path affect the robot stiffness at any point.
-The analytical results enable stiffness analysis of candidate robot designs with light simulations and suggest the ability to increase robot stiffness by using converging tendon paths.
-Finally, they extend the study to a non-linear Cosserat rod model with large deflections, handling prescribed tendon displacements, tendon stretch, pretension, and slack.
-Experimental validation on a prototype shows good agreement.
-Also, a novel stiffness modulation approach using a Cosserat rod model was proposed for a tendon-driven soft CTR (Xiao et al., 2022).
-The experimental validation shows an accurate shape estimation.
-They also demonstrate that the robot stiffness changes with a Nitinol backbone.
-In particular, the tip compliance can be halved, with potential applications in manipulation tasks.
-Concerning workspace estimation, the CRT is also applied to investigate equilibrium positions reachable by the end-effector of a planar tendon-driven CSM with a discrete actuation space (Walid et al., 2022).
-Experiments on a physical prototype efficiently estimated the robot workspace.
-Another example predicts the soft arm shape combining the EBRT, which considers the influence of external forces in bending, with the PCC model, which efficiently derives a relationship between cable length and arm curvature (Chen et al., 2022).
-Experiments on a cable-driven soft arm subject to external forces, or , yield a maximum error in the range .
-4.4.2 Bend & Stretch in Soft Pneumatic Actuators
-A study compares five models, including the Cosserat rod, a dynamic lumped system, and approaches with constant or variable curvature (Sadati et al., 2017b).
-The comparison made in simulation and on the STIFF-FLOP soft arm (Cianchetti et al., 2013) demonstrated that the lumped mass model better predicts the dynamics, while constant curvature methods are less accurate but are suitable for design optimization.
-Additionally, they modeled CSM actuated by braided pneumatic actuators (Sadati et al., 2017a).
-This specific design causes cross-sectional deformation, which was captured analytically through geometry deformation and CRT.
-This approach outperformed constant curvature approaches by and variable curvature by .
-A subsequent work presents a semi-analytical model for the dynamics of multi-segment CSM (Sadati et al., 2017c).
-The Cosserat rod static and Lagrangian dynamic models are derived and combined with the Ritz-Galerking method to reduce the infinite state space of CSM to a small finite number of states.
-Moreover, a static Kirchhoff rod model was proposed for a hybrid CSM characterized by pneumatic actuators with embedded elastic rods, computing its deformations using the minimum energy principle (Sun et al., 2019).
-In addition, the classical CRT modeled a variable-length pneumatic CSM (Gilbert and Godage, 2019).
-They optimize the cross-sectional area of the actuators and a parameter relating stiffness and damping in the constitutive law.
-The model is validated experimentally on dynamic bending deformations obtained by pressurizing the actuators with alternating square pulses.
-A study characterized the bending behavior of a cantilevered extensile fluidic artificial muscle (Garbulinski and Wereley, 2022).
-An Euler-Bernoulli rod determined static properties like the bending stiffness, which increases more than four times in the pressure interval psi.
-Experimental data are applied to assess the dynamic properties, such as natural frequency and damping.
-Also, it was developed an analytical model of a PneuNet actuator based on the Euler-Bernoulli finite strain rod (Sachin et al., 2022).
-They describe the actuator behavior in free motion and for tip contacts and accurately predict the actuator deformation and the force characteristics.
-Another work showcases a parallel extensible soft robot with three pneumatic actuators, each with a silicone body, inner air chamber, winding wiring, and sealing head (Wang et al., 2022c).
-The dynamic model of a single actuator uses the CRT, and an interactive constraint topological structure solves the constraint force and moment of the parallel robot.
-A linear mapping between the pressure and the extension force is established for each actuator.
-Validation measures the static step response of a single actuator and conducts dynamic experiments with the complete soft robot.
-The model accurately captures the actuator dynamics and the parallel robot constraints.
-Rod models met traditional robotics models where the EBRT augmented the kinematic model of a pneumatic soft bending finger treated as a planar serial rigid robot parameterized with the Denavit-Hartenberg convention (Flores-Martínez et al., 2022).
-The experimental validation compares the actuator trajectory versus the kinematic model and the flexion-extension motion trajectory of average human fingers.
-Manufacturing uncertainty could be significant in fluidic soft actuators.
-For instance, Eugster et al.
-(2022) described the kinematics of a soft arm using a CRT with a nonlinear pressure-dependent constitutive law, using the principle of virtual work and modeled the actuator with strain energy functions.
+Simultaneous bending and stretching behaviors are frequent in CSM actuated by cables or pneumatics (Figure 4).
+Notably, this category excludes twist and shear effects for several reasons: the actuators are linear and apply forces along the centerline, not generating torsional or shear stresses; interactions with the environment are either neglected or primarily induce axial and bending loads.
+As a result, the deformation behavior can be described through bending and axial stretching.
+We summarize several Bend & Stretch works in Table 2, whereas herein we focus the discussion on manufacturing errors, stiffness regulation, and fiber-reinforced actuators.
+Manufacturing variability and material degradation are critical concerns in soft robotic systems, as they can significantly affect actuator behavior, model accuracy, and control performance.
+Variations in geometry, material stiffness, or the positioning of actuators during fabrication can cause substantial differences between the modeled behavior and the robot’s actual motion.
+While several experimentally validated rod models were developed for pneumatic soft arms (sadati2017geometry; sadati2017mechanics; sadati2017control; gilbert2019validation; Wang2022DynamicsMod; Lamping2023ANovel) and fingers (sachin2022analytical; flores2022soft), relatively few have focused specifically on the effects of manufacturing irregularities.
+Below, we discuss two works using Cosserat rod models to address manufacturing uncertainties in pneumatic soft robots.
+eugster2022soft described the kinematics of a soft arm using a nonlinear pressure-dependent constitutive law, the principle of virtual work, and modeled the actuator with strain energy functions.
 Manufacturing imperfections are considered by scaling the relations of extensional and bending stiffness and by a constant shift of the actuator position .
 Assuming the pressure dilates the chambers, the model also considers cross-section deformations with a pressure-dependent diameter .
-The model is validated experimentally for static stretching and bending motions in vertical and horizontal mountings.
-They compare models with constant or pressure-dependent chamber radius, with or without chamber repositioning on the cross-section, and with a linear or Ogden material law.
-The strain energy for an Ogden material (Odgen et al., 2013) is
-Here, is a material parameter, and is the principal axial stretch.
-Similarly, Alessi et al.
-(2023a) presented a dynamic Cosserat rod model for a 3D-printed pneumatic CSM.
-The model considers the pressure transient with a parameterized smooth step.
-Also, it captures the manufacturing variability of the actuators using strain gains that tune the pressure-strain relation for each actuator
+The model is validated experimentally for static stretching and bending motions in vertical and horizontal mountings, comparing constant or pressure-dependent chamber radius, with or without chamber repositioning, and with a linear or Ogden material law.
+Similarly, alessi2023ablation presented a dynamic model for a 3D-printed CSM, where the manufacturing variability caused the actuators to have different stretching capability when subject to equal actuation.
+They captured this effect by tuning the pressure-strain relation for each actuator introducing strain gains that tune the pressure-strain relation for each actuator
 where is the interior actuator area.
-An extensive experimental validation includes quasi-static and dynamic motion patterns at various actuation frequencies.
-Finally, an ablation study evaluates the contribution of different model components, uncovering that neglecting gravity and manufacturing uncertainties could degrade performance up to , depending on the motion.
-Concerning fiber-reinforced actuators, Berthold et al.
-(2022) introduces a Cosserat rod model to deepen the understanding of the pressure-induced deformation behavior of a fiber-reinforced soft pneumatic actuator with three DoFs.
-By converting the inner chamber pressure into an equivalent force, they consider the compression effects of lateral pneumatic chamber walls.
-The validation compared the results to FEM simulations, obtaining a maximum mean tip position error of .
-Differently, Hanza and Ghafarirad (2023) introduces an inhomogeneous Cosserat rod model for a CSM with pneumatic fiber-reinforced actuators, using a nonlinear strain-force relationship to account for the effect of radial pressure.
+An ablation study conducted on dynamic motions investigated the contribution of different model components, uncovering that neglecting manufacturing uncertainties can degrade performance up to .
+In practice, manufacturing irregularities can affect every soft robotic system but remain unexplored outside the Bend & Stretch class.
+We attribute this to the fact that, for linear actuators, it is easier to quantify the expected motion and thus identify model discrepancies.
+Stiffness modeling and regulation are central to advancing the adaptability and precision of soft robotic systems.
+However, in many designs, limitations such as single-mode actuation, the absence of antagonistic mechanisms, or the lack of a structural backbone can restrict the ability to modulate stiffness effectively.
+To address these challenges, Sun2018hybrid presented a novel design of a hybrid continuum robot whose actuators combine pneumatic muscles with embedded elastic rods.
+The robot changes its stiffness through a locking mechanism, switching between large-scale movement enabled by pneumatics and fine positioning enabled by push-pull of the rods.
+Stiffness tests revealed that the robot increases its structural stiffness by 65.1% during fine positioning in the locked mode and reduces the repetitive positioning error by 62.8%.
+To model this behavior, they improved a static KRT, calculating the elastic deformation through the minimal total potential energy principle in an optimal control framework.
+Similarly, xiao2023kinematics addressed stiffness regulation for a tendon-driven concentric tube soft robot with a Nitinol tube backbone, which can slide inside the soft robot body for pose or stiffness regulation.
+A Cosserat rod model was validated in different scenarios by varying the joint-space tendon input and task-space external contact force, achieving an average tip position error below 1% of the total length.
+Simulation studies demonstrate that the Nitinol backbone insertion can enhance the kinematic workspace and increase stiffness by 57.7%.
+Two manipulation case studies demonstrated the potential application of hybrid actuation for stiffness regulation.
+Another compliant rod-driven mechanism enabling stiffness regulation was introduced in a soft arm with pneumatic actuation (dou2023design).
+While an Euler-Bernoulli model was adapted to demonstrate the benefit of the mechanism on experiments of way-point tracking, the investigation of the stiffening behavior remains a future work.
+Fiber-reinforced soft pneumatic actuators present significant modeling challenges due to nonlinear deformation under pressure.
+Several works leveraged variations of rod theories to more accurately capture their complex bending, elongation, and cross-sectional behaviors.
+sadati2017geometry; sadati2017mechanics; sadati2017control performed comprehensive investigations of several modeling approaches—all experimentally validated on the well-known STIFF-FLOP arm—addressing the challenge of cross-section deformation, captured analytically through geometry deformation and CRT.
+This approach outperformed constant curvature approaches by and variable curvature by .
+More recently, Berthold2022investigation introduced a Cosserat rod model to deepen the understanding of the pressure-induced deformation behavior of a fiber-reinforced soft pneumatic actuator with three DoFs.
+Specifically, they considered the compression effects of lateral pneumatic chamber walls by converting the inner chamber pressure into an equivalent force.
+Their model, validated against FEM simulations, obtained a mean tip position error up to .
+Conversely, Hanza2023mechanics accounted for the effect of radial pressure by proposing an inhomogeneous Cosserat rod model with a nonlinear strain-force relationship.
 The proposed method outperformed the standard CRT by about .
-Alongside, Liu et al.
-(2023) discusses another modelling approach of a pneumatic fiber-reinforced soft actuator with a semicircular cross section.
-Here, the EBRT is employed to explore the effects of self-gravity and external loads on the actuator configuration, to predict its shape.
-Dou et al.
-(2023) introduce a hybrid CSM with pneumatic chambers and an antagonistic rod-driven compliant mechanism enabling elongation, contraction, bending, and stiffness regulation.
-They derive a kinematic model by adapting the EBRT and demonstrate the benefit of the compliant mechanism with an experimental validation comprising way-point tracking.
-Recently, Lamping and de Payrebrune (2023) presented a study on soft robotic actuators that utilize pneumatic actuation with multiple chambers.
-The modeling approach uses the EBRT to characterize the deformation, providing a practical method for accurately determining the rod parameters essential for predicting the performance of these actuators, enhancing their design and control.
-Subclass | References | Rod Model | Contribution |
-Cable | Oliver-Butler et al. (2019) | EBRT/CRT | Stiffness analysis for various tendon stretch and routing paths . |
-Chen et al. (2022) | EBRT | Augments the PCC with a rod model to quickly solve the CSM deformation under . | |
-Pneumatic | Sadati et al. (2017c) | CRT | Utilizes the Ritz-Galerkin method to reduce the continuous state space of rod models. |
-Alessi et al. (2023a) | CRT | Ablation study of model addressing manufacturing uncertainties in a 3D-printed CSM. | |
-Dou et al. (2023) | EBRT | Design and modeling of a hybrid CSM incorporating an antagonistic compliant mechanism. | |
-Smart | Goldberg et al. (2019) | CRT | Simulated locomotion of a caterpillar-inspired soft robot actuated by SMA. |
-4.4.3 Bend & Stretch with Smart Actuators
-Regarding locomotion with smart actuators, Goldberg et al.
-(2019) presents a model for a caterpillar-inspired soft robot actuated by SMA, leveraging the work by (Bergou et al., 2008) related to the elastic rod theory and the Lagrange motion equation for systems of particles.
-A planar version of Bergou’s formulation is employed here to model the dynamics of a single SMA actuator.
-The actuator stiffness and damping parameters are first estimated and subsequently used to simulate the locomotion of the caterpillar robot fabricated from a series of SMA actuators.
-Interestingly, increasing the actuator stiffness by yields simulation displacements higher than the prototype by .
-4.4.4 Bend & Stretch in Parallel Manipulators
-The CRT also aided the design, statics modeling, and workspace analysis of an extensible rod-driven PCM with five DoFs at the tip (Wu and Shi, 2022a).
-A variant of the Levenberg-Marquardt method makes the PCM model suitable for real-time applications.
-Open-loop errors on the tip position in tracking tasks are lower than in PCC models.
-Also, they propose a numerical procedure for efficiently estimating partial workspace using a ROM.
+Meanwhile, LiuShengkai2023Modeling explored the effects of self-gravity and external loads on the configuration of an actuator with a semicircular cross-section through an EBRT.
 6.5 Bend & Twist & Stretch
-Five contributions explored bending, twisting, and stretching combined.
-For cable-driven actuation, it was presented a general geometrically exact static CRT for CSM (Renda and Laschi, 2012).
-The work highlights the role of torsion deformation.
-The variety of the design parameters allows the simulation of a wide range of continuum manipulators and the coupled tendon drive for multiple sections.
-Similarly, a cable-driven conic CSM made of silicone was modeled by combining a geometrically exact dynamic Cosserat rod and the Kelvin model of viscoelastic materials (Wang et al., 2016).
-An experiment with the arm fixed at the base and oriented horizontally, subject to gravity, without actuation (), validates the viscoelastic model, which matches the experimental data better than a pure elastic model.
-Finally, the validation executed 2D and 3D motions actuating different cable combinations ().
-Also, a discrete Cosserat rod combined with a FEM described the dynamics of cable-driven soft robots (Adagolodjo et al., 2021).
-In particular, they use FEM to model the robot structure and Cosserat for the robot actuation.
-The implicit backward Euler time stepper integrated the differential equations for numerical stability.
-Direct and inverse simulations demonstrated the suitability of the approach for modeling robots actuated by cables or rods.
-Aimed at colonoscopy, the CRT was employed to reconstruct the shape of a continuum endoscope given the force exerted on the distal end (Ryu et al., 2018).
-The experimental validation shows that longer endoscopes produce higher tip errors.
-Also, Cosserat rods modeled a six-link PCM (Ghafoori and Khalaji, 2020).
-The model, initially validated for a single rod obtaining an average tip position error of , is applied to a six-link robot with good performance.
-In a recent work (Roshanfar et al., 2023), the authors modelled and experimentally validated a soft robot with combined pneumatic and cable actuation.
-The robot is modelled using CRT where the tangent elastic modulus is expressed as a function of the internal pressure.
-In addition, the hyperelastic constitutive model was integrated into the Cosserat framework.
+A few works explored the combination of bending, twisting, and stretching, especially in cable-driven systems.
+renda2012general presented a general geometrically exact static CRT for cable-driven CSM.
+The work contributed with:
+The simulation of coupled tendon drive for multiple sections, enabled by various design parameters.
+Highlights of the role of torsion deformation.
+Although the study was conducted in simulation only, it provided a solid foundation for future full dynamic analyses (Sec.
+6.2).
+Similarly, wang2016three modeled a cable-driven conic CSM made of silicone by combining a geometrically exact dynamic Cosserat rod and the Kelvin model of viscoelastic materials.
+The experiments with real-world data included:
+Validation of the viscoelastic model with the arm fixed in a cantilever fashion, subject to gravity, without actuation ().
+Execution of 2D and 3D motions actuating different cable combinations ().
+They found that the Kelvin model matched the experimental data better than a pure elastic model. Also, adagolodjo2021coupling described the dynamics of cable-driven soft robots combining
+FEM to model the robot structure, and
+a discrete Cosserat rod to model the robot actuation.
+The implicit backward Euler time stepper integrated the differential equations for numerical stability, while direct and inverse simulations demonstrated the suitability of the approach for modeling robots actuated by cables or rods.
+Recently, Roshanfar2023Hyperelastic investigated stiffness regulation on a cable-driven CSM equipped with a central linear pneumatic chamber.
+They captured robot motion and stiffening behavior by augmenting the classic Cosserat rod with the hyperelastic material model.
+In conclusion, while the Bend & Twist & Stretch class comprises fewer works than the top four deformation classes, it has consistent contributions to cable-driven systems.
+This class could be considered a special case of the All Deformation, where the shearing strain was negligible and neglected.
+Other works in this class are reported in Table 2.
 6.6 Bend & Stretch & Shear
-Two works combine bending, stretching and shearing.
-For multi-backbone continuum robots, it was proposed a variable curvature statics-kinematic Cosserat rod model incorporating multi-backbone structural constraints (Chen et al., 2021).
+The Bend & Stretch & Shear deformation class, similar to the previous class, emerges by neglecting torsion.
+chen2021variable proposed a variable curvature static-kinematic Cosserat rod model for multi-backbone continuum robots, incorporating multi-backbone structural constraints.
 The model is validated on a robot with two multi-backbone continuum segments, each with four nitinol rods.
-The push-pull actuation bends the continuum robot.
+The push-pull actuation bends the continuum robot and transmits shear forces.
 The experiments evaluate the tip positioning accuracy for circular trajectories and the shape discrepancy when the robot is subject to an end-point load in two configurations.
@@
 The proposed model outperforms the PCC.
-Differently, a hybrid model combining the classical CRT with a data-driven stiffness estimation was developed (Bartholdt et al., 2021).
+Note that the torsion-free assumption did not hold in all experiments.
+Indeed, when out-of-plane external forces were applied to cause torsion, the tip position error raised to 19% of the length.
+Differently, bartholdt2021parameter developed a hybrid model combining the classical CRT with a data-driven stiffness estimation.
 Rather than solving numerically nonlinear constitutive equations, the model relies on a work point-dependent linear stress-strain relationship for computational efficiency.
-Experimental validation is carried out by identifying the stretching and bending stiffness of a soft pneumatic actuator and learning the actuation to stiffness maps using artificial neural networks.
+Experimental validation is carried out by identifying the stretching and bending stiffness of a soft pneumatic actuator and learning the actuation to stiffness maps using ANN.
 6.7 Bend & Twist & Shear
-The only work that performed an experimental validation for the compound class of bending, torsion, and shearing is by Grazioso et al.
-(2019), who proposes a novel FEM-inspired spatial discretization technique.
-In particular, the rod is divided into several nodes, defining a helicoidal shape function for the interpolation.
-Thanks to the specific choice of the shape function, the Cosserat rod dynamics was written with Lie Groups.
-The method, implemented in a new simulator called SimSOFT, is validated on the pure bending of a cantilever soft arm and the pure in-plane rotation of a soft arm with varying external conditions.
-In addition, they performed an experimental validation on the Princeton arm to benchmark the coupling of bending, torsion, and shear.
-The validation results report excellent performance in terms of accuracy and computational efficiency.
-Bentley et al.
-(2023) employ a Cosserat string formulation to derive a tissue and needle force model.
+Another special case is when robots bend, twist, and shear, without stretching.
+In this context, grazioso2019geometrically proposed a novel FEM-inspired spatial discretization technique where the rod is divided into several nodes, defining a helicoidal shape function for the interpolation.
+Thanks to the specific choice of the shape function, they formulated the Cosserat rod dynamics with Lie Groups.
+The proposed method, also implemented in a new simulator called SimSOFT, was validated on the pure bending of a cantilever soft arm and the pure in-plane rotation of a soft arm with varying external conditions.
+In addition, they experimentally validated the coupling between bending, torsion, and shear on the Princeton arm benchmark, achieving excellent accuracy and computational efficiency.
+bentley2023safer employ a Cosserat string formulation to derive a tissue and needle force model.
 It describes normal and frictional forces along the shaft as a function of the planned needle path, friction model and parameters, and piercing force.
@@
 6.8 Stretch & Shear
-Combined stretching and shearing deformations are characteristic of peristaltic locomotion (Fig.
-5).
-The peristaltic locomotion of earthworms was studied with a continuous model of compressible and incompressible slender bodies represented as Cosserat rods (Hemingway and O’Reilly, 2021).
+Combined stretching and shearing deformations are characteristic of peristaltic locomotion.
+The peristaltic locomotion of earthworms was studied with a continuous model of compressible and incompressible slender bodies represented as Cosserat rods (hemingway2021continuous).
 Incompressibility is enforced as an internal constraint using Green and Naghdi’s theory of a directed rod.
@@
 The stretching deformation successfully simulated peristaltic locomotion.
-The forward motion results from wave propagation through contraction and extension of the robot segments along the movement direction (Fig.
-5).
-With bioinspired robotics applications, Nguewou-Hyousse et al.
-(2019) modeled a caterpillar soft robot using a network of linear and torsional springs and massless rods, modeled as planar discrete elastic rods (Goldberg et al., 2019).
-4.10 Discussion on rod models
-The analyzed works demonstrate a steady interest in rod models for continuum and soft robots. As shown in Fig. 6, the number of published papers on rod-based models increases.
-4.10.1 On the Deformations
+The forward motion results from wave propagation through contraction and extension of the robot segments along the movement direction.
+With bioinspired robotics applications, Hyousse2019distributed modeled a caterpillar soft robot using a network of linear and torsional springs and massless rods, modeled as planar discrete elastic rods (goldberg2019planar).
+7 Model-based Controllers for Rod Models
+Controlling continuum and soft robots can be challenging due to their intrinsic features.
+Material nonlinearities (e.g., hysteresis), the infinite-dimensional configuration space, and the under-actuated nature of the system make the direct implementation of well-known controllers for rigid manipulators difficult or impossible.
+The control literature for rigid robots primarily focuses on controlling the end-effector in terms of pose, velocity, or force.
+For continuum robots, the natural equivalent task is tip control.
+To exploit the continuum nature of soft mechanisms and achieve tasks that are impossible for their rigid counterparts, the researchers proposed two other tasks:
+Multi-Point (MP) control, to regulate the pose of a set of cross-sections,
+Shape control, to regulate the robot backbone, fully utilizing the potential of continuum structures.
+To tackle the inherent properties of continuum robots, rod theories have been widely used for control design, resulting in an increasing trend of works over the last decade.
+These controllers generally fall into two categories: model-based and learning-based.
+This section deals with MB approaches, while the Sec.
+8 covers the learning-based.
+The MB control strategy involves leveraging prior knowledge of the robot to guide it toward achieving a specific task.
+The motivation behind developing MB controllers is the possibility of mathematically guaranteeing the stability and performance of the controlled system.
+However, implementing controllers in physical prototypes can be challenging, often requiring an extensive identification procedure.
+This context sees the spatial discretization technique as crucial to balance accuracy and computational efficiency.
+Therefore, the model selection and the discretization technique are integral parts of the control design process.
+To formalize the control problem, let us consider a generic configuration vector, whatever discretization algorithm is used.
+According to della2023model, the shape control problem can be particularized in two sub-problems: (i) regulation and (ii) trajectory tracking (TT).
+In the last case, the goal of shape control is to find an input such that
+where are the desired configuration and its time derivative, respectively.
+The condition (42) can be particularized in the regulation case, imposing .
+Similarly to rigid robots, it is possible to formalize the tip control problem, rewriting (42) in the task space, such as
+where and are the desired pose and velocity twist of the tip.
+Similarly, the tip pose regulation condition can be obtained, particularizing (43) with .
+Lastly, the MP control problem can be formalized by extending (43) in a set of desired poses and velocity twists of the cross sections.
+To solve these control problems, it is possible to represent the MB control framework in a general scheme, shown in Figure 5.
+The block “Sensors” refers to the sensors mounted on the robotic platform, mapping in the measurements .
+The “Estimator” block implements the Shape Estimation/Reconstruction algorithms, providing an estimation of , , defined as , .
+Finally, the MB controller block computes the input , processing the error between the estimated configuration and the desired one.
+Following this general scheme, many works on MB control occurred, thanks to the development of accurate and computationally efficient rod models.
+Below, we review the contributions, while Table 3 reports the key points.
+7.1 Collocated Form
+To address the underactuated nature of continuum robots, a recent method proposed by pustina2024input reformulates the dynamic model (14) in a collocated form.
+This approach decouples the actuated variables from the passive ones via a change of coordinates, .
+The resulting state vector is partitioned into actuated coordinates, , and passive coordinates, .
+A valid change of coordinates can be expressed as
+where is the coordinate transformation in the actuated coordinates .
+For continuum robots modeled with CRT, has a clear physical interpretation.
+The actuated coordinates correspond to the lengths of the individual actuators, such as
+where denotes the length of the -th actuator.
+The change of coordinate (45) is significant because it associates a subset of the generalized coordinates with a physically meaningful quantity that can be easily measured (pustina2024input).
+For example, for a tendon-driven CSM, the actuated variables can be directly obtained using encoders.
+In these collocated coordinates, the EoM take the partially decoupled form, such as
+where and are the inertial, Coriolis, gravitational, stiffness, and damping matrices, respectively, expressed in the collocated coordinates .
+Adopting this change of coordinates is promising for facilitating the transfer of control methods developed for underactuated rigid systems.
+For instance, pustina2024input demonstrates the application of classical controllers—such as PD with gravity and elasticity compensation (borja2022energy; borja2022role) and its variant P-satI-D (pustina2022p), a PID controller with integral saturation—directly to the collocated variables.
+Finally, this formulation not only enhances the understanding of the system dynamics and simplifies control design, but also enables the design of linear controllers in the collocated coordinates without approximations.
+Notably, the control laws remain nonlinear due to the inherent nonlinearity of the mapping .
+7.2 Control by Model Inversion
+The most direct approach to controlling a continuum robot is through its kinematic model.
+This strategy, known as Inverse Kinematics (IK), computes the actuation inputs required to drive the robot’s tip to a desired pose.
+Early studies demonstrated effective planar trajectory tracking and vibration suppression using simplified models such as EBRT (ataka2020model; gravagne2003Large).
+To better handle real hardware complexities, these models were extended with online observers, which compensate for phenomena such as variable stiffness caused by changes in inflation pressure (ataka2020model).
+To achieve higher accuracy and address complex spatial deformations, recent work has shifted towards the more comprehensive CRT.
+This framework enables advanced capabilities, including micrometer-level positioning in parallel continuum robots (mauze2021micrometer) and multi-point orientation control along a manipulator’s body (richter2021multi).
+The accuracy of CRT comes at a high computational cost, limiting real-time use.
+Researchers have therefore turned to quasi-static formulations and actuator nonlinearity compensation to improve efficiency (campisano2021closed).
+These kinematic models are continuously being generalized to accommodate more complex physical designs, including manipulators with arbitrary, discontinuous, or overlapping actuator routings (renda2022geometrically).
+While kinematic models determine the robot’s shape, dynamic models are required for high-speed maneuvers and interactions with the environment.
+Controllers based on Inverse Dynamics (ID) account for forces such as inertia, gravity, and external contacts.
+Such methods achieved full shape control, with some approaches combining geometric methods with EBRT to balance accuracy and computational speed (mbakop2021inverse).
+Because dynamic models are computationally demanding, recent work explored decentralized control architectures that distribute computation across modules, improving real-time feasibility (doroudchi2021configuration).
+7.3 Managing Nonlinearities and Under-actuation
+A central challenge in continuum robotics arises from the strong nonlinearities and inherent under-actuation of these systems.
+Feedback Linearization (FL) is a powerful technique that cancels nonlinear terms, transforming the system dynamics into a simpler, linear form.
+FL was successfully applied in cascade control architectures to compensate for gravity and bending effects, thereby improving tracking performance (falkenhahn2016dynamic).
+However, its effectiveness depends heavily on model accuracy.
+To address this limitation, rucker2022task proposed a hybrid approach that combines an inner-loop FL controller with a robust outer-loop strategy such as Sliding Mode Control (SMC), enabling robustness against model uncertainties (rucker2022task).
+Practical aspects also matter: research shows that the choice of discretization scheme can significantly influence numerical robustness, computational cost, and accuracy (caradonna2024model).
+An alternative approach is the Energy-Shaping Control (ESC).
+Instead of canceling dynamics, ESC reshapes the system’s potential energy such that the desired configuration becomes a stable energetic minimum.
+By reformulating the dynamics in a Port–Hamiltonian framework, this method can formally guarantee stability and passivity, allowing control of both under-actuated and hyper-redundant systems (caasenbrood2021energy; caasenbrood2022energy).
+ESC often produces smooth, natural motions that exploit the robot’s intrinsic compliance.
+It was applied even to complex bio-inspired systems such as octopus tentacles, where optimal control was used to identify shaping laws for advanced models that include non-rigid cross-sections (chang2023energy).
+Discrete-time control is necessary for practical implementations, even though most of the ESC literature concentrates on continuous-time models.
+Structure-preserving numerical techniques, such as Lie group variational integrators, are used to maintain the stability and passivity guarantees in a discretized system (tiwari2023discrete).
+These integrators allow for efficient stabilization of flexible beams while preserving the geometric structure of the manipulator’s configuration space.
+7.4 Robust and Optimal Control
+Robust controllers for continuum robots were designed to mitigate model uncertainties.
+SMC is a widely adopted strategy for ensuring robustness against model uncertainties and external disturbances.
+Adaptive variants of SMC were developed to handle limitations such as actuator saturation while guaranteeing marginal stability (alqumsan2019robust).
+A classical drawback of SMC is the high-frequency “chattering” in the control signal.
+This issue was mitigated by integrating fuzzy logic to smooth the discontinuous control law, leading to smoother actuation profiles (li2021modeling).
+The methodology continues to evolve with advanced variants, such as terminal SMC, which achieves faster finite-time convergence.
+It was combined with algorithms for high-level tasks such as obstacle avoidance (mishra2023trajectory).
+When performance is critical and constraints must be explicitly respected, Model Predictive Control (MPC) is the method of choice.
+MPC solves a finite-horizon optimization problem at each time step, enabling it to naturally enforce input bounds while optimizing complex objectives.
+This capability makes it particularly suitable for medical applications, such as catheter control, where the robot must apply a constant force to tissue while tracking a trajectory (soltani2017soft).
+MPC is also highly effective for motion planning, as it can generate the complex, oscillatory actuation signals required for behaviors like fish-like swimming, even under disturbances (barbosa2023motion).
+However, the limitation of MPC remains its computational cost, which often necessitates model simplifications (e.g., linearization) for real-time implementation.
+Beyond traditional methods, emerging paradigms include optimal control for bio-inspired motions (wang2021optimal), decentralized robust control for modular architectures (doroudchi2018decentralized), and synergetic control (hashemi2023robust), which has recently shown robustness to disturbances in continuum robots control.
+| Controller | References | Rod Model | Task | Validation | ||||||||||||||||||||
+| Inverse Kinematics |
+| Inverse Dynamics |
+| Feedback Linearization |
+| Energy-Shaping |
+| Sliding Mode |
+| Model Predictive Control |
+| Open-Loop Optimal Controller | wang2021optimal | CRT (DER) | Reaching, Fetching | Numerical | ||||||||||||||||||||
+| doroudchi2018decentralized | EBRT | Shape Regulation | Numerical | |||||||||||||||||||||
+| Synergetic Control | hashemi2023robust | CRT | Tip Position TT | Numerical |
+7.5 Shape/State Estimation
+A general MB shape controller requires feedback in terms of configuration vector and its time derivative (Figure 5).
+Depending on the spatial discretization, the vector could be difficult or impossible to measure.
+To tackle this issue, a general MB control framework requires an Estimator, which maps the measures to the estimated state .
+For continuum and soft robots, this problem is referred to as Shape/State Estimation or Shape/State Reconstruction.
+Shape Estimation methods can use geometrical models or Computer Vision algorithms (bezawada2022shape; albeladi2021vision).
+However, the following works show that incorporating prior knowledge of rod-based models can significantly enhance performance.
+The pioneering work of trivedi2014model suggests three approaches for Shape Estimation, based on CRT: (i) mounting load cells at the base, (ii) employing cable encoders, and (iii) mounting an inclinometer at the end of each piece.
+A maximum tip position error of is observed in simulations.
+All three approaches were then experimentally validated on the OctArm VI (grissom2006design).
+Another study by donat2021real proposes a real-time shape-estimation method based on the force-torque measured at the basis of the tubes for a CTR.
+It extends a shape estimation algorithm for elastic Kirchhoff rods.
+They modeled a CTR combining planar PCC segments lying on different equilibrium planes.
+The approach is evaluated with single and two combined additively manufactured tubes at high frequency, achieving a mean deviation of mm along the tube.
+Differently, lilge2022continuum developed a model for shape and strain estimation of a continuum cable-driven robot using a Gaussian Process regression to estimate continuous trajectories in .
+The idea involves substituting (i) time with arc-length and (ii) kinematic and dynamic laws based on CRT.
+This method efficiently estimates the robot’s shape using noisy measurements from sensors such as strain gauges, tracking coils, and an external camera.
+The real-world experiments provide excellent performances, with an average tip pose error of mm and .
+In parallel, aloi2022estimating developed a model-based framework for estimating distributed contact forces directly from shape measurements.
+The method involved fitting a Cosserat rod kinetostatic model with Gaussian load parameterization and solving it via an Extended Kalman Filter, enabling real-time, sensorless force estimation with sub-Newton accuracy.
+State estimation under uncertainty was also addressed in reconfigurable parallel continuum systems (anderson2017continuum), applying Cosserat rod modeling, multi-point constraints, and advanced filtering to improve tip accuracy and enable load inference.
+ferguson2024unified presents a surgical continuum manipulator with integrated FBG sensing for real-time 3D shape reconstruction, enabling sub-millimeter navigation without radiation in complex surgical settings.
+shi2016shape provided a comprehensive review of shape-sensing modalities for surgical continuum robots, highlighting complementary strengths of electromagnetic, optical, FBG, and vision-based methods.
+On the control side, underactuated discrete rod models were leveraged to achieve precise task-space trajectory tracking through FL and robust sliding-mode strategies, all while relying on minimal sensing (rucker2022task).
+In addition, yousefi2023model propose an optimization-based method that simultaneously estimates the shape and the forces acting on a continuum robot by employing the quasi-static CRT.
+Magnetic localization determines the position of multiple robot points.
+The method estimates the robot’s shape and force in a wide range of conditions validated experimentally, with and without knowledge of the contact positions.
+Recently, zheng2024full; zheng2024estimating proposed the Cosserat theoretic boundary observer, a mechanics-based dynamic state estimation algorithm to recover the infinite-dimensional robot states by measuring the tip velocity twist and the tip pose.
+Finally, feliu2024dynamic presents a novel method for estimating the dynamic state (pose and velocity) of tendon-actuated CSM.
+The approach uniquely combines a GVS Cosserat rod model with a nonlinear State-Dependent Riccati Equation observer.
+A key advantage of this technique is its ability to estimate the state using only internal sensor data, such as tendon displacements and actuator forces.
+The method was validated both in simulation and on a prototype, achieving an average dynamic tip position error of 1.79 cm, which is about 7%L.
+8 Learning-based Controllers for Rod Models
+ML methods are becoming a trend to control CSM (george2018control; wang2021survey; wang2022control; falotico2024learning).
+Learning-based controllers leverage the physics of rod models and utilize ML to derive parameterized control policies , which provide and end-to-end mapping between desired poses and sensor measurements to motor commands :
+The vector of policy parameters represents the weights of ANN, which can be trained with Supervised Learning (SL) or RL.
+We report learning-based controllers in chronological order to highlight the evolution of this growing field.
+8.1 Supervised Learning
+SL in simulation is a viable approach to address simple tasks with soft robots.
+It relies on pseudo-random motion data of actuations and corresponding task space to train an inverse model (47).
+For instance, thuruthel2017learning developed dynamic models of CSM using recurrent neural networks and presented a trajectory optimization method for task space control.
+They validated the controller on a Cosserat rod with PCS parameterization and a pneumatic CSM.
+Moreover, wang2022data developed a closed-loop controller of continuum robots, employing RNN to approximate the forward and inverse dynamics.
+They extended the model in (till2019real) to account for spine compression and collect motion data for training their forward models.
+A non-parametric Gaussian process regression compensates for the error between the real robot and the RNN.
+Then, a hybrid controller alternates between two stand-alone policies.
+They validated the control architecture on reaching and tracking tasks using a tendon-driven continuum robot.
+Control accuracy improved by combining simulated and experimental data.
+Recently, generative models were employed for the first time to achieve domain translation on a soft robot, supported by a Cosserat rod model (kushawaha2025domain).
+However, this was validated only in simulation for simple tracking tasks and required a substantial computational effort.
+8.2 Reinforcement Learning
+RL is one of the main ML paradigms and a versatile framework suited for decision-making and control (sutton2018reinforcement).
+The two fundamental objects of an RL problem are the agent and the environment.
+In soft robotic control, the agent is a learning-based controller implementing an ANN, whereas the environment is everything outside the agent, comprising the soft robot and other objects.
+Figure 6 depicts the agent-environment interaction highlighting the role of ANN and rod models.
+The agent and the environment interact at discrete time steps .
+At every interaction step, the agent observes the (possibly partial) state of the environment including sensor readings and task information.
+The agent acts according to a policy to select actions taking the form of low-level controls like torques or pressures.
+A policy can be a deterministic mapping from states to actions, .
+Alternatively, a policy can be stochastic, mapping a state to a distribution over actions .
+The environment transitions into another state according to the dynamics of the rod model.
+The agent also perceives a scalar reward signal from the environment, indicating the goodness of the current state specifying the control task.
+The goal of the agent is to maximize its return, the cumulative sum of rewards:
+where is a discount factor. Therefore, the goal of RL is to find an optimal policy that maximizes the expected return
+We review the following deep RL contributions in chronological order.
+These are also summarized in Table 4.
+satheeshbabu2019open applied deep Q-learning (DQL) with experience replay for the quasi-static position control of a pneumatic CSM that can bend and twist.
+The policy was open-loop and trained on a simulation of a static Cosserat rod.
+They validate it in simulation and on the physical platform, subject to various external loads.
+DQL can cope with the challenging state-action space of soft robots by employing an ANN to approximate the value function of state-action pairs (mnih2013playing), which is defined as
+Then, a deterministic policy can leverage to select the best action in a state as .
+The work was extended by increasing the dexterity of the CSM and developing a closed-loop policy, still for precise quasi-static positioning, via deep deterministic policy gradient (DDPG) (satheeshbabu2020continuous).
+They investigated the robustness of the control policy to loading and workspace discontinuity, and deployed the policy on hardware.
+The DDPG algorithm also derived a closed-loop controller for the quasi-static positioning of a pneumatic CSM modeled with the KRT (uppalapati2020berry).
+Interestingly, the CSM was integrated into a mobile platform including a rigid arm and a sensorized gripper.
+The system was teleoperated and validated in the agricultural task of picking berries using different maneuvering strategies.
+The potential of continuum mechanical models is fully exploited in physical interaction tasks.
+naughton2021elastica effectively applied several deep RL algorithms to train various control policies.
+They explored point reaching, trajectory tracking, and maneuvering through structured and unstructured obstacles.
+The controllers were learned and tested in simulation using a synthetic CSM based on a dynamic CRT.
+The rod was actuated by applying distributed internal torques, modeled via splines characterized by control points and vanishing values at the rod ends.
+The work showcased a captivating interaction between the rod and the environment.
+In particular, the maneuvering task required the free end of the rod to reach a position behind fixed obstacles.
+The task was solved with a reward function that only considered the distance to the target and a penalty term for the occurrence of numerical errors
+Interestingly, the rod learned to navigate and interact with the obstacles without an explicit reward term.
+The generalization capabilities of learned controllers are vital for successful deployment in the real world.
+In this regard, alessi2023learning applied proximal policy optimization (PPO) to train a closed-loop position control policy for dynamic trajectory tracking with a pneumatic CSM.
+The controller was trained leveraging a dynamic Cosserat rod model of the soft robot (alessi2023ablation).
+After training, four simulation tests evaluated how the policy generalized to new observations, dynamics, and tasks.
+The experiments included tracking new trajectories subject to unknown external forces or using different material properties (e.g., ).
+The policy generalized well within some boundaries and also transferred without retraining to intercept a moving object.
+Subsequently, alessi2024pushing extended the simulated environment to account for the dynamics of a dexterous CSM and its interaction with objects.
+The proposed closed-loop control architecture for pose/force control enabled dynamic pushing with CSM in the real world.
+They mitigated the significant sim-to-real gap in soft robotics with a novel adaptation of domain randomization (DR) (zhao2020sim) that emphasizes the randomization of soft material properties.
+This was the first sim-to-real transfer of a manipulation policy based on rod models.
+| Controller | References |
+Task |
+| Deep Q-Learning | satheeshbabu2019open | CRT (✗) | Tip Position TT (way-point) | ✓(None) | ||||
+| Deep Deterministic Policy Gradient | satheeshbabu2020continuous | CRT (✗) | Tip Position TT (way-point) | ✓(None) | ||||
+| uppalapati2020berry | KRT (✗) | Tip Position TT (way-point) | ✓(None) | |||||
+| Trust Region Policy Optimization | naughton2021elastica | CRT (✓) | Tip Pose TT, Obstacle maneuvering | ✗ | ||||
+| Proximal Policy Optimization | alessi2023learning | CRT (✓) | Tip Position TT, Object interception | ✗ | ||||
+| alessi2024pushing | CRT (✓) | Pushing (Tip Pose/Force) | ✓(DR) |
+9 Discussion and Emerging Challenges
+9.1 On the Rod Theories
+Following our mathematical treatment and review of rod models, we briefly reevaluate the strengths and limitations of the various rod theories within soft robotics.
+The Euler-Bernoulli model is simple, computationally light, and effective for slender, stiff structures subject to small bending deflections.
+Its nonlinear version makes it suitable for large deformations.
+However, it remains a planar model and neglects shear as the cross-sections remain perpendicular to the centerline.
+This assumption makes it unsuitable for modeling interactions with the environment in soft robotics.
+Nonetheless, some works extended its applicability beyond bending by coupling it with other models.
+The Timoshenko model improves on this by including shear deformation.
+However, Timoshenko is less suited to nonlinear actuator paths that induce torsion.
+To this end, the Kirchhoff–Love theory handles large rotations and displacements but assumes the rod is inextensible and unshearable.
+This assumption limits its ability to model extensible pneumatic actuators.
+Finally, the Cosserat-Reissner model and its extensions are the most general and suited for continuum soft robots.
+They can handle large deformations, shear, torsion, and stretching.
+However, they require more computational effort.
+Overall, each rod theory has its strengths and weaknesses.
+The choice depends on the system’s behavior and the level of detail needed.
+We suggest opting for a Cosserat rod model and tuning the space and time discretization to match the desired tradeoff between efficiency and accuracy.
+9.2 On the Deformations
 Researchers developed numerous rod-based models reformulating and adapting classical rod theories in the soft robotic field.
-Herein, we reviewed diverse models, which described different combinations of strain modes, giving rise to nine deformation classes (Fig.
-4).
-The most common class in literature is Bend, with the of the examined papers.
+Herein, we reviewed diverse models, which described different combinations of strain modes, giving rise to nine deformation classes.
+The most common class in literature is Bend, with about a third of the examined papers.
 This popularity is due to the propensity of slender bodies to bend.
@@
 While this represents a computational surplus for trivial bending structures, the generalization of Cosserat models pays off if the model gets deployed in physical interaction tasks.
-The next class is All Deformation, with the of the examined papers.
+The next class is All Deformation, with about a quarter of the examined papers.
 As expected, only models based on the CRT achieve all deformations (Table 2).
@@
 In addition, the most common actuation sources (i.e., pneumatic and cables) excite elongation and compression, causing a significant change in the dimension workspace.
-The third class accounting for of the investigated studies is Bend & Twist.
-These models mainly describe robots in which the unshearability and the unstretchability approximations hold, as summarized in Table 3.
+The third class accounting for about a fifth of the investigated studies is Bend & Twist.
+These models mainly describe robots in which the unshearability and the unstretchability approximations hold (Table 2).
 Including the twisting deformation allows the representation of backbones with generic 3D curves.
 However, the twisting deformation can be excited only by specific actuator paths (e.g., helicoidal cables) or asymmetric actuators.
-Another significant class, summarized in Table 4, is Bend & Stretch () for the wide diffusion of linear extensible soft pneumatic actuators and tendon-driven robots.
+Another significant class is Bend & Stretch thanks to the wide diffusion of linear extensible soft pneumatic actuators and tendon-driven robots.
 Such actuators extend or compress upon solicitation, which induces the continuum body to stretch and bend thanks to their radial disposition on the cross-section.
 The remaining five categories are less explored.
-While the Bend, Twist & Stretch () is a particular case of the All Deformations class, the other classes comprehend few works.
+While the Bend, Twist & Stretch is a particular case of the All Deformations class, the other classes comprehend fewer works.
 The lack of work is due to the difficulty of exciting these specific strain modes with the current actuation technologies or the small number of applications in which the assumptions are valid.
 In addition, fewer studies that explored shearing per se could be due to the tendency to consider shearing an undesirable or negligible deformation.
-However, shearing could be significant when there is a prominent environment interaction, as observed in peristaltic locomotion (Hemingway and O’Reilly, 2021) or complex manipulation (Grazioso et al., 2019).
-4.10.2 On the Spatial Discretization
-One characterizing feature of rod models is the spatial discretization method, which significantly impacts the model’s accuracy and complexity.
-For simulation and control purposes, the discretized model must be computationally efficient and as accurate as possible.
-In the analyzed literature, exist two main discretization techniques: Strain Parameterization (Sec.
-2.2.3) and Discrete Elastic Rod (DER) (Sec.
-2.2.4).
-The former is based on a discretization directly in the configuration space, allowing the user to choose the degrees of accuracy.
-In addition, it is possible also to neglect specific strain modes to reduce the computational time.
-The latter discretizes the length of the rod, considering also the external forces applied to the nodes.
-Furthermore, most of the existing spatial discretization algorithms are already implemented in dedicated simulators, such as SoRoSim (Mathew et al., 2022) for SP and PyElastica (Naughton et al., 2021) for DER.
+However, shearing could be significant when there is a prominent environment interaction, as observed in peristaltic locomotion (hemingway2021continuous) or complex manipulation (grazioso2019geometrically).
+Shearing also becomes important when the aspect ratio of the CSM becomes less slender and more tick.
+9.3 On the Spatial Discretization
+A characterizing feature of rod models is the spatial discretization method, which significantly impacts the model’s accuracy and complexity.
+For simulation and control purposes, the discrete model must be computationally efficient and as accurate as possible.
+In the analyzed literature, two primary discretization approaches are commonly employed: Discrete Elastic Rod (DER) (Sec.
+4.2.4) and Strain Parameterization (Sec.
+4.2.3).
+The DER method discretizes the rod’s length into nodes connected by rigid links, representing the rod CSM as a set of discrete poses.
+The accuracy of the solution relative to the continuous rod depends directly on the number of nodes.
+In contrast, Strain Parameterization discretizes the rod in the configuration space, constraining the robot’s backbone to a subset of admissible spatial curves.
+The latter approach permits the neglect of specific strain modes without imposing additional constraints on the EoM, as required in the DER formulation, thereby enhancing computational efficiency.
+Furthermore, Strain Parameterization naturally allows the forward dynamics to be expressed in Lagrangian form, which is particularly advantageous for control design.
+Both approaches are widely used and integrated into dedicated simulators, such as SoRoSim (mathew2022sorosim) for Strain Parameterization and PyElastica (naughton2021elastica) for DER.
 Straddling geometric and mechanical models, some examined works proposed to fit discrete points with functions, such as PH or Euler’s spirals, which is useful, especially for Shape Estimation.
-4.10.3 On the Identification and Validation
+9.4 On the Identification and Validation
 Identification and validation of rod models play a crucial role in accurately reproducing the experimental robot behaviors.
@@
 However, the literature needs rigorous mechanical tests to yield more accurate models, especially for unstructured physical interactions with the environment.
+Indeed, despite a careful model calibration on motion data, alessi2023ablation; alessi2024pushing observed real-world buckling during a manipulation task that was not present in simulation.
 Concerning performance metrics, rod models predicted experimental tip positions with Euclidean errors in the range .
@@
 In summary, the wide use of experimental data suggests an increasing need to reduce the reality gap to develop more accurate controllers.
-4.10.4 On the Impact of Materials and Manufacturing
+9.5 On the Impact of Materials and Manufacturing
 Most of the analyzed rod models employed continuum and soft robots built with diverse materials, including traditional elastomers (e.g., silicone, rubber) or other polymers (e.g., SMA, hydrogels).
@@
 To tackle this issue, many analyzed works presented models with a nonlinear constitutive law, using hyperelastic model such as Odgen.
-Manufacturing techniques provided continuum and soft robots with unprecedented dexterity (Wallin et al., 2018).
+Manufacturing techniques provided continuum and soft robots with unprecedented dexterity (wallin20183d).
 However, manufacturing uncertainties may impact model accuracy.
-Recent rod models addressed manufacturing uncertainties considering variations in material (Alessi et al., 2023a) and geometric properties (Eugster et al., 2022).
-Despite these advancements, some gaps persist, particularly in capturing hysteresis effects or integrating the dynamic restoring process of self-healing materials into rod-based models.
-Hysteresis, a prevalent phenomenon in soft materials, strongly impacts the accuracy and predictability of the robot’s behavior.
-Furthermore, incorporating self-healing materials can potentially enhance the durability and longevity of soft robots but remains relatively unexplored in rod theories (Terryn et al., 2017).
-Addressing these challenges could advance the capabilities of continuum and soft robots, enabling their integration into diverse applications and real-world tasks.
-5 Controllers using rod models
-Controlling continuum robots can be very challenging, due to the intrinsic features of this type of robots.
-Material nonlinearities (e.g., hysteresis), the infinite-dimensional configuration space, and the under-actuated nature of the system make the direct implementation of well-known controllers for rigid manipulators difficult or impossible.
-The literature on the control of rigid robots primarily focuses on controlling the end-effector in terms of pose, velocity, or force.
-In the realm of continuum robots, the natural equivalent of this task is tip control.
-To exploit the continuum nature of soft mechanisms and achieve tasks that are impossible for their rigid counterparts, the researchers have proposed two other tasks: Multi-Point (MP) control and shape control.
-The goal of MP control is to control the pose of a set of cross-sections, while shape control aims to regulate the backbone of the robot, fully utilizing the potential of the continuum structures.
-To tackle the inherent properties of continuum robots, rod theories have been widely used for control design, resulting in an increasing trend of works in the last decades, as shown in Fig. 6.
-The developed controllers fall into two different categories: model-based and learning-based.
-5.1 Model-based Controllers
-The Model-Based (MB) control strategy involves leveraging prior knowledge of the robot to guide it toward achieving a specific task.
-The motivation behind developing MB controllers is the possibility of mathematically guaranteeing the stability and performance of the controlled system.
-However, implementing controllers in physical prototypes can be challenging, often requiring an extensive identification procedure.
-This context sees the spatial discretization technique as crucial to balance accuracy and computational efficiency.
-Therefore, the model selection and the discretization technique are integral parts of the control design process.
-5.1.1 Formalization
-To formalize the control problem, let us consider a generic configuration vector, whatever discretization algorithm is used.
-According to Della Santina et al.
-(2023), the shape control problem can be particularized in two sub-problems: (i) regulation and (ii) trajectory tracking (TT).
-In the last case, the goal of shape control is to find an input such that
-where are the desired configuration and its time derivative, respectively. The condition (48) can be particularized in the regulation case, imposing .
-Similarly to rigid robots, it is possible to formalize the tip control problem, rewriting (48) in the task space, such as
-where and are the desired pose and velocity twist of the tip.
-Similarly, the tip pose regulation condition can be obtained, particularizing (49) with .
-Lastly, the MP control problem can be formalized by extending (49) in a set of cross-sections’ desired poses and velocity twists.
-To solve these control problems, it is possible to represent the MB control framework in a general scheme, shown in Fig.
-7.
-The block “Sensors” refers to the sensors mounted on the robotic platform, mapping in the measurements .
-The “Estimator” block implements the Shape Estimation/Reconstruction algorithms, providing an estimation of , , defined as , .
-Finally, the MB controller block computes the input , processing the error between the estimated configuration and the desired one.
-More in-depth, the control input is composed of two actions (Fig.
-7).
-The former is the feedforward action , depending on the chosen model, the estimated state and the desired trajectory.
-The latter is the feedback action that depends on the error w.r.t.
-the desired configuration and its time derivative.
-Finally, the specific control algorithm computes the control output combining the feedforward and feedback terms.
-5.1.2 Inverse Kinematics
-One of the first approaches proposed in the MB literature is the Inverse Kinematics control.
-The main idea is to use the inverse kinematic model to find the actuation input such that the tip reaches a desired pose or velocity twist.
-In the above case, the rod-based kinematic models enable the design of controllers that account for the continuum robots’ deformations.
-In the planar case, the EBRT provides a simple and efficient kinematic model.
-Ataka et al.
-(2020a) proposed a kinematic pose control for an inflatable eversion CSM for trajectory tracking in a planar environment.
-Assuming a pure bending robot, they used EBRT to model the effect of internal inflation pressure on each bending segment of the prototype.
-A further advancement in the modeling is the characterization of the variable structural stiffness caused by changes in inflation pressure.
-Additionally, an observer based on the Extended Kalman Filter estimates the unknown model parameters online from chamber pressure and bending angle measurements.
-These parameters approximate the actuator-space Jacobian matrix online, which controls the CSM’s tip pose under various stiffness conditions.
-Simulation and experimental validation using an inflatable CSM made of unstretchable fabric demonstrated that the proposed method achieved position control in a planar environment despite different robot inflation pressures (i.e., stiffness values).
-Mauzé et al.
-(2021) achieve micrometer positioning accuracy with a PCM made of three planar kinematic chains linked to a rigid moving platform.
-Each chain has a prismatic actuator and a flexible rod that deforms continuously, with bending and twisting deformations modeled with the CRT.
-A regular Cartesian controller based on inverse kinematics is implemented using the parameters identified for the forward kinematic model.
-The controller is validated for a circular and square trajectory.
-For both, large orientation displacements of the platform increase the angular error.
-A closed-loop multi-point orientation control strategy for discretely magnetized continuum manipulators was developed (Richter et al., 2021).
-A non-homogeneous magnetic field model conforming to Maxwell’s equations was formulated and applied to the actuation system to predict the exerted wrenches on the manipulator.
-The approach is demonstrated on a manipulator containing two permanent magnets actuated inside a non-homogeneous magnetic field.
-Here, a quasi-static CRT models the manipulator deformation under external wrenches, while a Jacobian provides the actuation inputs.
-Experiments in simulation and on a real robot demonstrate that the algorithm independently controlled the orientation of two interconnected magnets in a non-uniform magnetic field.
-However, the control frequency is limited to Hz due to the computational cost of shape reconstruction and the Jacobian.
-Campisano et al.
-(2021) propose a controller for CSM under tip follower actuation.
-In particular, the strategy is to couple a quasi-static regime and sensor feedback of the actuation sources.
-This approach achieves (i) rapid computation of the Cosserat rod model, (ii) actuator nonlinearities compensation, and (iii) oscillation reduction.
-In conclusion, the closed-loop control algorithm is validated experimentally and compared with an open-loop one, to show the effectiveness of the design process.
-Renda et al.
-(2022) proposed an inverse kinematics controller based on the GVS model, using the actuators-strain mapping in (39).
-The control algorithm is designed for trajectory tracking of a set of points in .
-The kinematic model is extended, including a map between the length of threadlike actuators and the configuration of the CSM.
-This formulation covers the general case of actuators’ routing, such as discontinuous or overlapping routing.
-The control algorithm is validated through simulations on a CSM with parallel and helical actuators’ routing.
-5.1.3 Inverse Dynamics
-To exploit the mechanical impedance of continuum robots, the controller requires a knowledge of the dynamics of the elastic rod.
-Furthermore, the dynamic models allow us to consider the influence of external forces, such as gravity or contacts.
-Therefore, an inverse dynamic model can be used, compensating undesired nonlinear effects and disturbances.
-In the case of shape control, Mbakop et al.
-(2021a) proposed an inverse dynamics controller for a soft finger.
-The control algorithm is based on a mix of a geometrical and mechanical approach, combining the PH and the Euler-Bernoulli model.
-While the former allows the reconstruction of the finger’s shape, the latter maps actuators’ magnitude to the positions of PH control points.
-They experimentally validated the control law with two prototypes: (i) a planar Fluidic Elastomeric Actuator soft finger and a pneumatic CSM.
-Applying an external tip load to the CSM proved that combining the PH and EBRT improves accuracy and the computational time, w.r.t.
-using only the PH model.
-The subsequent simulation study proposed a new decentralized Inverse Dynamic controller based on the CRT (Doroudchi and Berman, 2021).
-The controller considers all strain modes and incorporates the system dynamics inside different gain matrices.
-Each piece or unit of the soft arm is active and independent from the other pieces.
-This approach increases the computational efficiency and successfully tracks configurations requiring shear deformations with average normalized root-mean-square errors below .
-5.1.4 Feedback Linearization
-Continuum robots have an inherent under-actuation, which makes control design more difficult.
-Feedback Linearization (FL) is a widely used method for controlling under-actuated mechanisms (Spong, 1994).
-The main idea of this control algorithm is to linearize the system with a feedback control action, compensating all the nonlinear terms.
-The FL controller can be combined with a linear controller (e.g.
-Linear Quadratic Regulator), in order to achieve the desired performances.
-For a task-space control, Rucker et al.
-(2022) proposed a controller for continuum robots based on a highly under-actuated planar Kirchhoff rod (i.e., the EBRT).
-The main idea is to show that a maximal coordinate form of the Kirchhoff model is suitable for control purposes, exploiting the under-actuation of the system.
-The controller is derived by solving a constrained Lagrangian problem, discretizing the length of the Kirchhoff rod.
-The control algorithm is based on an input-output Feedback Linearization combined with an Sliding Mode Control (SMC) in an outer loop.
-An observer based on the constrained model formulation provides control feedback, adding corrective virtual external forces based on the estimation error.
-Lastly, simulations of an actual two-DoFs soft prototype evaluate the control framework and demonstrate the consistency of the algorithms during tip position trajectory tracking.
-The numerical examples of the SMC also demonstrate robustness to parameter variations.
-Caradonna et al.
-(2024) presented a 2D soft segment with a revolute joint at its base controlled by a FL controller.
-By analyzing the equilibrium map of the system’s zero dynamics, they proposed a method to employ the system for the pick-and-place task, accounting for the added mass from the object.
-They utilize the FL controller with the Curvature Parameterization (27) and Strain Parameterization (11) and compare the two discretization techniques emphasizing numerical robustness, computational cost, and accuracy.
-The control algorithm and pick-and-place strategy validated through simulations demonstrated consistency.
-5.1.5 Energy-Shaping
-Another well-known control algorithm designed for under-actuated systems is the Energy-Shaping Control (ESC) (Spong, 1996).
-The main idea is to find an input such that the system reaches a desired energy state.
-The rod models are particularly effective in this case, thanks to the Hamiltonian formulation of the elastic rods.
-Caasenbrood et al.
-(2021, 2022) proposed an ESC for CSM for tip pose regulation.
-To apply the control algorithm, the Authors reformulated the GVS model in a Port-Hamiltonian form, showing the passivity condition of the system and enabling the design of the ESC.
-The control law is validated through numerical examples, proving the capability of the controlled system to reach the desired tip pose.
-Furthermore, thanks to the properties of the specific control algorithm, both under-actuation and hyper-redundancy are effectively managed.
-The controlled system demonstrates smooth convergence to the target, and a pseudo-joint strategy similar to that observed in nature.
-Chang et al.
-(2023) introduced an ESC for a CSM inspired by an octopus tentacle employing the CRT with the DER discretization.
-They modeled three main types of muscles: transverse, longitudinal, and oblique.
-The transverse muscles necessitate the DER approach due to the invalid assumption of a rigid cross-section.
-Leveraging the system’s passivity properties, they designed an ESC by solving an optimal control problem based on Hamiltonian principles.
-Additionally, they developed an efficient method to design a task-specific equilibrium configuration by solving an optimization problem in the group.
-The model and ESC are validated on the PyElastica simulator, controlling the CSM for reaching and grasping a static object.
-Tiwari and Banavar (2023) proposed a ESC for a 2D flexible single-link manipulator, based on CRT.
-In particular, the model used is geometrically exact, i.e.
-the manipulator’s configuration space lies on an infinite-dimensional Lie Group.
-To implement the controller in the discrete-time domain, the Authors used a lie group variational integrator, derived from the Lagrangian variational principle of the flexible single-link manipulator (Tiwari and Banavar, 2023).
-The results in silico indicate that the discrete closed-loop system achieves shape regulation during point-to-point rotational maneuvers.
-5.1.6 Sliding Mode
-An alternative to the above controllers for under-actuated systems is the SMC (Xu and Ümit Özgüner, 2008), due to the inherent robustness.
-The main idea of this strategy is to control the nonlinear system with a discontinuous control signal, forcing it to a specific curve on the phase space (e.g., a line).
-An adaptive SMC based on CRT is proposed by Alqumsan et al.
-(2019).
-The controller is designed for planar tip position control, which considers the saturation of the actuation sources.
-Furthermore, the marginal stability of the proposed adaptive control law is provided.
-The validation is conducted in simulation, where the PDE of the Cosserat rod are solved by the Generalized- method (Chung and Hulbert, 1993), an implicit solver that assures unconditional stability and numerical dissipation for high-frequency noise.
-Subsequently, the high nonlinearity of a large-scale and ultralight two-section inflatable CSM made of TPU-coated Nylon is captured by a dynamic Cosserat rod recursively to predict the behavior of the thin-walled and incomplete continuum structure (Li et al., 2021).
-An extended inverse dynamic model compensates for the error in a feed-forward manner.
-Moreover, stiffness and damping matrices are calibrated on deflection data and a weight release experiment, respectively.
-The study also compares the rod model versus a rigid-body and a PCC model.
-Finally, three closed-loop model-based controllers for trajectory tracking are compared to further demonstrate the viability of the CRT.
-Here, an adaptive SMC generated the desired pressure differences in the presence of model uncertainties.
-The chattering phenomenon caused by the conventional SMC is attenuated by integrating a fuzzy logic control, using continuous membership functions to replace the discontinuous .
-Mishra et al.
-(2023) proposed a Terminal SMC (Venkataraman and Gulati, 1993) for tip position TT control.
-The model on which the controller is based consists of a combination of CRT and fractional Bounc-Wen model (Mishra et al., 2022), taking into account the hysteresis effect.
-In this work, the Terminal SMC is robust, fast, non-singular, and with exponential convergence.
-In addition, an obstacle avoidance algorithm is presented, using a combination of cumulative danger field and damped least square methods.
-The model and the controller are validated in a simulated version of Robotino-XT (Wang et al., 2022a).
-5.1.7 Model Predictive Control
-The Model Predictive Control (MPC) (Rawlings et al., 2017) is a well-known controller based on iterative and finite-horizon optimization.
-Furthermore, it is possible to optimize while satisfying constraints, such as actuators’ saturation.
-Towards MIS, Soltani et al.
-(2017) introduce a 3D, stable, non-linear model-based hybrid force/position controller to apply a constant force to the cardiac tissue while tracking the desired trajectory.
-A quasi-static Cosserat rod models the distal shaft of tendon-driven catheters, reformulated for controller design.
-The position controller is based on a non-linear MPC tracker satisfying input constraints and local stability conditions.
-The position control operates in an inner loop, while the force control is in an outer control loop.
-Simulations conducted with a bidirectional ablation catheter predict the control performance applied in catheterization.
-Finally, experimental validation is carried out on a real pig heart, though the authors suggest the necessity of more computational efficiency and incorporation of imaging modalities with position tracking systems.
-Recently, a novel motion planning method of fish-like soft robots actuated by macro-fiber composite pairs was developed (Barbosa et al., 2023).
-The structure should mimic oscillatory and undulation movements, accomplished if the amplitude of the tail motion is greater than that of the head motion.
-The motion planning adopts MPC for generating the actuation signals.
-This model-based strategy requires a small-sized model.
-Therefore, they extracted a linear modal state-space model from a free-free dynamic EBRT considering the electro-mechanical coupling of the actuator pairs.
-Moreover, design strategies like the addition of concentrated and distributed masses were addressed to mimic fish-like motion, guaranteeing suitable mode shapes for the structure.
-Four numerical case studies demonstrated the capability of the proposal for deriving bounded input signals that generate oscillatory and undulation movements even with disturbances.
-5.1.8 Miscellaneous
-An open-loop optimal control method to control an octopus-inspired planar arm modeled with CRT was proposed (Wang et al., 2021b).
-The main goal is to use the optimal control approach to replicate two stereotypical motions: (i) reaching (i.e., bending propagation from the base to the tip) and (ii) fetching (i.e., creation of pseudo-joints).
-The functional cost of the optimal problem is the elastic potential and actuation energy of the arm, with a terminal cost related to the specific task.
-An iterative forward-backward algorithm solves the optimization.
-The proposal validated in silico demonstrated the emergence of propagating waves, as observed in stereotypical strategies.
-Conversely, Doroudchi et al.
-(2018) designed three controllers based on the dynamics of a linear Euler-Bernoulli rod.
-Thanks to the linearized form of the EBRT, the authors can apply the optimal control theory easily.
-The novelty here is that the optimal controllers are decentralized, with a modular soft robotic arm where each piece has independent actuation and sensorization.
-They observe that ill-conditioned system matrices, common in discretized beam models, may hinder control convergence.
-The iterative algorithm is also very sensitive to the initial condition.
-Furthermore, a third controller based on the gradient descent method provides a bounded solution for the controller gains and reduces the sensitivity to initial conditions.
-Finally, simulations validated the approach.
-To control aerial continuum manipulation systems, Hashemi et al.
-(2023) proposed a terminal synergetic controller to accomplish tip position TT.
-After presenting a novel platform that integrates a quadrotor with a tendon-bent CTR, they designed a two-loop control architecture.
-In particular, the first loop implements a terminal synergetic controller to ensure the continuum robot’s tip follows the desired trajectory.
-The second loop’s controller manages the quadrotor’s position and velocity.
-In addition, this last control algorithm is based on a combination of synergetic control and a hybrid feedback law.
-For the first time, a synergetic controller is applied in the literature, showing robustness against disturbances and model uncertainties.
-Finally, they validate the control framework simulating an aerial continuum manipulation system.
-Controller | References | Rod Model | Task | Validation | ||||||||||||||||||||
-Inverse Kinematics |
-Inverse Dynamics |
-Feedback Linearization |
-Energy-Shaping |
-Sliding Mode |
-Model Predictive Control |
-Open-Loop Optimal Controller | Wang et al. (2021b) | CRT (DER) | Reaching, Fetching | Numerical | ||||||||||||||||||||
-Doroudchi et al. (2018) | EBRT | Shape Regulation | Numerical | |||||||||||||||||||||
-Synergetic Control | Hashemi et al. (2023) | CRT | Tip Position TT | Numerical |
-5.1.9 Shape/State Estimation
-As seen in Sec.
-5.1.1 and in Fig.
-7, a general MB shape controller requires feedback in terms of configuration vector and its time derivative .
-Depending on the spatial discretization, the vector could be difficult or impossible to measure.
-To tackle this issue, a general MB control framework requires an Estimator, which maps the measures in the estimated state .
-For continuum and soft robots, this problem is referred to as Shape/State Estimation or Shape/State Reconstruction.
-Currently, most Shape Estimation methods use geometrical models or Computer Vision algorithms (e.g., (Bezawada et al., 2022; AlBeladi et al., 2021)).
-However, incorporating prior knowledge of rod-based models can significantly enhance performance.
-In the pioneering work of Trivedi and Rahn (2014), the Authors suggest three approaches for Shape Estimation, based on CRT: (i) mounting load cells at the base, (ii) employing cable encoders, and (iii) mounting an inclinometer at the end of each piece.
-A maximum tip position error of is observed in simulations.
-Lastly, they use the OctArm VI (Grissom et al., 2006) to experimentally validate all three approaches.
-Another study proposes a real-time shape-estimation method based on the force-torque for CTR measured at the tubes’ basis (Donat et al., 2021).
-It extends a shape estimation algorithm for elastic Kirchhoff rods.
-They modeled a CTR combining planar PCC segments lying on different equilibrium planes.
-The approach is evaluated with single and two combined additively manufactured tubes at high frequency, achieving a mean deviation of mm along the tube.
-Differently, Lilge et al.
-(2022) develop a model for shape and strain estimation of a continuum cable-driven robot using a Gaussian Process regression, a mathematical tool for estimating continuous-time trajectories in .
-The idea involves substituting (i) time with arc-length and (ii) kinematic and dynamic laws based on CRT.
-This method efficiently estimates the robot’s shape using noisy measurements from sensors such as strain gauges, tracking coils, and an external camera.
-The real-world experiments provide excellent performances, with an average tip pose error of mm and .
-Finally, Yousefi et al.
-(2023) propose an optimization-based method that simultaneously estimates the shape and the forces acting on a continuum robot by employing the quasi-static CRT.
-Magnetic localization determines the position of multiple robot points.
-The method estimates the robot’s shape and force in a wide range of conditions validated experimentally, with and without knowledge of the contact positions.
-5.2 Learning-based Controllers
-Learning-based controllers leverage the physics of rod models and utilize Machine Learning (ML) to derive parameterized control policies , which provide and end-to-end mapping between desired poses and sensor measurements to motor commands :
-The vector of policy parameters are the weights of ANN trained via Supervised Learning (SL) or RL.
-We report learning-based controllers in chronological order to highlight the evolution of the literature.
-5.2.1 Supervised Learning
-SL is viable to address the soft robotic control problem for simple tasks. It relies on pseudo-random motion data of actuations and corresponding task space to train an inverse model (50).
-Initial learning strategies used rod models to rapidly prototype controllers.
-For instance, Thuruthel et al.
-(2017) developed dynamic models of CSM using recurrent neural networks and presented a trajectory optimization method for task space control.
-They validated the controller on a Cosserat rod with PCS parameterization and a pneumatic CSM.
-Recently, Wang and Rojas (2022) developed a closed-loop controller of continuum robots, employing RNN to approximate the forward and inverse dynamics.
-They extended the CRT in (Till et al., 2019) to account for spine compression and collect motion data for training their forward models.
-A non-parametric Gaussian process regression compensates for the error between the real robot and the RNN.
-A hybrid controller alternates between two stand-alone policies.
-They validated the control architecture on reaching and tracking tasks using a tendon-driven continuum robot with 3D-printed compressible spines.
-The results show that combining simulated and experimental data improves the control accuracy.
-5.2.2 Reinforcement Learning
-RL is one of the main ML paradigms and a versatile framework suited for decision-making and control (Sutton and Barto, 2018).
-The two fundamental objects of an RL problem are the agent and the environment.
-In soft robotic control, the agent is a learning-based controller implementing an ANN, whereas the environment is everything outside the agent, comprising the soft robot and other objects.
-Fig.
-8 depicts the agent-environment interaction highlighting the role of ANN and rod models.
-The agent and the environment interact at discrete time steps .
-At every interaction step, the agent observes the (possibly partial) state of the environment including sensor readings and task information.
-The agent acts according to a policy to select actions taking the form of low-level controls like torques or pressures.
-A policy can be a deterministic mapping from states to actions, .
-Alternatively, a policy can be stochastic, mapping a state to a distribution over actions .
-The environment transitions into another state according to the dynamics of the rod model.
-The agent also perceives a scalar reward signal from the environment, indicating the goodness of the current state specifying the control task.
-The goal of the agent is to maximize its return, the cumulative sum of rewards:
-where is a discount factor. Therefore, the goal of RL is to find an optimal policy that maximizes the expected return
-The following deep RL contributions are summarized chronologically in Table 6.
-Satheeshbabu et al.
-(2019) applied deep Q-learning (DQL) with experience replay for the quasi-static position control of a pneumatic CSM capable of bending and twisting.
-An open-loop policy was trained in simulation utilizing a static Cosserat rod.
-They validate it in simulation and on the physical platform, subject to various external loads.
-To cope with the high-dimensional state-action space in soft robotics, DQL (Mnih et al., 2013) employs an ANN to approximate the value function of state-action pairs, defined as
-A deterministic policy can leverage to select the best action in a state as .
-The following work increased the dexterity of the CSM and trained a closed-loop controller for precise quasi-static positioning via deep deterministic policy gradient (DDPG) (Satheeshbabu et al., 2020).
-Experiments in simulation investigated the robustness of the control policy for effects of loading, reachability, and workspace discontinuity.
-Finally, the controller was deployed on the prototype.
-The DDPG algorithm also derived a closed-loop controller for the quasi-static positioning of a pneumatic CSM modeled with the KRT (Uppalapati et al., 2020).
-Interestingly, the CSM was integrated into a mobile platform including a rigid arm and a sensorized gripper.
-The system was teleoperated and validated in the agricultural task of picking berries using different maneuvering strategies.
-The potential of continuum mechanical models is fully exploited in physical interaction tasks.
-Naughton et al.
-(2021) effectively applied several deep RL algorithms to train various control policies.
-The control tasks included point reaching, trajectory tracking, and maneuvering through structured and unstructured obstacles.
-The controllers were learned and tested in simulation using a synthetic CSM based on a dynamic CRT.
-The rod was actuated by applying distributed internal torques, modeled via splines characterized by control points and vanishing values at the rod ends.
-The work showcased a captivating interaction between the rod and the environment.
-In particular, the maneuvering task required the free end of the rod to reach a position behind fixed obstacles.
-The task was solved with a reward function that only considered the distance to the target and a penalty term for the occurrence of numerical errors
-Interestingly, the rod learned to navigate and interact with the obstacles without an explicit reward term.
-Finally, Alessi et al.
-(2023b) applied proximal policy optimization (PPO) to train a closed-loop position control policy for dynamic trajectory tracking with a pneumatic CSM.
-The controller was trained in simulation leveraging a dynamic Cosserat rod model of the CSM (Alessi et al., 2023a).
-After training, simulation tests evaluated the generalization capabilities of the policy to new observations, dynamics, and tasks.
-The experiments included tracking new trajectories subject to unknown external forces or using different material properties (e.g., ).
-In addition, the learned controller was transferred without retraining to intercept a moving object.
-The policy has shown good generalization capabilities for all four tests.
-Subsequently, Alessi et al.
-(2024) elaborated on the previous simulated environment to account for the dynamics of a dexterous CSM and its interaction with objects.
-They developed a learning-based control architecture for pose/force control, which enabled dynamic pushing with CSM in the real world.
-To mitigate the significant sim-to-real gap in soft robotics, they originally applied domain randomization (DR), an effective sim-to-real technique (Zhao et al., 2020b), emphasizing soft materials.
-Controller | References |
-Task |
-Open loop ANN | Satheeshbabu et al. (2019) | CRT () | DQL | Tip Position TT (way-point) | (None) | ||||||
-Closed loop ANN | Satheeshbabu et al. (2020) | CRT () | DDPG | Tip Position TT (way-point) | (None) | ||||||
-Uppalapati et al. (2020) | KRT () | DDPG | Tip Position TT (way-point) | (None) | |||||||
-Naughton et al. (2021) | CRT () | Several | Tip Pose TT, Obstacle maneuvering | ||||||||
-Alessi et al. (2023b) | CRT () | PPO | Tip Position TT, Object interception | ||||||||
-Alessi et al. (2024) | CRT () | PPO | Tip Pose/Force for Pushing | (DR) |
-5.3 Discussion on Rod-based Controllers
-5.3.1 On the Model-based Controllers
-The emergence of computationally efficient and accurate models enables researchers to develop model-based controllers that can accurately predict and react to the nonlinear nature of the elastic rods.
-As shown in Table 5, the controllers based on Inverse Kinematics are the most popular choice, achieving tip trajectory tracking tasks in real-world prototypes.
-Despite the simplified model, the feedback action is able to reject the unmodelled effects, showing the usability of that strategy.
-However, a dynamic model is essential to exploit the properties of continuum robots.
+Recent rod models addressed manufacturing uncertainties considering variations in material (alessi2023ablation) and geometric properties (eugster2022soft).
+Research gaps persist despite these advancements.
+First, capturing hysteresis—a prevalent phenomenon in soft materials that affects the accuracy and predictability of the robot’s behavior—remains largely underexplored in rod models.
+Second, incorporating self-healing materials can potentially enhance the durability and longevity of soft robots (terryn2017self).
+However, integrating the dynamic restoring process into rod models should still be investigated.
+Filling these gaps could advance the capabilities of continuum and soft robots, enabling their integration into diverse applications and real-world tasks.
+9.6 On the Model-based Controllers
+The emergence of computationally efficient and accurate models enables researchers to develop model-based controllers that can accurately predict and react to nonlinear elastic rods.
+We observed that controllers based on IK are the most popular choice, as they achieve tip trajectory tracking tasks in real-world prototypes (Table 3).
+Despite the simplified model, the feedback action can reject unmodeled effects, showing the effectiveness of that strategy.
+However, we should note that most IK-based controllers treat the system as fully actuated.
+This limits control to a smaller set of DoFs than the robot is inherently capable of exhibiting.
+A dynamic model is essential to exploit the properties of continuum robots.
 Most of the proposed controllers adopt well-known approaches from the under-actuation literature, such as FL, ESC, SMC, or MPC.
-However, only a few proposed controllers presented an experimental validation.
-The reasons behind this sim-to-real gap could be the lack of robustness or physical limits of the prototypes.
+However, only a few proposed controllers were validated experimentally.
+The reasons for neglecting the sim-to-real comparison could be the lack of robustness or physical limits of the prototypes.
 While the Inverse Dynamics, FL, and ESC are sensitive to unmodelled effects and parametric uncertainties, the discontinuous output of the SMC is difficult to implement with the most popular actuation technologies (e.g., tendons, pneumatic).
-Specifically, the actuators’ internal dynamics may limit the effective bandwidth of the system, hence filtering the discontinuous output of the controller.
-The MPC is a promising but not widely explored controller in the continuum robots literature.
-Another relevant aspect of the analyzed works is the discretization technique employed.
-Two main aspects eased the control design: (i) the advancement of discretization techniques and (ii) the tunability of the accuracy.
-The former allows the researchers to design the controller with a finite number of state variables and compute efficiently the robot’s kinematics or dynamics.
-Selecting appropriate discretization techniques is crucial, particularly for determining the minimum frequency necessary to ensure numerical stability.
-In addition, the DER technique takes into account also cross-section deformations, especially useful for modeling and controlling muscles in the case of bio-inspired robotic systems (Chang et al., 2023).
-The latter aspect allows for the choice of accuracy levels and the number of strain modes the model can predict.
-A common approach simplifies the model and designs a controller that treats unmodeled strains as external disturbances.
-This strategy can enhance the speed of the control loop and its robustness.
-Regarding research gaps, it is possible to note the lack of experimental validation for MB controllers.
-Furthermore, there are few studies involving external forces due to interactions with environments, particularly because of the complexity involved in modeling and estimating contact and friction forces.
-Further developments on model-based controllers could benefit from using physics-inspired ANN to lessen the computational burden of complex expressions (e.g., the Coriolis matrix) or use them to compute the EoM of the elastic rod, as already explored in (Lutter and Peters, 2023; Liu et al., 2024).
-Finally, the literature on rod-based planning (Mishra et al., 2023; Bentley et al., 2023) is sparse, likely due to the current maturity of the controllers.
-Recently, some planning algorithms using only geometrical approaches have been proposed (Rao et al., 2024).
-5.3.2 On the Learning-based Controllers
+Specifically, the actuators’ internal dynamics may limit the effective system bandwidth, hence filtering the discontinuous output of the controller.
+The MPC is a promising but not widely explored controller in the continuum robots literature, which could be attributed to the relatively high computational cost.
+Two key aspects facilitate the design of controllers for CSM: (i) the development of advanced discretization techniques and (ii) the tunability of model accuracy.
+The former enables the design of controllers with a finite number of state variables and allows for efficient computation of the robot’s kinematics and dynamics.
+Choosing appropriate discretization techniques is particularly critical for determining the minimum control frequency required to ensure numerical stability.
+Furthermore, the DER technique captures cross-sectional deformations, which are especially relevant for modeling and controlling muscle-like structures in bio-inspired robotic systems (chang2023energy).
+This feature was recently integrated within the GVS framework by sun2025real, who proposed an extended CRT formulation that models radial (in-plane) deformations through an additional strain variable.
+Their model also incorporates nonlinear hyperelastic and viscous effects, extending the GVS approach to more complex, bio-inspired dynamic scenarios.
+Equally important is the ability to tune model accuracy by neglecting specific modes or reducing the number of DoFs, as implemented in the Strain Parameterization.
+This flexibility allows trade-offs between accuracy and computational complexity while enabling the controller to treat unmodeled dynamics as external disturbances.
+Such a strategy increases the control loop frequency by reducing computational load and enhances its overall robustness.
+Regarding research gaps, we observed the lack of experimental validation for MB controllers.
+Furthermore, there are few studies involving external forces due to interactions with the environment, particularly because of the complexity involved in modeling and estimating contact and friction forces.
+Due to the current lack of maturity of the controllers, the literature on rod-based planning (mishra2023trajectory; bentley2023safer) is still sparse, although planning algorithms using geometrical approaches were recently proposed (rao2024towards).
+Future developments on model-based controllers could benefit from using physics-informed ANN to lessen the computational burden of complex expressions (e.g., the Coriolis matrix) or use them to compute the EoM of the elastic rod, as already explored in (lutter2023combining; liu2024physics).
+Another promising future direction is the application of the Collocated Form proposed in (pustina2024input) to the design of more advanced controllers, particularly those inspired by the rigid underactuated robotics literature.
+This specific change of coordinates can simplify feedback implementation, since the actuator lengths are directly measurable and easier to estimate, as also exploited by feliu2024dynamic.
+Furthermore, adopting this coordinate transformation can enhance IK-based controllers by naturally accounting for the system’s under-actuation.
+Finally, differentiable simulators and libraries for rod models could be utilized for optimal control.
+Recently, analytical derivatives for the GVS model were calculated in (mathew2024analytical), facilitating the implementation of model-based optimal controllers.
+9.7 On the Learning-based Controllers
 Learning-based controllers utilize the physics of rod models in various applications, with a clear distinction between SL and RL.
@@
 Therefore, this approach could fade away in the future.
-Conversely, the role of rod models is central for developing complex deep RL control policies, especially to achieve unstructured physical interaction with CSM.
-Deep RL agents exploited computationally feasible rod models embedded in simulated environments to learn optimal iteratively.
-Herein, RL not only replaces the burden of collecting and labeling datasets offline but also unlocks a new range of possibilities not fully explored yet.
-The literature review highlighted the role of rod models in RL policies.
-Indeed, as reported in Table 6, control tasks evolved from position control for quasi-static tracking to simulated pose control for obstacle maneuvering and pose/force control for dynamic pushing.
-For instance, initial policies limited the physical interaction to static payloads without addressing the sim-to-real gap (Satheeshbabu et al., 2020).
-On the other hand, efforts in developing simulated environments (Naughton et al., 2021) combined with careful reward engineering and sim-to-real techniques (Alessi et al., 2024) facilitated recent advancements in physical interaction with CSM.
-In summary, rod models combined with recent deep RL methods and sim-to-real transfer techniques (Zhao et al., 2020b) advanced the state of the art of soft robot control.
-Future developments could explore model-based RL.
-Here, the term model-based refers to the agent not only learning a policy but also learning a predictive model of the environment used for planning (Moerland et al., 2023).
-Alternative research lines can be borrowed from rigid robotics and tailored to the nonlinearities of continuum and soft robots (Ibarz et al., 2021).
+Conversely, the role of rod models is central for developing complex deep RL control policies to achieve unstructured physical interaction with CSM.
+Deep RL agents exploited computationally feasible rod models embedded in simulated environments to learn optimal actions iteratively.
+Herein, RL reduces the burden of collecting labeled datasets offline and unlocks new possibilities that are not fully explored yet.
+The control tasks addressed evolved from position control for quasi-static tracking to simulated pose control for obstacle maneuvering and pose/force control for dynamic pushing (Table 4).
+For instance, initial policies limited the physical interaction to static payloads without addressing the sim-to-real gap (satheeshbabu2020continuous).
+Conversely, efforts in developing simulated environments (naughton2021elastica) combined with careful reward engineering and sim-to-real techniques (alessi2024pushing) facilitated recent advancements in physical interaction with CSM.
+In summary, rod models combined with recent deep RL methods and sim-to-real transfer techniques are promising to advance the manipulation capability of continuum soft robots.
+We should note, however, that these works relied on motion capture systems to track the robot positions.
+Indeed, reliable soft robotic proprioception is still a challenge (pagliarani2025softtex).
+Future developments could explore model-based RL, where model-based here refers to the agent not only learning a policy but also learning a predictive model of the environment used for planning (moerland2023model).
+Alternative research lines can be borrowed from rigid robotics and tailored to the nonlinearities of continuum and soft robots (ibarz2021train).
 10 Conclusion
-This paper provided an overview of the modeling and control of continuum and soft robots using rod theories.
-First, we formulated the four main rod theories used in robotics under a unified mathematical framework.
-Subsequently, we provided mathematical tools for modeling the actuation in soft robotics.
-This was followed by a classification and summary of the rod-based models for continuum and soft robots across nine classes of deformations, offering a unique perspective on the evolution of the state of the art, together with a discussion on research trends and gaps that emerged.
-Furthermore, we reviewed recent control strategies leveraging rod models to perform various tasks, followed by a discussion on the encountered approaches as well as our insights.
-Through an in-depth exploration of various rod-based models and control strategies, the diverse range of applications of rod theories demonstrates their capabilities to capture the complex mechanical behaviors of continuum and soft robots.
-Despite the remarkable advancements in incorporating rod theories into soft robot models and controllers, there are still areas requiring further research.
-For instance, accurate modeling of processes inherent in soft materials, such as hysteresis, is still a challenge.
-Related to the control, addressing efficient sim-to-real policy transfer in learning-based approaches is essential for real-world robot deployment.
+This review paper explored the modeling and control of continuum and soft robots with a focus on rod theories.
+We surveyed the state of the art with a vertical approach spanning mathematical formulations of rod theories, rod-based models of continuum and soft robots, and control strategies with rod models.
+The review of the mathematical background facilitated comparison between different variants of rod theories while highlighting connections to other models.
+A comprehensive review of rod-based models uncovered their versatility in supporting various studies, from bioengineering design to experimental validation of continuum and soft robots.
+Meanwhile, our literature classification on deformation classes offered new perspectives on the evolution of the state of the art, especially in relation to robot deformations and the model.
+The review of controllers highlighted the direct applicability of advanced model-based algorithms for trajectory tracking tasks, while RL coupled with sim-to-real strategies emerged as a promising approach for manipulation.
+Despite remarkable advancements in integrating rod theories within soft robot models and controllers, limitations persist.
+For instance, modeling hysteresis effects of soft materials and comprehensive experimental validations in contact scenarios remain challenging.
+Related to the control, addressing efficient sim-to-real policy transfer is essential for robot deployment.
+This paper offered a comprehensive treatment of rod models in continuum and soft robot control.
+Nonetheless, we believe our insights can benefit various soft robotics communities and could serve roboticists dealing with continuum or soft structures.
 This work was supported by the Future and Emerging Technologies (FET) program of the European Commission, through the PROBOSCIS project (G.A.
 ID 863212), by the US Office of Naval Research Global under Grant N62909-21-1-2033, and in part by Khalifa University under Awards No.
-RIG-2023-048, RC1-2018-KUCARS
+RIG-2023-048, RC1-2018-KUCARS, by the Singapore-Italy Cooperation Project DESTRO (A*STAR grant number R22I0IR124, _______ MAECI), by the MOE Tier 2 Project REBOT (grant number MOE-T2EP50221-0010), by the NUS grant RoboLife, and by the PNRR MUR project PE0000013-FAIR.
 Appendix A Query for models and controllers